A Complete Bibliography of Publications in Journal of Computational Physics: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 1.10

Title word cross-reference

(γ, n) [MY73]. (z − p^{2}/2m)^{−1} [HRY73]. 10 [Ano72d]. 11 [Ano73a]. 12 [Ano73b]. 13 [Ano73c]. 14 [Ano74e]. 18 [Ano75c]. 19 [Ano75d]. 2 [BD79]. 20 [Ano76b]. 21 [Ano76c]. 22 [Ano76d]. 23 [Ano77b]. 24 [Ano77c]. 25 [Ano77d]. 26 [Ano78c]. 27 [Ano78d]. 28 [Ano78e]. 29 [Ano78f]. 30 [Ano79b]. 31 [Ano79c]. 32 [Ano79d]. 40 < Z < 90 [Ban73]. 5 [Ano70a, MN74]. 6 [Ano70b]. 7 [Ano71a]. 70 [DGH73]. 8 [Ano71b]. 9 [Ano72c]. ^{3} [DGH73]. _{2} [PPS74]. B [AK72, Dav78]. χ [SG79]. δ(μ − μ_{0}) [HBH78]. E_{1}(x), x > 0 [SZ77]. γ [Chi71]. H [EJK^{+}73, Nes71]. I_{n}(x + jy) [Sca71]. ∫_{0}^{∞} t^{n} exp(−t^{2} − x/t) dt [CP79a]. J_{n}(x + jy) [Sca71]. K_{iζ}(x) [BO75]. L^{2} [HRY73]. m^{†}(h − ϵ)^{−1}m [Nes71]. μ [PPP76]. N [AC73, MC76, Mil70b, RSM70, Ves77, FJ71a, Mil71b, Mor78, RCB77, Str77]. P [GW78]. ϕ^{3} [HH77]. Q [Vie73]. S_{2} [AHMB74]. S_{4} [AHMB74]. S_{n} [AHMB74, SB73]. σν [CMM^{+}78]. × [MN74]. V (a, x) [RL76].

-algebra [Chi71]. -alkanes [RCB77]. -body [Mil71b, AC73, Mil70b]. -D [BD79]. -dimensional [FJ71a, RSM70]. -molecules [PPP76]. -particle [Mor78, Ves77]. -splines [Dav78]. -Wave [GW78].

12-6 [RNP75].

2-D [For77b].

360 [Ama77, Col74].

60 [Tem79a].

A. [Saf78]. Abel [Bru73]. abnormal [WL71]. absorption [CW76]. Accelerated [CGR72]. Acceleration [Nis72, GT78b]. accelerator [JD70]. Accuracy [Hof73, Kov70b, Abr75, iCS78, DGH73, GM79a, Ng74, OS70, Rus70, SC79, Ste79a, Ste75, TAG76, WH74]. Accurate [Coh78, Kni76, Thi78, Abr73, Ada77, BJ79, Dic72, EGZ72, Gaz73, Gaz75, HZ79, Hau79, Hir75, MM73, MM74, OS74, PR72, SZ77]. acid [MS70a]. acoustic [BMP^{+}72, CC79a, CT77, FM78]. adapted [Ben79]. adaptive [Lep78]. Addendum [Ano72a]. additional [Bel75]. ADI [HD74, She79, Smi75b]. adiabatic [BNGF73]. adjoint [CHS78, HL75]. adjusting [HZ72]. Adjustment [WGS72, WS73]. admittance [Blu71a]. adsorption [RNP78]. advection [KB79a, MB77]. advective [GL76]. aerodynamics [Bel70]. aide [Lav73]. Airy [SAG79]. albedo [HSM74, KSV70, Shu73]. Alfvén [BV76]. algebra [Chi71]. algebraic [Ixa72]. ALGOL [Tem79a, PKPR72, RB71]. Algorithm [MS78b, Mor78, Nes65, SAG79, Tem79a, AAL79, AW74, BS79, BW76, BB73, BKL75, BP73, Cha76b, CT77, CC79b, Fri78, GL76, HJKS78, HP78, Kal76, Lep78, Les73, Nis72, PKP79, Pex76, PB76b, Raf79, Raj73, SL72, Sha70b, Two75]. Algorithms [Chi71, Per78, Tem75b, BL73, BB76, Eas75, Han72, Lan73, Lew72, MC78, Run75, Zal79]. alkanes [RCB77]. almost [BP73, NR71]. almost-Lagrangian [BP73]. along [Bel75]. Alterman [BL78]. Alternating [LK73, Gus71, NC71]. alternative [Ben79, VW77]. Ambartsumian [Cor75]. ambipolar [Fre73]. amplitude [CS70, Dal77]. amplitudes [CP76, dAPM77]. analog [PPP73, PPP76]. Analysis [Gil72, Lan79, VT77, ABGR75, BR71a, Bus75, CC79a, CH70, CH73, CLS78, DG76, Dow70, Eis78, EJK^{+}73, Fle73, GJ73, Han72, HLB79, Ixa72, JN71, Jon79, KB79a, KDS79, KL76, MPB73, MS70b, PF74, Pes77, PMB73, SV76, SK78, SS78, Sto77, Tak76, VP75, WH74, WO75]. Analysis-of [VT77]. Analytic [Bun71, Bun73a, FH74, OCS72, BBA73, FPW73, Ter79a]. Analytical [AHMB74, Sha70a]. analyzing [TR76]. angle [Dav79b, Hol77]. angular [BB71, MBM76, PL70]. anisotropic [And75, CS76, HK72, KSV70, MN73, Shu73]. Announcement [Ano72b, Ano74a, Ano74b, Ano75a, Ano76a, Ano77a, Ano78a, Ano78b]. Announcements [Ano79a]. annular [MS79b]. annuli [Sch75b]. annulus [Quo76b]. any [CP75, TTM74, TTM77, Wil78]. Appearing [Squ70, GL77]. Application [AKK70, FRT79, MC76, Pex76, RL72, SOH72, Ste78b, Tak76, TM70, BL72, Chi76, DE70, ER74, HU74, KDS79, MB77, MGRK73, PH78a, Pfe79, SE76, Ste76a]. Applications [Lyn76, Bal76, Cho77, Smi75b]. applied [AM79, AG75, Bak79, DT77, Gar76, GH73, HH78, HD79, Jon79, PPP73, Sul72, Tak71, Two75]. applying [Wei78]. approach [ABGR75, BLV76, Ben79, BMP73, BMP^{+}72, CS77, Cri77, GA72, GM79b, KB77, MPS75, ML73, Mor70a, Pev71, PS72, PB75, Sam71, SBD75, Smi75a, Tro76, Van77b, WH74, ZG73]. approaching [Bey71]. approximants [FPW73, Gru79]. approximate [SS71]. Approximating [KP76]. Approximation [Squ70, Cam72, CJ74, Cho78b, Chu71, FJ71b, Fos73, FO73, GW73, Hai79, Har76, TG77, Unc74]. approximations [Bal76, BG70, Chi74, Chi75, CGR72, CR75b, FGW73, GH77, GM79a, Lew70, OJ74, RB77, Ste76b, WO75, Wil71a]. Arakawa [Jes74]. arbitrary [Cha75, CC70b, DS75, FB76, HAC74, RS72b, Sca71, SS76, TTMW77, TI79, TTM74, TTM77, WJW78]. areas [Bud79]. Argon [RNP75]. argument [dAPM79, Sca71, de 77]. arising [All70, Ede73, KH75a, Lon72, SH71b, Voj73, WB77, Wim78]. Aron [Sun75]. article [Chi74, Voj73]. articles [Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q]. artificial [GK79, Lut71, Roa72, SG79, Str78, Whi73a, Whi73b, Whi74]. aspects [AS72]. assemblies [KB71]. assessment [SGP76]. associated [Cam71, Chi75, Mor70b, Mor71]. Asymptotic [HK76, Kaz70, AH73c, CS76, Che71a, ON73]. atmosphere [Kan76, MB77, MY71]. atmospheric [WS78]. atom [FRT79, HP75, LNRY73, Smi75a, TO79]. atomic [Bru74, CP75, CFU74, Jos71, Mil74, ML73, Per75, WGS72, WS73]. atomic-like [ML73]. atoms [Ban73, Bur78, Fro78]. attainment [FR77]. attempt [CP71]. Author [Ano70a, Ano70b, Ano71a, Ano71b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano74e, Ano74c, Ano74d, Ano75b, Ano75c, Ano75d, Ano76b, Ano76c, Ano76d, Ano77b, Ano77c, Ano77d, Ano78c, Ano78d, Ano78e, Ano78f, Ano79b, Ano79c, Ano79d, Ano79e]. Automated [CF78a, RT79b]. Automatic [PKPR72, Sas76a, SS79, TTM74]. Automation [HH77, Ros74, Ros76a]. Average [Ban73]. averaging [SW76]. axial [Pac74]. Axisymmetric [And75, BKM73, Coh76, GJA79, GJ73, GJ75, Gru78, Hof73, JJGG78, JDG^{+}79, MS74]. azimuthally [HSM74].

Ball [War79]. Banach [Mor79]. baroclinic [Kan76, MP77, Quo76b, Yee72]. barotropic [Cro70a, Cro70b, Kan76]. based [Gaz73, Mil70a, Mor79, Run75]. basis [AMHH73, Dav78, ED74, FB76, KP76, PS73, PH78b, Rei74a]. beam [BTS79, WSW78]. beams [WJW78]. bedrock [Båt78]. behavior [Foo73, Kaz70]. behind [Pat78]. bell [TM70]. bell-shaped [TM70]. belts [Dav79b]. BERKELEY [LBM^{+}78]. Bessel [BO75, MB71, Sca71, Tal78, Tem75a, Tem76, Tem79a, Wil71c, de 77]. beta [And75]. Bethe [Fle73]. between [Che71c, DS78, HW75, Hof74, LM73, dAPM79, VW77]. bicharacteristics [CH73]. biharmonic [GM79b, Jac73]. binary [TA77, WCK71]. binding [PPP73]. biography [BL78]. birefringence [Net71]. Blasius [WB77]. blended [Cav75]. Bloch [Kal76]. block [BE76, DMWV75]. block-implicit [DMWV75]. blood [Pes77]. blunt [JPC72]. bodies [Bra76, BMP^{+}72, Chu70, PL75, TM70, TTMW77, TTM74, TTM77]. body [AC73, BWSE76, JPC72, Kno75, Lav73, Mil70b, Mil71b, OFM71, Sco79, TTMW77, TTM74]. body-fitted [TTMW77, TTM74]. boiling [Ste79b]. Boltzmann [SK77, Taf76]. Boolean [Les73]. Borel [Gru79, Wim78]. Borel-type [Gru79, Wim78]. born [Ste76b]. both [Sca71]. bottleneck [Ben72]. bound [BLV76, CR75b, Lav73]. boundaries [ML77, SW70]. Boundary [CS74, CB71, GT78a, GM79a, GK79, Hau79, Rid79a, Sun75, Ada77, AW74, BJ77, Bos70, BGP73, BL72, CK71, Ceb79, CKW79, Cel76, Che73, Chi75, CN75, Cho78b, CF78b, Dav73, DL72, Det79, Dic73, Eas72, FRSW79, FPM76, Gar78, GH77, Gus80, Hai79, HP76, Hed79, HD79, JB75, Jon73, KC72, KH75b, LC72, Lin77, Lin75b, MS78a, MAT74, NTK79, NG72, NR79b, NH71, Orl76, PB76b, Quo76a, Roa70, RS72a, SS76, Slo77, Smi74, SS79, Tay70, Thi78, TTM77, TR76, Tro76, Tru72, WM72, WH75]. Boundary-distribution [Rid79a]. Boundary-fitted [Hau79, TTM77]. boundary-layer [CK71, CKW79, KC72]. boundary-value [Chi75, Det79, MAT74, NG72, NR79b]. bounded [DD79, HG79, Vie71]. bounds [RSM70]. box [FRT79, LS76]. branching [KH75a]. Brandow [Kal76]. bremsstrahlung [MY73]. bremsstrahlung-produced [MY73]. brickwall [Bey71]. Brillouin [HMB73, Hol77, LH72, Sin72]. Brillouin-zone [LH72]. Bubble [Gaz70, LWH78]. Bubble-chamber [Gaz70]. buckling [AKK70]. bundles [Mey75]. Burgers [CG77].

C [Sun75]. Cagniard [Abr78]. Calculating [Dav79b, HC72, NH73, Pra75, Bai78, BK73, Cha76a, DIC79, FC71, Gil72, God79, HCB70, HN70, HP78, PB75, Pra71, Rit70, Ste70, WCL78, Wil70]. Calculation [And71, Bel75, BS70, CM73, CS76, DS75, DS78, Fel74, Ore73, Ste79b, Wei78, Abr73, Bar73, Bro79, BK76, CZ79, CCQ78, Cho78a, Cri77, Cri78, Dav75, Ein76, GL77, HA71, HA75b, HRY73, JCM74, Kau72, Kau73, LR74, McD74, MS77, Mil70b, OS70, OFM71, Per78, PCJ76, PPP73, PPP76, Pop78, RL76, RDW74, RT79b, RG75, RNP75, SBPH73, Shu73, Sni72, Ste79c, Vie73, War78, WK79, WL71, ZG73]. Calculations [WSW78, AH70, Bar70a, Bas72, Bee76, Ben72, BS73, CKW79, Cho71a, Cla73, DGH73, FR77, Fin78, Har71, HA74, HAN76, JE74, Joh73, Lad71, LK73, MT78, NH71, OCS72, PT75, PL70, RB72, RNP78, RK77, Tol75, TAE73, Yos73]. Campylotropic [AS78]. Canonical [God75, Ros71, RNP75, RNP78]. capacitance [CM73]. capacitor [CM73]. capillary [Bro79]. Carlo [Bir71, Ben76, BL75, BKL75, BW79, CD78a, CHS78, Che71b, Cho71a, Cho71b, CM76b, FC71, FLRZ71, Gil78, GS71, HLM^{+}76, HP78, KK75, Kno75, MH79, Mat70, MT78, RR70, RNP75, RNP78, TO79, TV77, VW77, WK79, WL71]. Cartesian [AMHH73, MS78a, MD78, RCB77]. cascade [Eis78]. cascades [Kor78]. case [BBA73, Det79, Ros74]. Cashion [Dic73]. catalytic [EGV78]. Cauchy [LM74, TI79]. Cauchy-type [TI79]. cavitation [LWH78]. cavity [BD79, BD73, God79, Quo76a]. Cebysev [EEJ79]. cell [Bak79, HAN76, HR74, LSW72, SS71]. center [BS73, Bot70, CFUW74, JKAR77, MPS75, PS76, RT79a, SK75, Sha70a]. Centered [LS76, Bel75, Van77a]. centering [For73]. centerline [GJA79]. certain [CD78b, NG72, SL72, SW75]. chains [FD79, JD74]. chamber [Gaz70]. Chandrasekhar [Cor75]. changing [Per78]. channel [Mud76]. channels [Sch75b, Ste79b, Tol75]. characteristics [Cho73, GH77, Hof73, SV76, Van73]. characterized [MAT74]. charge [DS75, Kus73, MWW76]. Charged [AL76, Abr72, BTS79, MR73]. charged-particle [BTS79]. charges [Bel75]. charmonium [CS78]. Chebyschev [RB77]. Chebyshev [CCV72, HZ79, Hai79, PB76a, SL72]. checking [Ore73]. chemical [Ede73, Gil76, TAE73]. chemically [CH73, Wes78]. chemistry [AF79]. Cherenkov [God74]. choice [Cho76, Cho77, CP79b]. Cholesky [Ker78]. Cholesky-conjugate [Ker78]. Choosing [VW77]. Chorin [Saf78]. Chu [Sun75]. CI [LBM^{+}78]. circuit [BTS79]. circular [Bea71, Ceb79, Ken70, Pat78, Raj73]. circulation [HW75]. circulations [Cro70a, Cro70b]. class [Bru73, Cam72, CJ74, KB77, MM73, MM74]. classical [Ama77, Bru74, Col74, Sti71, Van74]. clausing [Cho71a]. close [AYMS71, Ale76, Bey71, ML71]. close-coupled [Ale76]. close-packed [AYMS71]. closed [RS76]. cloud [Bak79]. clustering [SS79]. clusters [Sto78]. CNDO [JCM74]. CO [PPS74]. code [BTS79, Bur78, BNPLD77, BK78, Dev74, HMR77, LTD79, OFM71, Sho78, TA77, TTM77, Whi79]. codes [CS74, CB71, God74, God75, Ste77b, Sun75]. coefficient [Dea71, Det79, HW75, Sal77, Sni72, SS71]. coefficients [Bel75, Del74, GW78, GW73, SS71]. Cohn [McK71]. Cold [Ros74]. colliding [Bra76, GDZK70]. collision [BL75, Hoc71, MS79b, TA77]. collisional [RD75]. collisionless [Lew72, Taf76]. collisions [Ale76]. collocation [Lau71, Rea70]. columns [AKK70]. combination [RS72a]. combined [vL74]. combustion [LC72, Wes78]. Comment [Cel76, Gus80, Win78, Cho78a, Saf78]. Comments [Bir71, Yee70]. common [dAPM79]. commonly [Roa70]. compact [Ada77, CLW78, Hir75]. Comparative [CGG^{+}78]. Comparison [BEH76, Dic73, EJK^{+}73, MKS77, RDW74, Ste77a, Tan79, Two75, Win78, And74, CCV72, CM76b, DM79, EJ73, GM70, LSW72, LN75, Wil71a]. Comparisons [DC74, Tem79b]. compensation [Sha72]. Complex [SAG79, HG79, KS73, Pfe79, Rig79, Sca71, TI79]. complicated [ML77]. composite [HU74]. compound [NR79a, NR79b]. compressed [FT77, NRH71]. compressibility [LL75]. compressible [BM77, Fle79, FE72, JB75, KW72, PR72, Raj73, SGP76, Van77a]. compression [BTS79]. Compton [SB73]. Computation [AK72, Cam72, CJ74, Che70, FK77, Lap71, Mos70, Pat71, PH78b, Sco79, Tem79a, TLM73, Abr72, AH73c, BM79, BJ77, Cho71b, CF78a, For73, FH74, Fox77, Gaz73, GL77, KV76, MN74, OB79, Ryb70, SFG71, TM70, Wal75, WB77, Wim78, ZR73]. Computational [iCS78, BNGF73, Che73, FE72, Gar74, GJA79, Lin75a, Lin77, MC76, SW76, SC79]. computations [CM76b, GM79a, HZ72, MH79, QB73]. compute [Gru78, Sca71]. computed [Job78]. Computer [CF76, Cum75, DD79, JN71, MC78, Aha72, BBA73, BK71, Blo73, CP71, CH70, CS70, FJ71a, Fle73, Hoc71, HGE74, Hof74, JD74, Lin70, Ore73, Ros71, Sto78, Sto77]. computer-simulated [Lin70]. Computerized [Boy74]. computers [EEJ79]. computing [Bud79, CMM^{+}78, EGZ72, Fre73, HAC74, Jen78, Pra75, TO78, Vie71]. concept [DE70]. concepts [Bru74]. condensed [HG79]. condition [CB71, HP76, Lin77, McK71, Orl76, Roa70]. conditions [Ada77, Bar78, BGP73, Che73, CS74, Eas72, Gar78, GH77, GT78a, GK79, Gus80, Hed79, Lin75b, McK71, NTK79, NH71, Roa70, SS76, Slo77, Sun75, Tay70, TR76]. conductances [Mos70]. conduction [Jon79]. cone [Dav79b]. cones [HN70]. confidence [Joh76]. configuration [Ben79, Bey71, CK76, Emp70, PH78a, Ste78a, Yos73, LBM^{+}78]. confined [BKM73]. conform [WGS72]. conformal [VT77]. conformation [Cri75, Cri77, SFG71]. congruential [HR78]. conical [Chu70]. conjecture [HL76, Ise73]. conjugate [Ker78]. connected [VT77]. Conservation [Lin75a, PW70, Ros76b, Vin74, BW76, CR78, MM73, MM74, Mey75, Sod78, vL74]. conservation-law [BW76]. conservative [AG72, Van77a, Van77b, Wei78, vL74, vL79]. conserving [KB79a, Lan73, Lew70]. consistent [BNPLD77, JCM74, Kus73, PT75, Ste70]. constant [Bel75, Det79, LL75, TAE73]. constant-coefficient [Det79]. constant-gain [TAE73]. constants [Mos70]. constitutive [Gat75]. constrained [Two75]. constraints [CP74, RCB77]. constructing [Fos73]. Construction [BDF78, Yos73, CM76a, Fin79, PT73]. containing [Coh78, CC70b, TTM74, TTM77]. containment [WJD70]. continuation [FPW73]. Continuous [PPP73, PPP76]. Continuum [RT79a, And71, Bot70, PS76]. CONTOR [RS72b]. Contour [Har71, ZHR79]. contours [RS72b]. contracted [PH78b]. contraction [GL78, Sni72]. Contribution [Job78]. contributions [FD79]. convection [DR79, HD74, LS76, PW70, ST76, Van77b]. convective [FE72, For77a, Gaz73, Men74]. conventional [Ben79]. Convergence [And78, CF77, Rig79, BD79, BDPH76, CGR72, Ise73, Nis72, SW75]. convergent [Kaz70, Run75]. conversion [GW78, Sal77]. converter [EGV78]. Cooley [Dic73]. Coordinate [GO77, HU74, And75, Cla77, Cla73, Eis78, Eis79, GCS75b, Raf79, TTMW77, TTM74, TTM77, Vin74]. coordinates [AS78, Cla73, Cri78, GJA79, GJ75, Hau79, HB77, Mey75, PT75, Pop78, PT73, WGS72]. core [Che71c]. cores [KK75]. Corner [WM71, SAK75]. corners [GW78, Rid79a]. corrected [BBH75, BB73, BB76, CC79b, Zal79]. Correction [Sni72, Stu78]. corrections [AF79, CP71, RNP78]. corrector [Hud72]. correlation [Ama77, Col74, KV76]. Correlational [Ama77, Col74]. Correlations [Blo76, PL70]. corresponding [Dav75, SBPH73]. Couette [GH73]. Coulomb [Bel75, Bot70, DS75, dAPM79, Stu78, Tol75, War78, Wil71c]. Coulombic [FD79]. coupled [Ale76, All70, BMP73, Gil76, Hat70, MdVD73, Smi75a, Tol75]. coupled-channels [Tol75]. coupling [SOH72]. covariance [Kov70a]. covered [MR71]. Cowell [Van74]. Crank [CCV72]. creep [And73, Pra71]. creeping [BBA73, ST76]. Critical [MPS75, WCK71]. cross [CM73, MY73]. crystal [Bel75, Blo73, Bud79, Flo77]. Crystalline [MS70a]. crystals [CC79a, CCQ78, MC78, Net71, PB75]. cubic [Bel75, CS73b, Fle74, FK77, HMB73, LH72, Net71]. cumulative [Fos73]. currents [Har78]. Curve [Che70, Squ70, Jen78, Pat71, SC72, Sul72]. curves [Job78]. Curvilinear [Lau79, AH73a, Bar70b, HB77, Mey75, PT75, TTM74, TTM77, Vin74]. curvilinear-orthogonal [Mey75]. cyclic [Det79, Tem75b]. cylinder [Ceb79, Ken70, LR74, Pat78, Raj73, RL76]. cylindrical [AG72, Bur78, CO77, Cor74, GJA79, Hew74, HP78].

D [BD79, For77b]. damage [Cha73]. damped [BDPH76]. data [Ben76, BNP78, Fos73, KL76, Kov70b, Mac73, SC73, TG77, WGS72, WS73]. Davidson [BK76]. Dawson [GL76]. de-Vries [GM76]. decay [Bun78, PL70]. decaying [Fre73, Vie73]. Decomposing [MS78b]. deconvolution [ER74]. defects [Ama77, Col74]. definite [RLT78, SL72]. deformation [And73]. deformed [AMHH73, ML71]. degraded [JP78]. degree [Ste76a]. degrees [FS73]. delay [RLT78]. delta [HBH78]. densities [AMS73]. density [ASM71, Bey71, Bud79, CT77, LL75, RB77, RG75, Tak76]. dependence [OU78, RR70, HR78]. dependent [FC71, Gil78, GT78a, GK79, HSM74, HPZ78, JJGG78, Lin75b, Mar78, Mud76, Pac74, Pre78, RB72, SGP76, Wat75, Zin73]. Depolarization [Stu78]. derivation [Ros74, Ros76a]. derivative [Gaz75, Joh73, War79]. Derivatives [Tem79a, And78, CM76a, Gaz73]. descent [KB77]. describing [EGV78, Kan76, KW72, Unc74]. design [Kor78, MC72, Sas76b]. designed [LBM^{+}78]. Determination [Dic72, FJ71a, HF71, dAPM79, WCK71, de 77, CS77, JDG^{+}79, KC72, MS70a, Ter79a]. determining [McK71, SK72]. Development [Chu71, GW73]. device [JD70, NG72]. devices [AK72]. DF [PPS74]. Diagnosing [Bun78]. Diagnostic [Cur70]. diagonal [HR74]. diagonalizatio [Pos73]. Diagonalization [Nes65, BH78, Sch70]. diagrams [CH70, JN71, Kal76]. Diatomic [Sch70, Bas72, Bir77, VP75]. dielectric [CW76, KB71, Ros76a]. Difference [Chi70, Ste79a, Ada78, Ant79, BW76, BM70, CC79a, CFU74, CFUW74, CC70a, Chi76, Chi74, Chi75, Chu71, CS74, CR75b, DIC79, DC74, DT77, Dic73, FMW76, For77a, GA72, Gar74, Gar78, God79, GT78a, Gus80, HL75, HH76, HS76, HR74, Hir75, HB77, Ila78, Lau79, Lip77, Mas73, McK71, OJ74, PS70, PW70, Quo76b, Rus70, Sas76b, Sch75a, Sch75b, SS78, Sod78, SGP76, SW75, Ste75, Sun75, Thi78, Tru72, TAG76, Van74, Van77a, Van77b, WH74, Whi70, Wil70, Wil71a, Yee76, Zin73, dR72, vL74, vL79]. differences [Ben76, CG76]. differencing [Gaz76, Hir75, LS76]. Differential [BKC72, All70, Ant76, Ant79, AHMB74, BG70, Bud79, BMP^{+}72, Bus75, Cam71, Cel76, CD78b, CF78b, DG76, Ede73, GH70, HLRP78, HPZ78, Hym77, JSK72, KB79b, KDS79, Le 72, MPB73, PMB73, PH70, RB71, Ros76b, Sco73, SC72, Slo77, SB79, SS79, Ste77a, Van70, Whi79, WH75]. differentiation [SK74b, Tak71]. differentiator [Eva79]. difficult [Dav77]. diffractive [War79]. diffuse [Ryb70]. diffusion [BK78, CS76, Dav79b, Fre73, HL75, KDS79, LS76, Mar75, OA78, PF79, SG79, SS71, Ste71, WS78, WL71, YL72]. Diffusively [Quo76a]. digital [EEJ79, Gaz70, RLT78]. dilute [Aha72, OS70]. dilution [Mos70]. dimension [FB76]. Dimensional [Sto77, ASM71, AB77, BR74, BJ77, BJ79, Bri74, BGL74, BK78, CO77, Cho73, Dav78, DG77, Dic73, EH74, FRSW79, FR77, FJ71a, FO73, FRRT78, GS77, God79, HMR77, HC72, Hoc71, Lar78, Lau79, LN75, LK73, MH79, McD74, MB75, MS77, MS70b, NR71, NC71, NH73, O’B69, PR72, Pra75, Rid79b, RL72, RS72b, RSM70, SK75, SA73, Sch75a, SK78, SG77a, SG77b, SG78, SB75, Sni72, Str78, TTMW77, TTM74, TTM77, Tru72, WCL78, Wei78, Yee70]. dimensions [AG76, Bar70a, Blu71a, CH73, DIC79, HH77, Lad71, Str77, Vre79, WE77a, Wil70, ZHR79]. dipole [KDR73, MC78, SOH72]. Direct [Bir77, GM79b, KH75a, Swe73, Tem79b, WE77a, BLV76, DG76, LM74, O’B69, RDW74, Sch70, SS76, Swa74, Yee70]. direct-numerical-integration [RDW74]. direction [Bel75, Gus71, LK73]. directional [CM76a, NC71]. Dirichlet [KR71]. discharge [Bus75]. Discontinuous [WO75, BJ79, HH78]. discrepancy [BW79]. Discrete [Den74, LG74, Cho78a, HJKS78, ML77, MS77, Mil70b, Nic75, PT73, Sam71, SB73, Swa74, Tem79b]. discrete-variable [Sam71]. Discretization [CB73, Moc76, Sch77]. discretized [Cha74]. discs [Che71c, CW76]. disk [Hoh72, LSM71, LM73, Mil76, Sti71]. disks [Hof74, KB71, ML71]. Dispersion [Chi75, Lin70, Can73, Job78, Mor70a, Rei74b, Ros74, Ros76a]. dispersive [AM79, Sch77]. displacement [Ics70, OU78]. dissipative [MM74]. dissociating [Bir77]. dissociation [Mos70]. Distance [Cri77, Cri78, MPS75]. distribution [AAL79, BK73, CMM^{+}78, Dav79b, DL72, LL75, MN74, OS74, Rid79a, SB73, Sti71, WK79]. distributional [Fos73]. distributions [DS75, Fri71, Jam77, MWW76, Shu73, TV77, Two75, Wil71b]. divergence [Gra76]. diverse [BH72]. domain [Chu71, MPB73, PMB73]. domains [VT77]. dominance [HR74]. Double [Sca71, WO78]. double-minimum [WO78]. doublet [JCM74]. doubly [VT77]. downstream [Hai79, Roa70]. drift [SW79]. driven [Cro70a, Cro70b, DC74]. drop [Foo73]. duct [MC72]. ducts [Bri74]. due [Oku72, Sho78]. DuFort [Lip77]. during [HCH72]. dynamic [Abo78, CS74, Cla77, GS78, GJA79, Sun75, YK71]. Dynamical [JJGG78, Boy74, FS73]. dynamics [Aha72, Bee76, Ben75, DR79, Fin78, Har71, HA71, HAN76, HCB70, KB71, KV76, KL76, Lar78, LR78, Lut71, MGRK73, Mas73, QB73, RCB77, Ves77, ZHR79].

Earth [Dav79b]. economical [GW73]. Eddington [PS70]. eddy [Dea71, HW75]. edge [RL72]. effect [LSM71, Sch77, TR76, ZR73]. effective [Owe71]. Effects [Lan70, AMI75, CLB74, Den74, GW72, GDZK70, RD75, RVD78]. efficiency [Gar74]. Efficient [Ben76, CM76a, KL76, SFG71, CCQ78, CP74, FLRZ71, Kni76, MN74, OS70, Yee72, Yee76]. Ehrenfest [WL71]. EI-Gendi [Ste78b]. eigenfrequencies [LS78]. eigenfrequency [SK72]. Eigenfunctions [BB71, Slo77, Sch70, Sod77]. eigensystems [And78]. eigenvalleys [Sca73]. Eigenvalue [BS70, Ant76, Ant79, BH78, BGP73, Dav77, Dic73, GH70, KS74b, NR79a, Raf79, Run75, Sco73, SB79, SK74b, Tru72, TT78, Was72]. Eigenvalues [Sod77, Bai78, BK76, Dav75, Dic72, Gas78, Har76, KS73, MAT74, Nis72, Rig79, Sha70b, SBPH73, WB77]. Eigenvector [BS70, BH78, ER74]. Eigenvectors [Fel74, BK76, CS77, Dav75, Emp70, HF71, Nis72, Sha70b, SBPH73]. Einstein [Pac74]. elastic [Abo78, Bal76, BMP^{+}72, Chi70, Ila78, KK75, LSM71, OU78, RDW74, Sam71, WO75, Whi74]. elastic-plastic [Sam71]. elasticity [BR74]. electric [FRT79, HL75, SH71b]. electrodynamics [FH74]. Electromagnetic [Mil71a, BL73, Chi71, CW76, God74, Lew72, Ste77b]. electron [Cha76b, Dow70, ED74, Fro73, HL75, HP75, Jos71, KS74a, LNRY73, MD78, Per75, PH78b, RT79a, Smi75a]. electron-atom [LNRY73, Smi75a]. electron-ion [Smi75a]. electronic [Ban73, CFUW74, FJ71a]. electrons [BKM73, SB73, WSB74]. electrostatic [LN75, MC76, Ore73, VT77]. element [ABGR75, BS79, BE79, BJ79, BTS79, Bro79, FGW73, Fle79, FT79, FRRT78, FRT79, HG79, HH78, HD79, HLB79, JB75, Jes74, PB76b, ST76, SB75, THM78, WM71, WO75]. elements [AMHH73, BJ77, Cav75, GFSW78, Gru78, GL77, HRY73, LFMM78]. Ellipsoidal [Cla73]. elliptic [AH73c, CF78b, Det79, GFSW78, HLRP78, Hym77, JSK72, MdVD73, NP74, Nic75, SS79, WM71]. emergent [Shu73]. ended [LBM^{+}78]. energies [BM79, CS73b, DS75, LW78, MC78]. Energy [Lan73, Lew70, AL76, Ben76, CR78, CS73b, HP75, Lie78, MS70a, Ore73, PPP73, PB75, SH71a, SFG71, TV77]. Energy-conserving [Lan73, Lew70]. Engineering [Smi75b]. ensemble [MT78, RNP75, RNP78]. entrance [MC73]. Entropy [EJK^{+}73, AAL79]. equation [AIC76, AYMS71, All70, Bar72, Bea71, BMP73, Blu71a, BGP73, BMP^{+}72, Bus75, CG70, Can71, Can73, CP73a, CGF74, CG77, CFUW74, CK71, CD78b, CK76, Che77, CW76, CS78, CP79b, Cor75, DS74, Dic72, DG76, EB79, EM75, EM77, EJ73, EHMM70, FRSW79, FM73, Fle74, Fle73, FRRT78, Fro73, GS77, Gar76, Gas78, Gat75, Gaz75, GS78, GM76, GS70, GS71, GSD72, GM79b, HZ79, HKN74, Han72, HE78, HH76, HK72, Hil78, Hil79, HB77, Ixa72, Jam77, JKM71, KB79a, KM78, Kus73, Lau71, Lin75b, LoD70, Mar75, MWW76, ML77, McD74, MPB73, ML73, Mor70b, Mor71, O’B69, OA78, PMB73, PH70, Rea70, RB77, Rid79a, Rid79b, RF71, Sch77, SS76, SK77, SK74a, SG76, SG77a, SG77b, SG78, SS78, Ste71, Ste78b, SV78, Str77, SC73, Swa74]. equation [Swe73, Taf76, Tak76, TI73, THM78, Tem79b, TO79, Tra71, WH74, War79, Wat75, Whi70, WE77a, WO78, Yee70, Yee76]. equations [AS72, Ale76, AM79, All70, And74, Ant79, AB77, AH73c, AHMB74, Bar78, BBA73, BKC72, BG70, Bos70, BM77, Bru73, BM70, Cam71, CHS78, CCV72, Cel76, Cha73, CC70a, Chi76, Chi74, Chi75, Cho73, CS73a, CLW78, CF78a, CI74, Coh78, CM74, CF78b, CR75b, Dav78, Dic73, Ede73, EH73, Fro72, Fro78, GCS75a, GCS75b, Gel72, GT78a, Gra76, Gus71, Hat70, HS76, HLRP78, HPZ78, Hym77, Ics70, JB75, JSK72, Kan76, Ker78, KB79b, Kni76, KDS79, KH75a, Le 72, LM74, MdVD73, MGRK73, MS79a, MR73, Mey75, NP74, OJ74, Pac74, PS70, PR72, PB76a, PB76b, Rig79, RB71, Ros76b, RCB77, Sch75a, Sco73, SC72, SW70, SH71b, SBD75, Slo77, SB75, Smi75a, SW79, SS79, Ste77a, TI79, Tho73, Tru72, Unc74, Van70, Vin74]. equations [Wei78, Whi79, WH75, ZHR79, dR72]. equilibria [AK72, JDG^{+}79]. Equilibrium [TR72, AH73b, Ste78a, WCK71]. equipartition [CR75a]. equivalence [Cel76, WH75]. equivalent [HRY73, Mos70]. ergodicity [CF76]. Errata [Ano72e, Ano78g]. Erratum [Ano70c, Ano75e, Ano77e, Ano78h]. error [BR71a, BB76, Fri78, GT74, Ixa72, Tur74]. Errors [CD71a, Chi74, OJ74]. ESR [Pos73]. estimates [Ise73]. estimation [Ben76, TV77, Two75]. Euclidean [PKP79]. Euler [BM79, ZHR79]. Eulerian [Cha75, HAC74, JB75, Mas73, Pra75]. evaluate [CP71, SL72]. evaluating [BEH76, HMB73, RSM70, Win78]. Evaluation [AMHH73, CP79a, ED74, EF77, HP75, HLRP78, KK75, LH72, LN73, Mil74, Nes71, Per75, SAG79, Sin72, Tem75a, AEH73, AL73, BR71a, BO75, CP73b, Cha76b, HSM74, KSV70, KS74a, KH75a, MB71, Mor78, OCS72, Pan75, Pev71, RR70, Sha70a, SZ77, Sha70b, Ste78a, Ste76b, Tem76, Ter79b, Wil71c, MS79b]. event [CP74, FLRZ71, Fri71]. Evolution [Hoh72, AH73c, Gil76]. Ewald [CF77]. Exact [CR78, CF77, GH77, Pac74]. example [CB73]. exchange [LN73]. excitation [Tol75, War78]. excited [HH76]. expanding [JH74]. expansion [BWSE76, CO75, CMM^{+}78, Cor74, HZ79, HRY73, Ics70, KDR73, MC76, Mod75, Owe71, WSB74]. expansions [Cho71b, ER74, HK76]. Experimental [Mil71b, AG75, BNP78, Fos73, Mac73]. Experiments [FJ71b, Han72, Dev74, GW78, JH74, JD74, KV76, Sto78]. explosions [KH75a]. explosives [KH75a]. Exponential [WE77b, BG70, Pat71, SZ77]. Exponential-sum [WE77b]. exponentials [Bos70, TLM73]. exponents [WCK71]. exposed [Mey75, NH73]. exposition [PF74]. extend [Dow70]. Extension [CP75, Hof74, Sal77]. extensions [Hun78]. exterior [KM78]. extracted [WJW78]. Extrapolation [CR75b, Ein76, Lyn76].

F [Cho78a]. face [Bel75]. face-centered [Bel75]. factor [Cho71a]. factors [Ban73, CGN70, Chi71]. Failure [Joh76]. Falkner [CK71]. false [MdVD73]. family [ML71]. Fast [JP78, LM74, Pos73, BT79, HPZ78, Le 72, Tol75, Abr73]. fcc [EJK^{+}73, Mod75, SH71a]. FCT [BB76]. few [Dav75, Kno73, Lav73]. few-body [Lav73]. Feynman [CH70, JN71, Sas76a]. FFT [Roa78]. field [Abr72, AGGM79, Bra76, BKM73, Cor74, FPM76, FT77, FRT79, HH77, JB72, JD70, Kus73, MR73, NTK79, SH71b, TTM74, TO79]. fields [Bun73d, HL75, Lie78, PT75, Ste70, TTM77, VT77, Wei78]. film [Bar73, Blu71a]. filter [Mud76]. filtering [Cam71, Gaz70]. filters [RLT78]. finding [AAL79, Blu71a, Pfe79, Wat71]. Finite [Ant79, Bro79, CC79a, DIC79, FRRT78, Gru78, HH76, HLB79, ST76, SB75, THM78, Tru72, Whi70, Wil70, AMI75, Abr75, Ada78, ABGR75, BS79, BW76, BE79, BJ77, BJ79, BTS79, Cav75, CFU74, CFUW74, CS70, Chi76, Chu71, CS74, CG76, CR75b, Dal77, DC74, DT77, Dic73, FMW76, FGW73, Fle79, FT79, FRT79, GA72, GT78a, GFSW78, HL75, HB77, Ila78, JP78, JB75, Jes74, Lar78, Lau79, Mas73, McK71, Mor79, PB76b, Quo76b, Sas76b, Sch75a, Sch75b, Sod78, SGP76, SW75, Ste75, Sun75, Thi78, Van74, Van77a, WM71, WH74, WO75, WK79, Yee76, Zin73, dR72]. finite-amplitude [CS70]. finite-circuit-element [BTS79]. Finite-difference [Ant79, DIC79, HH76, Ada78, BW76, CFUW74, Chi76, DT77, FMW76, GA72, GT78a, HB77, Sas76b, SW75, Thi78, Van77a, WH74, Yee76, dR72]. Finite-element [Bro79, SB75, FT79, FRT79, Jes74, PB76b, WO75]. finite-particle [Lar78]. finite-sized [AMI75]. firehose [BV76]. first [AIC76, HMB73, HH76, RB77, SH71a, Sca71, Tak76, Wil71a]. first- [SH71a, Wil71a]. first-excited [HH76]. first-order [AIC76]. fitted [Hau79, TTMW77, TTM74, TTM77]. Fitting [Mac73, JMW77, WE77b]. five [Bun71]. five-point [Bun71]. fixed [Cho73]. flame [Mar78]. flames [OB79]. Flow [HA75a, PL75, Pes72, AH70, BD73, Bri74, Bun73d, CKW79, Cho77, Cho78a, Cho73, CH73, CD71b, DS78, Dev74, DE70, FR77, Fle79, FT79, GL78, GM79a, Hai79, Har71, HA71, HA74, HA75b, HE78, Hau79, HAC74, Hof74, HN70, JB75, JPC72, Kus73, LC72, MS78a, MB75, Men74, Mey75, MS77, MC73, NR71, NH71, Pes77, PS72, Raj73, RT76, RL72, RK77, Sch75a, SAK75, SB75, Ste79b, Ste79c, TNM72, THM78, TTMW77, TR72, Van77a, Vre79, Wei78, WJD70]. flows [BBA73, BD79, BE79, Bos70, Cha75, Cha76a, Chu70, Dav73, DG77, DIC79, DC74, FE72, God79, GH73, HC72, Hof73, HLB79, HCH72, KC72, Ken70, KW72, Lap71, MS74, MKS77, MS70b, NC71, NH73, Orl76, Pop78, Pra71, Pra75, Sch75b, She79, SGP76, Tan79, TM70, Tay70, Thi78, TO78, Vie71, WCL78, Wes78]. Fluid [GJ75, AH70, And74, BBA73, BB73, BMP^{+}72, CS74, Dal77, Dav73, DR79, GL78, GP70, GJA79, Har71, HA71, HA75b, HAN76, HMR77, HR74, Hir75, HCB70, HD74, HCH72, LTD79, LR78, Mas73, Mil71a, NR71, Pra75, RT79b, SV76, SK78, Sho78, Ste79c, Sun75, Tay70]. fluid-flow [Har71]. fluid-hammer [SV76]. fluid-immersed [BMP^{+}72]. fluid-transient [SK78]. fluids [HLM^{+}76, Zal79]. Flux [BBH75, BB73, BB76, And75, CC79b, DM73, MB77, Zal79]. Flux-corrected [BBH75, BB73, BB76, CC79b, Zal79]. Fock [CFU74, CM74, FMW76, Fro72, Fro78]. focusing [JD70]. Fokker [CC70a, MKMR77, RB77, Whi70]. Foldy [Ros71]. following [Cla77]. follows [JB75]. force [Bun73d, EH74, Sho78]. forces [BL73, Fin78]. forcing [TR72]. form [Ban73, BW76, CM76a, Chi71, Fos73, Gra76, GH73, HG79, Lin75a, RS76, SG79, Whi73a, Whi73b, Whi74]. formic [MS70a]. formula [Del74, GL77, Han72, Sal77, WSB74]. formulas [Ada78, CF77, KB73, LW78, dAPM77, PCJ76, Str77]. formulation [Bar78, Chi76, Fle79, HLB79, TAE73, Wei78]. formulations [GM70, MB77]. forthcoming [Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q]. FORTRAN [Ada78, Les73, RS72b, Sca71, VP75]. four [BGL74, PF74, Per75]. four-electron [Per75]. four-index [PF74]. Fourier [Abr73, Abr73, Abr75, Ein76, EJ73, FD79, HPZ78, Lad71, Le 72, MB71, Ng74, SS78, Tal78, TLM73]. fourth [FH74, TAG76]. fourth-order [TAG76]. Fox [MS78b]. Frankel [Lip77]. Fréchet [War79]. Free [Lin75b, SH71a, Ben76, BJ77, CC70b, Cro70a, DG77, EC72, FE72, Fre73, HCB70, HD74, JB75, KS74a, LFMM78, MS74, dAPM77, NH71, NH73, PL70, TV77]. free-electron [KS74a]. free-energy [TV77]. free-particle [LFMM78]. Free-space [Lin75b]. free-surface [Cro70a, EC72]. freedom [FS73]. frequency [FC71, Fri71, LO78, PB75]. Friedrichs [Rei74b]. frog [AG76, Wil78]. Fröman [GL77]. fuel [Mey75]. Fully [Zal79]. Function [Tem75a, ASM71, Ama77, BR71a, Bar78, BM79, BK73, BO75, CM76a, Col74, GW73, HBH78, HLB79, JMW77, KS74a, LFMM78, Lon72, dAPM77, RL76, Roa70, Tem76, Voj73, WK79]. function-vorticity [Bar78]. functional [Cha73, Han72]. Functions [SAG79, Tem79a, AMS73, AL73, BR71a, Blo73, Cam72, CJ74, CMM^{+}78, ED74, FGW73, Flo77, Fox77, GL77, HN77, HK76, Jos71, KV76, KP76, LR74, Lyn76, dAPM79, MD78, MS70a, ON73, PL75, PS76, PH78b, RT79a, RL76, Rei74a, RB77, Ros76a, Ryb70, Sca71, Sco79, Tak71, Ter79a, Wal75, War78, Wil71c, WE77b, de 77].

G [Cho78a]. G2M [BK78]. gain [TAE73]. galaxy [Mil76]. Galerkin [AM79, Dav78, DC74, DGGJ75, GH73]. gamma [Ter79a]. gaps [CM73]. gas [Ama77, Bir77, BL75, Bus75, CF76, Cho77, Col74, HP78, Lar78, Lut71, MB75, MT78, OS70, RNP75, SK77]. gas-liquid [RNP75]. Gasdynamic [SC79]. gasdynamics [Vin74]. gases [Aha72, Che71c]. Gaussian [ED74, LFMM78, MD78, PCJ76, PH78b, Ter79b]. GAUSSIAN70 [CP75]. Gendi [Ste78b]. general [AH73a, Chi76, Chu71, Coh76, Fro78, Gil76, HG79, HU74, LG74, Pop78, SZ77, TT78, PKPR72]. generalization [Abr78]. Generalizations [BBH75]. Generalized [Che70, Jos71, Sul72, Ant79, BM77, Cha75, Ste71, TI79, Wes78]. generate [AH73c]. generated [BBA73, Blo76]. generating [AH73a, Bar70b, Kal76]. Generation [BK71, JD74, YLY71, CP74, Eis79, FLRZ71, Fri71, Kno75, Sas76a, SS79, TTM74, TTM77]. generator [Ama77, Bar70a, Col74]. generators [FT77, HR78]. geodesic [Wil71a]. geometrical [PB75]. geometry [AHMB74, Bun73c, CP73a, Cri77, EHMM70, HSM74, MS79b, WGS72, WS73]. geophysical [Dav73]. geostrophic [Yee72]. Gibbs [Chi75]. given [KC72]. Glauber [KK75]. Global [Cri75, HH78, HD79]. Glow [Che70, Squ70, Bus75, Jen78, Pat71, SC72, Sul72]. Godunov [TM70, vL79]. Goldstone [Kal76]. Gordon [SV78]. governing [Wei78]. gradient [KC72, Ker78, SW76]. gradients [Bar72, Bar73]. grand [MT78, RNP75, RNP78]. graphical [HPS74, Pfe79]. graphs [Sas76a]. gravitating [LL75]. Gravitational [Mil70b, AC73, Bra76, JH74, Mil71b]. Green [AL73, KS74a, LFMM78, PL75]. Grid [AMI75, And75, Bea71, CLB74, Dow70, HW75, JJGG78, Lan70, MB77, Nic75, Oku72, SS76, SS79, Swe73, Wil71a]. grids [AH73a, BT79, dR72]. Ground [Emp70, HH76]. Ground-state [Emp70]. group [BK71, Boy74, CZF78]. groups [FJ71a]. growth [Bun78]. Guiding [JKAR77, SK75]. Guiding-center [JKAR77]. gyrating [BTS79].

half [CC70b]. Hamiltonian [Rei74a, Yos73]. hammer [SV76, Les73]. Hard [AYMS71, Bey71, CF76, Che71b, Che71c, KK75, RF71, WK79]. hard-core [Che71c]. harmonic [HD79, MC76, SH71a, PS73]. harmonics [FK77]. Harris [SOH72]. Hartree [CFU74, CM74, FMW76, Fro72, Fro78]. hcp [EJK^{+}73, SH71a]. head [Mor70b, Mor71, MN73, MN74]. heart [Pes72, Pes77]. heat [GH73, Men74, Mor70b, Mor71, MN74, PS72, Smi75b]. heating [Hoc71, Sul72]. helicon [WM72]. helium [HH76]. helix [JD70]. Helmholtz [Yee70, Bun73c, Hil78, Hil79, O’B69, Rid79a, Rid79b]. Hermite [AMS73, Cho71b, EJ73, MH79]. Hermitian [CS77, FB76, Nes71, Thi78]. heterogeneous [SK77]. heuristic [GP70, TR76]. hexagons [Che71b]. High [Bey71, CCQ78, And75, BNGF73, HGE74, Lie78, Raj73, Ste75]. high-energy [Lie78]. high-order [BNGF73]. High-resolution [CCQ78, HGE74]. Higher [For77a, Hir75, Ros76a]. highest [SBPH73]. Highly [Ada77, CS76, DC74, HN77]. Hilbert [AMS73, Sod77]. histogram [Fos73]. homing [BK76]. Homogeneous [Eas72, JB72, OU78, RT76]. Hopscotch [MN73, GM76]. hot [Ros76a]. Householder [MS78b]. Houston [Sin72]. Hybrid [BCCH78, Gru78, HZ72, HCH72, She79]. Hydrodynamic [GDZK70, Bar70a, PT75, SK74b]. hydrodynamics [Chr73, EC72, FB79, SA73, SAHK75, Str78, Zin73]. hydrogen [Bur78, FRT79, NRH71, PPP76, TO79]. hydrogenic [Bur78, Jos71, Rei74a]. hyperbolic [AG72, BW76, BM70, Chi74, Chi75, Cho76, CP79b, Gar76, Gar78, GT74, GM70, Gus80, Hed79, HS76, Orl76, OJ74, Rus70, Sod78, Tur74]. hyperosculatory [Sal75, Sal76]. hypersonic [JPC72, NC71, RL72]. hypersurfaces [YLY71].

I-complex [KS73]. IBM [Ama77, Col74]. ICE [Wes78]. ideal [Ama77, CGG^{+}78, Col74, Gru78, Van77a]. idempotent [HF71]. Identifying [Sto78]. II [BBH75, CJ74, DM73, EM77, GT74, LN73, MS79b, MS70a, Mor71, PMB73, RL76, Ros76a, WS73, vL74]. III [BB76, HU74, Van77a]. ILUCG [PKP79]. imbedding [AW74, AKK70, SW79]. immersed [BMP^{+}72]. implicit [Ada77, BW76, BP73, BM77, CLW78, DMWV75, FC71, God79, Gus71, HR74, HB77, Jac73, LR78, LK73, MP77, Mas73, MM74, NC71, PB76b, SA73, SAHK75]. implicitization [Gra73, Gra76]. Improved [NH71, SAHK75, BW79, FRSW79, Fro72, HBH78, SG79]. improvement [FH75]. Improvements [GSD72]. impulsively [Ceb79, PL75]. impurity [OA78]. inadequate [SC73]. including [Lew72, Tol75]. inclusive [EF77]. incomplete [Ker78, Ter79a]. incompressible [AH70, BE79, Cha75, Cha76a, DIC79, FT79, HCB70, HLB79, NR71, NH71, SB75, Vie71]. incompressible-flow [NH71]. incorporating [Bar78]. incorporation [GH77]. increasing [Joh76]. indentation [Abo78]. independent [Bud79]. index [Ano70a, Ano70b, Ano71a, Ano71b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano74e, Ano74c, Ano74d, Ano75b, Ano75c, Ano75d, Ano76b, Ano76c, Ano76d, Ano77b, Ano77c, Ano77d, Ano78c, Ano78d, Ano78e, Ano78f, Ano79b, Ano79c, Ano79d, Ano79e, PF74]. indirect [Two75]. induced [MC78, Quo76a]. inelastic [Ale76]. inequalities [Les73]. Inertial [Lip77]. infinite [AEH73, BEH76, Hof74, Mos70, Win78]. inflection [dAPM79]. Information [Sha72]. inherent [WD78]. inhomogeneous [AW74, CG76, FM73, Hew74, Mil71a, Rei73]. Initial [EC72, Ste77b, Wil71b, BS79, CGF74, Chi75, CN75, Dav73, GH77, KR71, KS73, Moc76, NR79a, Sas76b, Sco73, WD78, Wat71]. initial-boundary [Dav73]. initial-value [BS79]. instabilities [CGG^{+}78, DK77, God74, God75, JJGG78, Oku72, RT79b, WD78]. instability [BV76, BGL74, Kan76, Sho78]. insulating [CM73]. integer [Sca71]. Integral [Ben72, Che70, CW76, CP79a, Squ70, Bos70, Bru73, Cam72, CJ74, Cha73, CP73b, CI74, Coh78, DS74, FRSW79, FH75, GL77, HE78, HK72, Ics70, Jen78, KS73, Kni76, Kus73, Mod75, Mor78, Pat71, Per75, PB76a, Sas76b, SH71b, SBD75, Ste77a, Ste78b, Sul72, Tak76, TI79, War79, Whi79]. Integrals [Bot70, Rei74a, Abr73, AEH73, AL73, BR71a, BEH76, BO75, Cha76b, ED74, FJ71b, FH74, HMB73, HP75, Ise73, Jos71, LN73, MD78, MS79b, Mil74, Mor70a, OCS72, Pan75, PCJ76, PH78b, RR70, RSM70, SL72, Sha70a, SZ77, Sin72, Ter79b, TLM73, Win78]. integrand [Lyn76]. integrands [Ise73, Pan75]. Integration [Bea71, BNP78, HN77, Str77, AC73, BW79, Cha76a, Cha76b, CK76, Che77, Gol72, Hol77, JKM71, KD76, Lan79, Lep78, OS74, Pac74, PH70, RDW74, Ros76b, RCB77, SK74a, SW70, TI73, Ter79a, Van74, WO78]. integrations [Mud76]. integro [BMP^{+}72, CD78b, MPB73, PMB73, PH70]. integro-differential [BMP^{+}72, CD78b, MPB73, PMB73]. integro-partial [PH70]. Interaction [EM77, LBM^{+}78, Ben79, MC76, MC78, MS70a, PH78a, PL70, Yos73]. interaction-free [PL70]. Interactions [LM73, Lew72, Owe71, SH71a]. interactive [Whi79]. interface [HU74]. interfaces [Cha75, JB75, PF79, RT76]. interferometer [Cum75]. internal [NR71]. internuclear [MPS75]. interpenetrating [HA75a]. interpolation [HM75, RK77, Sal75, Sal76, Sha72]. interpreting [HPS74]. interval [Cho73, Mil70a]. intervals [Bol78]. Intrinsic [Abr72]. invariant [AW74, AKK70, CM76a, Mor78, Sas76b, SW79]. invariants [BNGF73]. Inverse [Sal76, KC72, KB77, LC72, Pex76, SW75]. inverse-matrix [SW75]. Inversion [Bun73c, Bun73b, Kov70a, Bun71, Bun73a, DM79, DW78, Hym77, Lon72, Ste70, Two75, Voj73]. inverted [RB72]. investigation [Kan76, Rei74b]. inviscid [And74, Fle79, MB75, TNM72]. involved [Smi75a]. involving [Bot70, Cel76, PH78b, SL72, WSW78, WH75]. ion [Smi75a, WJW78, WSW78]. ionization [OA78, Per78, Pre78]. ionized [SK77, Unc74]. irradiations [Cla73]. Irregular [Hat70]. Ising [BKL75]. isolated [EB79, Jam77, MWW76]. isothermal [MY71]. issued [PL70]. Iteration [Ste70, Cor75, RS74]. iterations [CGN70]. Iterative [BS70, Fro72, Run75, And78, BGP73, Dav75, Ker78, Mil71c, Sha70b, SBPH73, TT78, Two75]. IV [Van77b].

J [Saf78]. Jacobian [MAT74]. jet [Bos70, EF77, KW72]. JWKB [Kau73].

Kalman [Cam71]. Kármán [Pat78]. kernels [Coh78, TI79]. kidney [MS79a]. Kind [Tem75a, Sca71, Tak76, Tem76]. Kinematically [PT75]. kinetic [Gel72, Ros77, Ste71, Wat75]. kinetics [Ede73, YK71]. Klein [SV78, WSB74]. known [JMW77]. Korteweg [CG77, GM76]. Krook [Wat75]. Kullback [JMW77].

Lagrangian [Bar70a, BP73, Cha75, For73, FB79, HP76, HCB70, HAC74, HJKS78, KB79a, Pra75, Rit70]. lake [Tak71]. laminar [Ceb79, Mar78, OB79, PS72]. Landau [Can73]. Langdon [GL76]. Langmuir [RVD78]. Laplace [Voj73, AG75, Bun73c, DM79, DW78, FM73, HPZ78, Lau71, Lon72, Pex76, Rea70, RB72, Sal75, Sal76, SS78]. Large [Fel74, Nes65, BD79, BT79, BK76, Cho78a, CS77, Dav75, dAPM79, MAT74, MS77, Nes71, Nis72, Raf79, Sha70b, SBPH73, TT78, Yos73]. large-order [dAPM79]. large-scale [TT78]. larger [de 77]. laser [Cla73, Gol72, PPS74, Sho78, TAE73]. laser-plasma [Sho78]. lateral [HW75]. lattice [AL73, Boy74, Che71b, Che71c, CF77, KS74a, ML71, Mod75, MT78, Pev71, Stu78]. lattices [EJK^{+}73, HMB73, JD74, LH72, SH71a]. law [BW76, EH74]. laws [MM73, MM74, Sod78]. Lax [GM70, Hun78]. layer [Bos70, BL72, CK71, Ceb79, CKW79, KC72, LC72, MS78a, PB76b, Thi78, Vre79]. layered [LS78, Sam71]. layers [Cho78b, Hai79, Quo76a, Ste70]. layout [MC72]. leading [RL72]. leap [AG76, Wil78]. leap-frog [AG76]. Least [CP76, JPC72, BH78, BDPH76, Cur70, Dow70]. Least-squares [JPC72, BH78]. Legendre [Del74, Ics70, WSB74]. length [HR78]. lengths [CS76]. level [McK71]. Lewis [Bun73b]. light [CHS78, Mil71a, YK71]. like [AL73, ML73]. limit [AYMS71]. limitations [Kov70b]. limited [CD78a]. limiting [War78]. limits [AEH73]. line [WCK71]. Linear [Sch75a, AM79, AW74, Ant76, CCV72, HPZ78, JD70, Ker78, KB79b, Les73, NP74, NR79b, Rig79, Ros77, Sha72, Slo77, Smi75b, Tay70, Two75, Vie73]. lines [Gar76, HS76, JSK72]. linked [CB71]. Liouville [Bai78, Har76]. liquid [Foo73, Lad71, LL75, RNP75, WCK71]. liquids [Blu71b, ME72, SS71]. liquids-approximate [SS71]. LISP [CJ74]. List [Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q]. lobe [ED74]. Local [Cav75, Chi74, GFSW78, OJ74, DE70, Hym77, KB79a]. locally [WCL78]. located [Bel75]. location [HU74, Lav73]. log [Joh73]. log-derivative [Joh73]. logarithmic [PCJ76, Tal78]. Long [Bal76, Fin78, RVD78, RNP78, EM75, EM77]. Long-range [Fin78, RNP78]. Long-time-scale [RVD78]. long-wave [EM75, EM77]. Lorentz [Sho78]. loss [Dav79b, Sha72]. lossy [CW76]. Low [JD70, BW79, DG77, FE72, GJ73, GJ75, HP75, LO78, Mod75, PS72, RT76, She79]. low-discrepancy [BW79]. low-energy [HP75]. low-frequency [LO78]. low-pressure [GJ73, GJ75]. low-speed [RT76]. low-temperature [Mod75]. Lowest [BS70, Dav75, SBPH73]. lying [dAPM79].

MAC [AH70, Dev74, Eas72]. MacCormack [DT77, GT78b, Jon79]. Mach [FE72, HA74]. Machine [Chu71]. Madelung [Bel75]. magnetic [Abr72, CR78, Cor74, DM73, FT77, JD70, Lie78, Lut71, MR73, NTK79, SH71b, TO79]. magnetic-field [JD70]. magnetized [BCCH78, Che77]. magnetohydrodynamic [CGG^{+}78, DGGJ75, DK77, Gru78, JDG^{+}79, LTD79, LK73]. magnetohydrodynamics [ABGR75, BP73, Lin75a, Lin77]. magnetostatic [BNPLD77]. magnitude [CM76a, Dea71]. manipulation [Han72]. many [Emp70, Sco79]. many-body [Sco79]. mapping [CP74]. mappings [VT77]. marching [FR77]. Marker [For73]. Markov [Ros77]. Mass [Ben75, AL76]. master [LoD70]. matching [GH77]. material [HA75a]. materials [WO75]. Mathematical [Bus75, PB76a]. mathematics [Sto77]. Mathieu [Can71, EH73, HK76]. Matrices [Fel74, Nes65, BH78, BK76, CZ79, Cri78, Dav75, FB76, HF71, NR79a, NR79b, Nis72, Ore73, RG75, Sha70b, SBPH73, VW77]. Matrix [BS70, EH73, LNRY73, AMHH73, BG70, Blu71a, Cam71, CS77, GH70, GL77, HG79, HRY73, Kov70a, LFMM78, MAT74, MN74, Nes71, Pos73, Raf79, Ros77, SW75, Yos73]. maximal [AAL79]. Maxwell [CM76a]. Maxwellian [SB73]. MCHF [Fro73]. means [Blo73, FJ71a, Mey75, SW79]. measured [ASM71]. Measurements [Hoc71, WCK71]. mechanical [Ben72, SK72]. mechanics [And71, BBA73, CM76b, FPM76, HR74, Hir75, Hub71, HK76, LG74, PS76, VW77]. media [Bal76, GJY75]. medium [CG76]. Meeting [Ano76e]. Mellin [DW78]. Memorial [BL78]. Mesh [CS73a, Hym77, Bar70a, BR71b, Cav75, iCS78, CR78, CD71a, EH74, Eas75, FB79, GFSW78, HP76, Pra75, SS79, Van70]. mesh-point [SS79]. meshes [Bar70b, HJKS78]. mesic [PPP73]. meteorology [Rit70]. Method [AF79, TG77, Abr78, AIC76, AG75, And78, And73, AH73b, Ant79, AB77, BH72, Bar70b, Ben79, BH78, BK73, BJ79, BBH75, Bri74, BM77, BL72, Bud79, BK76, CG70, Cha75, Cha76b, CN75, Chi70, Chr73, CW76, CLW78, CH73, CP79b, CMM^{+}78, CF77, Dav79a, Dav78, DS78, DL72, DGGJ75, Dic73, DG76, DR79, Eas72, Eis79, EHMM70, EJK^{+}73, FRSW79, FR77, Fle74, FH75, Foo73, FE72, For73, Fos73, Fre73, FLRZ71, FRRT78, FRT79, GH70, Gar76, Gaz75, Gil76, GM70, Gra76, GM76, GSD72, GH73, Gus71, HU74, HA71, HH78, HD79, HS76, HCB70, HAC74, Hof73, HN70, Hud72, HB77, Hun78, Ixa72, JCM74, JB75, JJGG78, Jes74, Joh73, JSK72, KS73, Kau72, Ker78, KW72, KB73, KDR73, Lau79, Lau71, LR78, LNRY73, MdVD73, MBM76]. method [MGRK73, MC72, Mas73, MY73, Men74, Mil70a, MPB73, NR71, NR79a, NR79b, OA78, Pat78, Pes72, Pfe79, PPP73, PPP76, Pra71, QB73, Rea70, RL72, Rit70, Ros77, SK75, SK72, Sca73, SE76, SS76, Sco73, SK77, SZ77, SV76, SK78, Shu73, Sin72, SB79, SOH72, SK74b, Ste78a, SW75, SB73, Ste75, Ste78b, Tak76, TI73, TM70, TI79, Thi78, Tol75, TR76, Tru72, TT78, Van74, Van73, Vie71, WCL78, Was72, Wat75, Wes78, Yee76, vL79]. Methods [RSM70, Ada77, AM79, Bee76, Bel70, BE76, BEH76, Bro79, BDPH76, BM70, CCV72, CK71, Cho77, CS73b, CM76b, DM79, DIC79, DC74, EM75, FPM76, FM78, FT79, FT77, Gaz76, Gil72, GT74, GT78a, HMB73, HE78, HR74, HLRP78, JKM71, KK75, Kus73, LSW72, ME72, MM73, MKS77, MN73, PS73, Ros76b, RK77, SG76, Smi75b, Sod78, Ste76a, Ste70, Swe73, Tan79, TNM72, Tem79b, TO78, Tur74, Van70, WM71, WD78, WH74, Whi79, Whi70, Win78]. MeV [DGH73]. microinstabilities [LO78]. Milinazzo [Cho78a]. minicomputer [HLM^{+}76]. minicomputers [LBM^{+}78]. Minimal [BB76]. Minimal-error [BB76]. Minimization [CS73b, BH72]. minimizes [PKP79]. minimum [AK72, RLT78, WO78]. minimum- [AK72]. minimum-delay [RLT78]. mirror [AK72, BKM73]. Mixed [FT79, And71, HD79, Quo76b]. mixing [CF76, EEJ79]. mixture [BL75]. MO [JE74]. Modal [BMP73]. mode [ABGR75]. model [AM79, AGGM79, BL75, Bus75, CT77, Cla77, Cro70a, Cro70b, CLS78, Dal77, DD79, DK77, DGH73, EF77, GW78, GP70, HL75, HN78, JN71, Jon79, LO78, MP77, Mar75, MS78a, MKMR77, Mud76, Quo76b, Rei74b, Sch75b, SS71, TA77, War79, WJD70, WL71, Yee72]. modeling [BTS79, Gat75, Gol72, SC79]. models [BGL74, CO77, CS78, EH74, FLRZ71, HW75, HGE74, LN75, MB77, ST76, Ste79c, THM78]. modes [Che73, Job78, OU78]. Modification [BK76, Sha70b, TI73, Kau72]. Modified [Cho71a, Tem75a, Bai78, BO75, CF78a, Sca71, WH74]. Molecular [KB71, Aha72, Bee76, Ben72, Ben75, Bir77, Bru74, Fin78, HLM^{+}76, KV76, KL76, QB73, RR70, RCB77]. molecule [Sch70]. molecules [Bas72, Blu71b, PPP73, PPP76, Ves77, VP75]. moment [KB79a]. momenta [God75]. moments [AG75]. momentum [Abr72, AS72, BB71, CD78a, MBM76]. momentum-space [AS72]. monoenergetic [AHMB74]. monotone [Cor75]. monotonic [For77a]. Monotonicity [vL74]. monovacancies [LSM71]. Monte [Ben76, Bir71, BL75, BKL75, BW79, CD78a, CHS78, Che71b, Cho71a, Cho71b, CM76b, FC71, FLRZ71, Gil78, GS71, HLM^{+}76, HP78, KK75, Kno75, MH79, Mat70, MT78, RR70, RNP75, RNP78, TO79, TV77, VW77, WK79, WL71]. motion [And74, JB72, MR73, RCB77]. motions [Sam71]. moving [And75, Bar73, Vie71]. multi [Eis79, HMR77, MH79, McK71]. multi-dimensional [MH79]. multi-fluid [HMR77]. multi-level [McK71]. multi-surface [Eis79]. multibody [Hau79]. multichannel [Joh73]. multiconfiguration [FMW76]. Multidimensional [TAG76, BWSE76, BW79, BM77, HN78, HL76, Ise73, Lep78, Lyn76, OB79, Smi75b, Zal79]. Multifluid [HA74, And75]. multifrequency [PS70]. multinephron [MS79a]. multinomials [CP76]. multiparameter [CLS78]. multiperipheral [FLRZ71]. multiphase [HA75b]. multiple [Chi70, HPS74, Nic75]. multiple-scattering [HPS74]. multiplying [CGN70]. multipole [CM76a, CO75, MC76, MC78, Owe71]. multipole-expansion [CO75]. multistep [Bee76, God79, GT78a, GM70, SG77a, Van70].

natural [Cla73, GJ75]. Navier [AB77, Bar78, BM77, Chi76, GCS75a, GCS75b, HB77, PR72, SB75]. near [SS79, War79]. nearcharacteristics [SK78]. nearly [Blu71b]. neighbor [SH71a]. neighborhood [WCK71]. neighbors [QB73]. Nesbet [Nis72, Sha70b, SOH72]. network [Blu71a, MKS77]. NEUCG [Bur78]. Neumann [FM73, SS76]. neutral [Ban73, HP78]. neutron [CP73a]. neutrons [LW78]. Newton [CFU74, CFUW74, FMW76, PPP73, PPP76]. Newtonian [GL78]. Nicolson [CCV72]. nine [Bun73a]. nine-point [Bun73a]. Nishina [WSB74]. no [Bar78, CM73, Tay70]. no-slip [Bar78, Tay70]. node [DL72]. noise [Ste77b]. noises [Oku72]. Non [BS70, AM79, GL78, GJA79, HL75, KB79b, Roa78]. non-axisymmetric [GJA79]. non-linear [AM79, KB79b]. non-Newtonian [GL78]. non-periodic [Roa78]. non-self-adjoint [HL75]. Non-Symmetric [BS70]. noncoupled [GM79b]. nondivergent [Cro70b]. nonequilibrium [CH73]. nonfunctional [Fos73]. nonhomogeneous [LM74]. Nonlinear [BMP73, CLS78, RZ75, AH73b, BKC72, CD78b, CN75, CP79b, DM73, DC74, EJ73, FC71, Fle74, GM70, HW75, Hed79, LoD70, MC72, Moc76, PH70, Rus70, SE76, SW70, Sod78, SV78, Sul72, SC73, Two75, War79, ZR73]. nonlinearly [WO75]. nonlocal [GS70, Ste71, Tho73]. nonorthogonal [KP76]. Nonphysical [BL73, Oku72, TV77]. Nonreflecting [Hed79, Smi74]. nonrelativistic [HH77]. nonspherical [Rei73]. nonstationary [WSB74]. nonsteady [Tay70]. Nonuniform [BR71b, CR78, CD71a, dR72]. nonunique [McD74]. norm [PKP79]. normal [ABGR75, Bar73, Job78, WL71]. normalized [BR71a]. Note [Cho78a, Cri78, Dav75, HBH78, Sal75, SB79, Sun75, Win78, AG76, Fri78, Ryb70, Wil78]. Notes [Chi74, Voj73]. notice [Ano76e]. novel [Cri77]. nuclear [AMHH73, HH78, OFM71, Tol75]. nuclei [Owe71]. nucleon [BLV76, Fle73, GW72]. nucleon-nucleon [Fle73, GW72]. number [Ama77, CP75, iCS78, Cho78a, Col74, DG77, FE72, HR74, HR78, MS77, PS72, TTM74, TTM77, Wil78]. numbers [Blo76, HA74]. Numerical [AG72, AS72, BR71a, Bar70b, Bel70, BJ77, BO75, BD73, Bra76, BKM73, Bri74, Can71, Can73, Che73, Chr73, CR75a, CI74, CP79b, Cor74, DM73, DM79, DS74, Den72, Dev74, DGGJ75, DW78, EM75, EM77, EGZ72, EGV78, FPW73, FPM76, FT77, FS73, Fro78, GCS75a, Gaz73, Gaz75, GS78, God74, Gol72, GO77, GL77, HMB73, Har76, HA75b, Hew74, HSM74, Hil78, Hil79, Hol77, HCH72, JB75, JH74, JDG^{+}79, JKM71, KSV70, KC72, KS74a, KD76, KW72, Kor78, KM78, Lad71, LWH78, Lon72, ME72, MS74, Mar75, McD74, MS79a, MS70b, MC73, NR71, NP74, ON73, Pac74, Pes77, PB76a, PPS74, PH70, RB77, Rei74b, Rid79b, RLT78, RCB77, SAK75, SC72, SV76, Sho78, SK74a, SG76, SG77b, Ste76b, SV78, SC73]. Numerical [Taf76, Tak71, Tak76, Tal78, Tem75a, TTMW77, Tho73, TAE73, Unc74, Voj73, War78, War79, WM72, Abo78, AIC76, AC73, All70, AGGM79, And74, And73, AB77, Bal76, BLV76, BR74, BD79, BE79, BK73, Bru74, Bru73, Bus75, BNPLD77, Cam71, CGF74, Cel76, CGG^{+}78, Cha76b, CD78b, CT77, Coh78, CF78b, Cro70a, Cro70b, Dal77, Dav77, DK77, DGH73, DR79, Ein76, ER74, Eva79, FRSW79, Foo73, For77b, GS77, Gat75, God75, GS70, GSD72, HKN74, HW75, Har71, HA71, Hau79, Hof73, HN70, HLRP78, HPZ78, Ise73, Job78, Joh76, JB72, JSK72, Kan76, Kau72, Kau73, Lav73, Lew70, Lew72, LoD70, LM74, Lut71, MP77, MB71, MGRK73, MWW76, MKS77, Mil71b, MPB73, Mil74, ML73, Mor70a, Mor70b, Mor71, Mor78, Ng74, NR79b, NH71, O’B69, OB79, PR72]. numerical [Pes72, PCJ76, Pra71, Quo76a, RT76, RDW74, RL72, RZ75, Ros76b, SK77, SG77a, SG78, SW70, SBD75, Smi75a, SW79, Ste78a, Ste77a, TM70, Tay70, TNM72, Tem76, TI79, Thi78, TTM74, TTM77, Tra71, TR76, Van74, Van77b, Vre79, WCL78, Wat75, Wil71c, WJD70, WH75, WO78, Yee70, Yee72, ZG73, PMB73]. numerically [CCV72, Gil76, RB72]. Numerov [Fri78].

objective [JMW77]. objects [JP78]. observation [Tak71]. obstacle [NR71]. obtaining [MC76]. ocean [CC71, CT77, Cro70a, Cro70b, HW75]. oceans [MP77]. odd [SB79]. Oldroyd [Gat75]. Oldroyd-type [Gat75]. Olver [SL72]. One [Bot70, MD78, BJ79, BGL74, Cho73, Dav78, Dic73, GS77, KSV70, Lad71, MS70b, O’B69, RT79a, THM78, Tru72, Yee70]. One- [MD78]. One-center [Bot70]. one-dimensional [BJ79, BGL74, Dav78, Dic73, GS77, MS70b, O’B69, Tru72, Yee70]. one-electron [RT79a]. one-equation [THM78]. one-speed [KSV70]. open [LBM^{+}78]. operator [Bun71, Bun73a, Bun73c, Bun73b, BB71, CLW78, Sod77]. operators [KS73]. Optical [Net71, DGH73]. optical-model [DGH73]. Optics [WJW78, Fle74, Hil78, Hil79]. Optimal [Eas75, Bar70a, Cel76, DL72, Gar74, WH75]. optimisation [JD70]. optimization [Cri75, PKPR72, Rig79]. Optimized [Bas72, DMWV75, BH78, GM70]. Optimum [Kau73, Kau72, Kov70b]. orbital [BB71]. orbitals [CP75]. order [AIC76, BNGF73, BM79, BJ79, BGP73, BM70, Cha76a, EGZ72, For77a, FH74, GW72, Gol72, GT74, Got74, Hir75, dAPM79, MM73, MM74, MC76, PR72, Ros76a, Rus70, Sca71, Sco79, SB79, Ste79a, Tur74, TAG76, Wil71a, de 77, vL74, vL79]. Ordinary [Tem79a, Ant79, DG76, GH70, Ros76b, Sca71, Sco73, Slo77, Tem76, Van70]. origin [dAPM79]. Orr [Gas78]. orthogonal [Bar70b, FB76, GW73, HB77, Mey75, Pop78, PT73]. oscillation [Foo73]. oscillations [LS78, Lip77, MY71]. Oscillator [PS73]. oscillators [BMP73]. oscillatory [AEH73, BEH76, BO75, Bun78, HN77, Pan75, Tra71, Win78]. osculatory [Sal75, Sal76]. other [BE76, Jon73]. Over-relaxation [DT77, BE76]. overlaps [WS73]. overrelaxation [Rig79]. overtones [LS78].

P [Cho78a]. packed [AYMS71, Bey71]. packings [ML71]. Padé [FPW73]. Pair [Sti71]. paper [Cho78a, Sun75]. parabolic [Bea71, Bos70, CCV72, CS73a, CLW78, CF78b, LR74, RL76]. Parallel [CP73a, CK71, HL75, Mey75, MC73, Raj73, RF71]. parameter [SG79]. parameters [TI73]. parametric [RT79b]. part [Chu71]. Partial [Fle73, Gra73, Mil71c, BKC72, CF78b, CS73b, Gra76, HLRP78, HPZ78, Hym77, JSK72, KB79b, KDS79, Le 72, MY73, PH78a, PH70, RB71, SS79, Ste76b, Ste77a]. partial-wave [Ste76b, Ste77a]. partially [JMW77, Unc74]. particle [AMI75, Abr72, AL76, BTS79, BNPLD77, CF76, CS76, Den74, EH74, Eas75, For73, God74, God75, HAN76, JB72, Lar78, LTD79, LFMM78, LSW72, LW78, MR73, Mor78, RVD78, Ste77b, TA77, Two75, Ves77, Wil71b]. particle-in-cell [HAN76, LSW72]. particle-mesh [EH74, Eas75]. particles [Bey71, Dav79b, Kno73, Lew72, PL70, Sto78]. particularities [Job78]. particulate [GS78]. partition [HE78]. passage [RB77]. past [CC71, HN70, Ken70, PL75, TM70]. path [FJ71b]. pattern [LS78]. patterns [Pes72]. PDE [EGV78, Gar74]. peaked [AS72]. Peclet [PS72]. penalty [HLB79]. penetrating [MR71]. penetration [DM73]. Performance [CJ74]. Periodic [Van73, Bal76, BGP73, Dal77, Den72, EH73, Roa78]. periodically [GJY75]. peripheral [BE76]. permanent [MC78]. perturbation [Hof74, Mor79, PH78a, RDW74, RS74, Sni72]. perturbations [CS78, WB77]. perturbative [AIC76]. perveance [WJW78]. Petit [MT78]. Phase [GT74, Tur74, Cam72, CJ74, CD78a, Den72, EF77, GL77, Kau72, Kau73, LR78, Mil70b, Mor78, RT76, RDW74, Ste79c, Whi79]. phase-integral [Cam72, CJ74, GL77]. phase-space [EF77]. phases [And71, HA75a, dAPM77]. phenomena [BCCH78, HPS74]. phenomenon [Chi75]. photoelectric [CCQ78]. photoneutron [MY73]. physical [Bol78]. physics [AG75, Le 72, Pfe79, PB76a, SL72]. picture [Gaz70]. piecewise [Bru73, Smi75b]. piezoelectric [CC79a]. pinch [Ste78a]. pipe [CD71b]. Pitch [Dav79b]. Planck [CC70a, MKMR77, RB77, Whi70]. plane [GH73, Ila78, Sch75b, Smi74, VT77]. planes [CC70b]. planetary [Bea71]. Plasma [Kno73, KS74b, AMI75, AL76, BL73, BGL74, BCCH78, CO75, Che77, Cor74, Cum75, DD79, Fre73, GJ73, GJ75, Gru78, Hew74, HN78, Hoc71, HGE74, KDR73, Lan73, Lan79, LSW72, Lew72, LN75, MKMR77, NTK79, Oku72, Pfe79, Ros74, SL72, Sho78, Tak76, TA77, Unc74, WJW78, Wil71b, WJD70]. plasma-simulation [BL73]. plasmas [And75, Bur78, BNPLD77, CLB74, CO77, Den72, Gol72, HP78, JKAR77, Lan70, LTD79, Lew70, Lin70, Per78, RD75, Ros76a, RT79b, SK75]. plastic [Sam71]. plastically [ML71]. plastically-deformed [ML71]. plate [Abo78]. plates [MC73]. plot [RS72b]. Point [Det79, AW74, Ant76, Bun71, Bun73a, CF76, CN75, Chr73, DL72, Gru79, Jon73, KH75b, Lew72, dAPM79, MAT74, NR79b, RS72a, SS79, Tro76]. points [KH75a]. Poiseuille [CD71b, GH73]. Poisson [Bun71, Bun73a, Bun73c, Bun73b, BT79, EB79, HZ79, Jam77, MWW76, ML77, McD74, SS76, Swa74, Swe73, Tem79b, WE77a, Yee76]. Poisson-solver [BT79]. polarizable [Ves77]. polarized [CHS78]. poles [Coh78, MC76]. policy [Cel76]. polyatomic [BL75]. polygonal [Rid79b, VT77]. polyhedra [BDF78, Fin79]. polymer [FD79, SFG71]. Polynomial [RK77, HBH78, RB77, WSB74]. polynomials [Bru73, HZ79, KD76, MH79, PB76a, SL72, Sal75, Sal76, Wat71]. polypeptide [CS73b, Cri75]. polytope [RSM70]. polytropes [HU74]. population [Pat71]. positive [MAT74, RLT78]. positive-definite [RLT78]. positive-real [RLT78]. positron [HP75]. positron-atom [HP75]. postignition [HCH72]. Postscript [Whi73b]. potential [AMHH73, AB77, Bar73, BWSE76, Dic72, FRSW79, GS70, HE78, HH78, Lau79, SOH72, TTMW77, WO78]. potentials [AS72, ML73, Tho73]. power [Blo76, EJ73]. practical [CC70a]. practitioner [Gel72]. precipitating [Dav79b]. precision [Sca71]. predicting [Bri74]. predictor [Hud72]. predictor-corrector [Hud72]. preferred [Fri71]. premixed [Mar78]. prescribed [OS70]. presence [Dav79b]. pressure [Eas72, GJ73, GJ75, KC72, SW76]. primitive [Dav78, Fle79]. principal [CP73b, Mor70a]. principle [Bar78]. principles [PKPR72]. print [Mor70b, Mor71, MN73, MN74]. print-head [Mor70b]. probability [Gil78, JMW77, MS79b, RB77]. probed [YK71]. problem [AG72, Abo78, AC73, AGGM79, Ant79, BWSE76, BJ77, BJ79, BL72, CFU74, CN75, Dav73, DE70, FM73, FMW76, HD79, HSM74, Ila78, Jon79, JPC72, KR71, KSV70, KC72, KS74b, KH75b, LWH78, MPS75, Mil71b, Moc76, MN73, OFM71, PS76, RS72a, RS76, Sco73, Shu73, Tro76, TT78, Wat71, WM72]. problems [AW74, AH73b, Ant76, BS79, Bar72, BLV76, BR74, Bel70, Boy74, Bru74, Cel76, CP73b, Chi75, Dav77, DL72, Det79, FRSW79, FPM76, FM78, FT79, FB79, GA72, GH70, GH77, GT74, GFSW78, GJA79, GO77, GK79, Hai79, HE78, Hau79, HR74, Hir75, Jac73, Jon73, KM78, KB77, LS76, Lau79, MAT74, NG72, NR79a, NR79b, Pfe79, Raf79, Rit70, Roa78, Sas76b, Sch75a, SH71b, SB79, Smi74, Smi75a, SK74b, TNM72, Tur74, WM71, Was72, WS78, Wil70, WH75, YL72, Zin73]. Procedure [BS70, BNGF73, BK76, Fin79, Fro73, Jac73, LC72, MC76, PPS74]. process [Ste71, Wei78]. processes [Bir71, Cor75, EEJ79, KDS79, Lie78, Mat70]. processing [Cum75, Gaz70]. produced [Gol72, MY73]. Professor [BL78]. profile [Tak76, WB77]. profiles [CT77]. progonka [Wei78]. Program [Tem79a, Aha72, Dow70, Les73, LBM^{+}78, VP75]. programming [MC72]. programs [MC78, PKPR72]. Projection [MS78b, Blo73, BB71, MBM76]. Projective [SH71b]. Propagation [CC70b, CC79a, CC71, CG76, Ila78, SE76, Smi74]. properties [Båt78, Che71b, Gil72, LSM71, Mar75, MS79b, NRH71, OS70, OU78, PW70, Ros76b, SW76, War78]. proposal [Joh76]. proteins [WGS72, WS73]. proton [KK75]. proton-proton [KK75]. protons [LW78]. Prüfer [Bai78]. pseudo [Roa78]. pseudo-spectral [Roa78]. pseudorandom [HR78]. Pseudospectral [FO73, WS78]. PSF [JP78]. PUFL [CB71]. pulsating [Coh76]. pulsed [PPS74]. pulses [CT77]. punch [Abo78].

Q.S [Ste77b]. QED [Sas76a]. Quadratic [Lut71]. quadrature [BKC72, CGR72, Del74, HRY73, HL76, Kni76, Lyn76, Mil70a, PCJ76, SK74b]. quality [SGP76]. quantal [GL77]. Quantum [AF79, Ben72, BK73, FPM76, FH74, PS76, WK79]. quark [KK75]. quarter [Ila78]. quasi [AG72, BW79, KB79a, Yee72]. quasi-conservative [AG72]. quasi-geostrophic [Yee72]. quasi-Lagrangian [KB79a]. quasi-Monte [BW79]. quasibound [PPP76]. quasilinear [NP74]. Quasilinearization [YL72]. quasineutral [BCCH78, HN78]. Quiet [HGE74, Ste77b].

Racah [KB73]. radial [AIC76, Dic72, DE70, HKN74, MKMR77, ML73, OU78, Tra71, WK79]. radially [Hew74, Lap71]. radiation [CGN70, CW76, Dav79b, FC71, SA73, SAHK75, Zin73]. radiative [Bar72, CP71, CHS78, PF79, SC73, WE77b]. Random [Cho76, Cho77, Fri71, Ama77, AF79, Blo76, CP74, Cho78a, Col74, CP79b, DS74, EEJ79, Gil78, MS77, TR72, YLY71]. range [BEH76, Fin78, RNP78, Win78]. Raphson [CFU74, CFUW74, FMW76]. Rapid [Cri78, Wal75, BKC72, CR75a, FR77, Fox77, Tra71]. rapidly [Per78]. rarefaction [Kaz70]. rates [CR75a, Gil78, Sul72]. Rational [BG70, Squ70, CCV72, Gru79]. ray [BS73, Blu71b, WGS72, WS73]. Rayleigh [GW78]. Rayleigh-wave [GW78]. Re [BD79]. reactant [CMM^{+}78]. reacting [Cho77, CH73]. Reaction [YK71, KDS79]. reactions [Gil76]. reactive [HCH72, Wes78]. Real [BS70, BK76, Dav75, RLT78]. real-symmetric [BK76, Dav75]. realistic [Owe71]. reciprocal [CF77]. reciprocal-lattice [CF77]. recirculating [Men74, Pop78, TO78]. recombination [OA78]. rectangular [CM73, Cav75]. Recurrence [dAPM77, CGF74]. Recursion [CZF78, Fox77, Wil71c]. recursive [Sha70a]. reduced [Blu71a, KM78, RG75, Ste79a]. reducing [Fos73, HP76, MY73]. Reduction [CLB74, KR71, HG79, MH79]. reductions [Det79]. reference [BM79, OU78, RB77]. Refinement [WGS72, WS73, Cav75, GFSW78, Hym77]. refinements [CS73a, Mil71c]. reflection [Chi70, Ryb70]. reflections [HP76]. region [Bar73, Men74, MC73, Rid79a, RNP75, Tra71]. regions [GO77, Rid79b]. regression [Cel76, CP76, WH75]. regularized [EM75, EM77]. related [BR71a, CS78, Nic75]. relation [Che71c, Lin70, Rei74b]. relations [CZF78, FH75, Fox77, Mor70a, Ros74, Ros76a, Wil71c]. relationship [HW75]. Relative [ML71]. Relativistic [HAN76, PL70, BKM73, MPS75, Mor78, Sho78]. relativity [Coh76]. relaxation [AH73b, BE76, Bir71, DT77, DMWV75, HU74, KB71, Mat70, Raf79]. Reliable [Ter79b]. relieve [WS73]. remainder [Che71a]. Remarks [EB79, MWW76, MB71]. remote [CT77]. removal [Dav79a]. renal [MKS77]. renormalized [CP71]. repeated [BR71a]. Reply [Saf78]. representation [FD79, Gas78, HBH78, HPS74, Mor79]. representations [Mar75]. repulsion [ED74, PH78b]. required [HP75, LN73]. residence [Gil78]. residue [Blo76]. resistive [DK77]. Resolution [Hai79, Bun73d, CCQ78, HGE74, HR78]. resolvent [Rei74a]. resonance [Dow70, PL70]. Resonant [Rei73, Kno75, RVD78]. respect [GW73]. rest [Ceb79]. restoration [JP78]. restricted [CP76]. results [Joh76, MGRK73, PMB73, Cam72]. Review [TO78]. Reynolds [Cho78a, iCS78, DG77, HR74, MS77]. rezoning [CC79b, HJKS78]. Riccati [Dav79a, Slo77, SB79]. Richardson [Ein76]. Riemann [LM74]. rigid [Abo78, LSM71, LM73, ML71, Sti71]. ring [CM73]. ro [Ale76]. ro-vibrationally [Ale76]. rocket [HCH72]. rod [Mey75]. role [HR74]. roll [Bak79, CB73]. roll-up [CB73]. Romberg [Mil70a]. Root [Wat71, BK76]. roots [Pfe79]. rotating [BL72, DS78, Hof74]. rotation [CZ79, Fox77, Wal75]. rotational [Bas72, Dal77, MS74, SS78, VP75]. rough [DS74, HC72, KB71]. roundoff [Fri78]. Rudeanu [Les73]. Rusanov [Got74]. Rys [KD76].

Saffman [Cho78a]. Salpeter [Fle73]. Salzer [Pex76]. sampling [TV77]. scalar [AB77, HPS74]. scale [BK78, Cla77, Dea71, Mud76, RVD78, Sch75b, TT78]. scale-dependent [Mud76]. scaling [OB79]. scattering [AS72, BS73, Blu71b, Bru74, BMP^{+}72, CC79a, CP73b, CW76, DS74, GW72, HK72, HP75, HPS74, Joh73, KK75, LN73, LNRY73, Mil71a, Mil74, OCS72, ON73, Rei73, RDW74, Sco79, Smi75a, SB73, Tho73, War79, Wim78, YK71]. scheme [AG76, AC73, AH73a, Cha76a, CC70a, CO75, DT77, FC71, GS77, Gol72, Got74, GT78b, HS76, KB79a, LS76, Lar78, Lip77, NC71, PS70, PR72, Van77a, Van77b, Wil70, Yee72, Zin73, vL74, vL79]. schemes [EGZ72, For77a, Gar74, Gar78, Gaz73, Gaz76, God79, Gus80, HZ72, Ila78, MM74, McK71, PW70, Rus70, Sas76b, SG78, SS78, SGP76, Ste79a, SW75, TAG76, Van77a, Wil78]. Schrödinger [AIC76, All70, CG70, CFUW74, CS78, Dic72, FRRT78, GSD72, HKN74, Hat70, HH76, Ixa72, ML73, Ste77a, Ste78b, TI73, Tra71, WO78]. Schrödinger’s [Fro73, GS70, GS71, TO79]. Schur [McK71]. search [Dow70]. searching [QB73]. second [BM79, BGP73, Cha76a, EGZ72, Gol72, GT74, KB79a, MM73, MM74, PR72, SH71a, Sco79, Ste79a, Ste76a, Tem76, Tur74, Wil71a, vL74, vL79]. second-moment [KB79a]. second-neighbor [SH71a]. second-order [Cha76a, EGZ72, Gol72, MM73, MM74, PR72, Sco79, Wil71a, vL74, vL79]. section [CM73]. sections [MY73]. seismograms [Job78]. selected [SGP76]. Selection [Hud72]. Self [HZ72, Kus73, BL73, BNPLD77, HL75, JCM74, LL75, SS71, Ste71, Ste70]. Self-adjusting [HZ72]. Self-consistent [Kus73, BNPLD77, JCM74, Ste70]. self-diffusion [SS71, Ste71]. self-gravitating [LL75]. semi [LR78, MP77]. semi-implicit [LR78, MP77]. semiconductor [Ste70]. Semiempirical [LW78]. semiquantitative [GH77]. sensing [CT77]. sensitive [Tro76]. Sensitivity [DG76, KDS79, CLS78]. Separable [BWSE76]. separated [KW72, Slo77]. Separating [CKW79]. sequel [vL79]. sequence [BW79, HR78]. sequences [CGR72]. Sereny [Sun75]. Series [Gas78, Bol78, Che71a, Del74, DS78, Hof74, Mor79, Pat78, Sal77, TLM73]. several [CS77, SBPH73, Sod78]. shallow [Gus71, SW70, Vre79]. shallow-water [SW70, Vre79]. shape [DS75]. shaped [TM70]. Shaping [EH74]. sharp [Cha75, RT76]. SHASTA [BB73]. shear [KC72]. sheet [CB73, Cho78b, ZG73]. sheets [Bak79]. shell [LS78]. shells [Ryb70]. shifts [Kau72, Kau73, RDW74]. Shock [CC71, CG76, AG72, CC70b, HZ72, WO75]. shocked [Lap71]. shocks [RZ75, Vie73]. Shooting [CK71, BL72, CP73a, Jon73, KH75b]. Shuman [SW76]. sided [DW78]. signals [Cum75]. significant [ME72]. significantly [HP76]. similarity [DE70]. Simple [Foo73, AH73a, CT77, CC79b, Gar76, MS78a, NR71, Orl76, TR76, Tro76]. simplest [TR72]. simplex [MC72]. simplified [AH70, Pev71]. simulated [Lin70, Wil71b]. simulating [Gil76]. Simulation [Bir71, Har78, Mat70, RD75, AMI75, And75, Bir77, BL75, BL73, BKL75, Bra76, BKM73, BGL74, BNPLD77, Cam71, CLB74, CO75, CO77, Chr73, Den72, Den74, EEJ79, Gil78, GJ75, HLM^{+}76, Hew74, HN78, JKAR77, JB72, Kno73, KS74b, KDR73, Lan70, Lan73, Lan79, LTD79, LO78, LSW72, Lew72, LN75, NTK79, Oku72, RVD78, SK72, Ste77b, TA77]. simulations [BCCH78, GJ73, JJGG78, Joh76, Mil76, Sch75b]. simultaneous [Les73, Raf79]. single [HG79, LW78, RS72b]. single-particle [LW78]. single-valued [RS72b]. singular [CI74, DS74, FPM76, FGW73, Ise73, PH70, TI79]. singularities [CF78b, Dav79a, TI79, WM71]. singularity [Lyn76, PCJ76, War79]. sinh [McD74]. sinh-Poisson [McD74]. Sixth [GW72]. Sixth-order [GW72]. size [HW75, SS76, Two75]. sized [AMI75]. Skan [CK71]. slab [Bun73c, EHMM70, HSM74, KSV70, Shu73]. slabs [Cla73]. slightly [Bai78]. slip [Bar78, Tay70]. sloping [SW70]. slow [Dev74]. Slowly [Coh76]. Small [MY71, Cla77, Hub71]. small-scale [Cla77]. Smooth [CS78, Fos73, RS72b, TG77]. smoothing [ASM71, Bol78, Den72, Ste79a]. solid [Hol77, KH75a, NRH71]. solids [And73, Gil72, GDZK70, LSM71, LM73, Sam71, Vie73, Whi74]. solitary [Cha74, EM77]. Solution [BM77, Cho73, HK72, Jon73, KH75b, ML77, MR73, MAT74, PR72, RS72a, AG72, Abo78, ASM71, AIC76, AW74, All70, AGGM79, Ant76, AB77, AHMB74, BS79, BR74, BKC72, BE79, Bos70, BGP73, Bru74, Bru73, Cam71, CG70, Can71, Can73, CP73a, CGF74, CFU74, CFUW74, Cel76, Cha74, CD78b, CN75, Cho76, CP79b, Cor75, CM74, CF78b, Dav77, Dav78, DS74, EB79, Ede73, EGV78, EHMM70, EH73, FRSW79, FM73, FMW76, FB79, Fro72, Fro73, Fro78, GS77, GCS75a, GCS75b, Gaz75, GS78, GH77, GS70, GS71, GSD72, GO77, GM79b, HZ79, HKN74, HE78, Hat70, Hau79, HH76, HS76, Hil78, Hil79, HD74, HPZ78, Jam77, JB75, JSK72, JPC72, Kan76, KSV70, Ker78, LWH78, Les73, LC72, LoD70, LM74, MdVD73, MGRK73, Mar78, MS74]. solution [Mar75, MWW76, MM73, MS79a, ML73, MC73, Mor70b, Mor71, O’B69, PB76a, RB77, Rid79a, Rid79b, RB71, RS76, SS76, SK77, SC72, She79, SG76, SG77a, SG77b, SG78, SH71b, SBD75, SB75, Smi75a, SW79, SV78, Sul72, Swa74, Swe73, Taf76, Tak76, Tay70, THM78, Tem75b, Tem79b, TI79, Tho73, TO79, Tra71, TT78, War79, WS78, WH75, Yee70, Yee76, YL72]. Solutions [Gra76, Ale76, And74, AH73c, BBA73, BD79, Coh78, DM73, EJ73, Han72, Hir75, HCH72, KM78, McD74, Sch77, SAK75, Ste77a, SC73, TTMW77, Thi78, WM72, WE77a, ZR73]. solve [OA78]. solver [BT79]. Solving [Mey75, AH73b, Blu71a, CCV72, CK71, DL72, DC74, Dic73, DE70, FPM76, Fle74, FRRT78, Gus71, HB77, Ixa72, KB79a, KB79b, Lau71, Le 72, MM74, Nic75, Rea70, SK74b, TNM72, Tru72, Was72, Whi79]. Some [Bee76, Hun78, MB71, OU78, Tem79b, AM79, Bel70, BEH76, BL73, Dav73, FPM76, FH74, HP75, Job78, LSW72, Lyn76, SS78, Win78]. Sommerfeld [Gas78]. source [Bar72, Jam77, OA78, WSW78]. space [AS72, Bar70a, BWSE76, BJ77, CD78a, CK76, Den72, EF77, FJ71a, Gaz73, Gaz75, GM70, Kus73, Lin75b, Mil70b, Mor79, Mor78, Str77, Wil70, Wil78]. space-charge-flow [Kus73]. space-time [BJ77]. sparce [Raf79]. spatial [BK73, HH77, Lan70, Oku72, OS74, Sch77]. special [GH73, OU78]. specified [Dic72, WGS72, WS73]. spectra [Bud79, CCQ78, DGGJ75, Dow70, EF77, MY73, PB75, VP75]. spectral [Gil72, Quo76b, Roa78, SK75, SE76, Tan79]. spectroscopy [JCM74]. spectrum [ASM71, BM79, Gru78]. speed [KSV70, RT76]. speeds [HA71, HAC74, Pra75]. sphere [AYMS71, LL75, OU78, Sti71, Swa74, Yee76]. spheres [DS78, MR71, Tan79, WK79]. spherical [AHMB74, Blu71b, CP73a, Kaz70, LS78, dAPM77, dAPM79, MC76, Ryb70, Vie73, Wil71a, Wil71c, Zin73, de 77]. spheroidal [OU78]. Spin [Ben79, BKL75, BB71, Dow70]. Spin-adapted [Ben79]. spinor [Hil79]. spline [Fle74, Tak71]. splines [Dav78]. Splitting [SG78, GS77, Jon79]. splittings [Run75]. square [Bey71, CP76, Quo76a]. squares [BH78, BDPH76, Cur70, Dow70, JPC72, RF71]. stabilities [ML71]. Stability [Bru74, GL76, GJ73, Ila78, Jon79, Ste79c, BL72, CD71b, Fro72, GT74, Got74, McK71, Mil71b, SW76, Sch75a, SK74b, Tay70, TR76, Tur74, WH74]. Stabilization [CM74, Van74]. stabilizers [EGZ72]. Stable [Ale76, AK72, HS76, PS70]. staggered [SS76, Swe73]. standard [Ama77, BDPH76, Col74]. stars [Hoh72]. start [Ste77b]. started [Ceb79, PL75]. state [AYMS71, BLV76, CGF74, CR75b, Emp70, FR77, HH76, Lad71, Ros77, RF71, Sch77, YL72]. states [AMS73, Bas72, Bud79, JCM74, Kno75, Lav73, PPP76]. static [JN71]. stationary [Hoh72, MR73, Wei78]. statistic [JMW77]. Statistical [BL75, Hub71, VW77, BK73, Cha73, CM76b]. statistics [MR71, TR72]. Steady [GL78, Ros77, Bri74, DIC79, FR77, Hof73, HN70, Sch77]. steady-state [Sch77]. steam [HM75]. steep [Bar72]. Stefan [BJ77, BJ79]. stellar [Taf76]. step [Gar74, Hud72]. stepsize [Hud72]. Stiff [Gel72, BG70, Ros77]. stimulated [Har78]. STO [Rei74a]. stochastic [Gil76]. stochasticity [FS73]. Stockmayer [Ves77]. Stockmayer-type [Ves77]. Stokes [AB77, Bar78, BM77, Chi76, GCS75a, GCS75b, HB77, PR72, SB75]. stopping [KB73]. strategy [BH72]. stratified [BL72, Cha76a]. stream [Bar78, BGL74, Roa70]. street [Pat78]. strength [CC70b]. strengths [Dic72]. stress [AH73b, Chi70]. stretching [HU74]. strong [Lie78]. strongly [Jac73]. Structural [Båt78]. structure [ME72, WS73]. structured [GJY75]. structures [HC72, MPB73, NH73, PMB73]. studied [BGL74]. studies [CGG^{+}78, Hof73, Mil71b, NP74]. study [BD73, CF76, CS70, Che71b, Cor74, DGGJ75, EM75, EM77, For77b, FS73, Men74, Quo76a, Quo76b, TNM72, Vre79]. studying [Foo73, FT77, LO78]. Sturm [Bai78, Har76]. style [RB71]. subdivision [CS73b]. Subgrid [Bun73d, Sch75b, Dea71]. submerged [NH73]. subprograms [Ada78]. subroutine [RS72b]. subroutines [Sca71]. subsonic [KW72, Raj73]. Successive [BE76, AF79, Rig79]. such [BO75]. Sufficiency [Roa70]. sum [WE77b]. Summary [Bol78]. sums [CF77, Gru79, LH72, Pev71, Stu78, Wim78]. superconducting [Bar73]. superconductor [AGGM79]. superconductors [DM73]. superposition [FJ71b]. Supersonic [Chu70, CH73, FR77, Hof73, HN70, MB75, Men74, SAK75, TM70, WCL78, Wei78]. Suppression [WD78]. surface [CC79a, CC71, CC70b, Cro70a, DS74, DG77, EC72, Eis79, HCB70, MS74, NH71, NH73, RS72b, Swa74, Yee76]. surfaces [Bro79]. survey [DM79, Sod78]. Sweden [Båt78]. swirling [LC72]. SYMBAL [CJ74].

Symbolic [CH70, AH73c, OFM71, PKPR72, RB71, Sto77]. Symmetric [BS70, BH78, BK76, Dav75, HF71, Lap71, SBPH73]. Symmetry [Blo73, Flo77, Bud79, JE74, Pac74, Sch70, SS78]. system [Bra76, CF76, Cla73, Eis78, EGV78, GW73, Hat70, RT79a, RCB77, SK72, SS78, TTM74]. systems [AG72, BW76, BR71b, BG70, BKL75, Cho76, Coh76, CD71a, CLS78, DG76, EB79, FS73, Fro73, GS78, Gel72, GM70, Har76, Hed79, Hub71, HCH72, JH74, Ker78, MM73, MM74, MR71, Nic75, Rig79, Ros76b, Ros77, RNP78, Rus70, Sco73, Slo77, Sod78, Sti71, Sto78, Taf76, Tem75b, TTMW77, TTM77, Vin74, Wes78]. systems-The [AG72].

tables [HM75]. targets [Rei73]. technique [Abr73, AG75, AH70, Bak79, BKC72, BB71, CP74, Cho71a, CC79b, God79, GJA79, HAN76, HE78, HRY73, Hir75, JD74, Jon79, KB79b, KH75b, Mil70a, MPB73, PMB73, RT76, Roa78, RS74, SW76, SG77a, YL72]. techniques [Dow70, EC72, Jon73, LK73, Lyn76, Mil74, Nic75, RDW74, RS72a]. Telenin [Men74]. temperature [Mod75, RR70, Tak71]. temperatures [WK79]. Tensor [DR79, Ben75, CB71]. terms [OA78]. terrain [Cla77, HC72]. terrain-following [Cla77]. testing [Joh76, RLT78]. their [Tem79a]. theorem [EJK^{+}73]. Theory [Che70, CO75, Squ70, BH78, Boy74, Cha73, DS74, FPM76, HH77, Jen78, KSV70, KH75a, LN73, ME72, Mil71a, OCS72, ON73, Pat71, PH78a, Sco79, Wim78]. thermal [Hoc71, Mor70b, Mor71, MN73, MN74]. thermally [DC74, Har78]. Thermodynamic [NRH71, Che71b]. thermodynamical [Mod75]. thermodynamics [BK73, MT78]. thermoluminescence [Har78]. thermonuclear [AL76]. thin [Bar73, Blu71a, CW76, MPB73, PMB73]. thin-film [Blu71a]. thin-wire [PMB73]. Third [BM70, Tem75a, BJ79, Got74, Rus70, Ste79a, Ste76a]. third-order [Got74]. thorough [PF74]. those [Tem79a]. Three [DG77, MB75, AG76, AB77, BLV76, Bri74, CO77, CH73, DIC79, God79, HC72, Lad71, Lar78, Lau79, LN75, NR71, NC71, NH73, OFM71, Pra75, RL72, RS72b, SK75, Sch75a, WCL78, Wei78, WE77a, BWSE76]. three-body [OFM71, BWSE76]. Three-dimensional [DG77, MB75, AB77, Bri74, CO77, God79, HC72, Lar78, Lau79, NR71, NC71, NH73, Pra75, RL72, RS72b, SK75, Sch75a, WCL78, Wei78]. threshold [BH78]. tidal [MGRK73]. tilted [Quo76a]. Time [Gaz76, Mar78, Moc76, Pre78, Bol78, BJ77, BK78, Cha76a, Cho73, EGZ72, FR77, FC71, For73, Gar74, Gil76, Gil78, GT78a, GK79, HPZ78, JJGG78, Jon79, Lan79, Lin75b, MPB73, Pac74, PMB73, RVD78, RB77, RB72, SGP76, Ste71, Wat75, Wil70, Zin73]. time-centering [For73]. Time-dependent [Mar78, Pre78, GT78a, HPZ78, JJGG78, Pac74, RB72, Wat75, Zin73]. Time-differencing [Gaz76]. time-domain [MPB73, PMB73]. time-marching [FR77]. time-splitting [Jon79]. times [Hoc71]. tokamak [BK78, HMR77, KB79b, MKMR77]. tokamaks [DGGJ75, JJGG78]. TOMCAT [TTM77]. topography [GCS75a, MS78a]. toroidal [CO77, GJ73, GJ75, JKAR77, JDG^{+}79, Ste78a, WJD70]. torsional [LS78]. Trace [FB76]. Traces [Lie78]. transfer [Bar72, CHS78, CGN70, GH73, Men74, PS72, PPS74, PF79, Smi75b, SC73]. Transfinite [HM75]. Transform [EJ73, AG75, DM79, Gaz76, Lon72, Pex76, SG76, Voj73, Abr73]. transformation [AMS73, Bai78, Bar72, Chu71, Cla77, Flo77, GCS75b, PF74, PH78a, Ros71]. transformations [Ben72, EEJ79, HG79, OB79, RZ75, Slo77, ZR73]. transforms [Abr75, DW78, Ein76, GO77, HPZ78, Lad71, Le 72, MB71, Ng74, RB72, Sal75, Sal76, Tal78]. transient [FB79, MdVD73, Pra71, Sch77, SK78, SB75, Ste79b]. transients [HCH72]. transition [Fre73, Gil78, RNP75, VW77]. transmission [WE77b]. Transonic [Ken70, Eis78, Kor78]. Transport [PF79, SS71, And75, AL76, BBH75, BB73, BB76, Bur78, CP73a, CS76, CC79b, EHMM70, FC71, HMR77, HK72, HP78, KSV70, KB79b, MKMR77, OS70, RR70, RB72, Zal79]. Transport-diffusion [PF79]. transport/Fokker [MKMR77]. transsonic [Bel70]. transverse [CD78a]. trap [Pat71]. trapped [Dav79b]. treatment [CHS78, CR75a, CI74, JE74, Kov70b, LG74, MB77, NR71, RZ75]. tree [WL71]. triangular [Che71b, FB79]. tridiagonal [Tem75b]. TRIGMAN [CJ74]. triplet [JCM74]. Troesch [CN75, Jon73, KH75b, RS72a, RS76]. truncated [Che71a, JMW77, TLM73]. truncation [DS78, Pat78]. turbulence [For77b, FO73, GA72, GP70, JB72]. turbulent [KW72, LC72, Pop78, Sch75b, THM78]. Two [BS73, HMR77, OB79, SA73, SK78, Vre79, AG76, ASM71, AW74, Ant76, AB77, Bar70a, BR74, Blu71a, BJ77, BGL74, BK78, CF76, CFUW74, Cha76b, CN75, CC70b, CM76b, DS78, DL72, DW78, EH74, FRSW79, FR77, FO73, FRRT78, FS73, Fro73, God79, GM70, Gru79, HH77, Hoc71, Hof74, Jon73, Jos71, Kno75, KH75b, Lad71, LN75, LR78, LK73, MB77, MPS75, McD74, MD78, MAT74, MS77, NR79b, PR72, PS76, RT76, RT79a, Rid79b, RS72a, SK75, Sha70a, SG77a, SG77b, SG78, SB75, Ste79c, Str78, TTMW77, TTM74, TTM77, Tro76, WCL78, Wil70, ZHR79]. two- [AB77, God79, SK75]. Two-center [BS73, CFUW74, MPS75, PS76, RT79a, Sha70a]. Two-dimensional [HMR77, SA73, SK78, BR74, BJ77, BK78, EH74, FRSW79, FR77, FO73, FRRT78, Hoc71, LN75, LK73, McD74, MS77, PR72, Rid79b, SG77a, SG77b, SG78, SB75, Str78, TTMW77, TTM74, TTM77, WCL78]. two-electron [Cha76b, Fro73, MD78]. two-fluid [Ste79c]. Two-layer [Vre79]. two-particle [CF76]. two-phase [LR78, RT76, Ste79c]. two-point [AW74, Ant76, CN75, DL72, Gru79, KH75b, MAT74, NR79b, RS72a, Tro76]. two-sided [DW78]. two-stream [BGL74]. type [DM73, Gat75, Gru79, Ics70, LN73, TI79, Thi78, Ves77, Wim78].

ultimate [Van77a, Van77b, vL74, vL79]. Umbrella [TV77]. unbiased [Kov70b]. unbounded [GO77, Orl76]. unconditionally [PS70]. uncorrelated [EF77]. underwater [FM78]. unequal [Bol78]. unfolding [ASM71]. unified [GP70]. uniform [DS75, OS74, Ste79a]. Uniformity [HR78]. unsteady [Cho73, FE72, OB79, PR72, TM70, YL72]. unsymmetrical [HN70]. upon [Mor79]. Upstream [Van77a]. Upstream-centered [Van77a]. Use [AH73c, Bos70, Dow70, Ein76, Le 72, MC72, RB72, SL72, SK75, Bee76, CGN70, CC79b, CD71a, Dav78, FGW73, GCS75b, GL77, JCM74, McK71, Mud76, SK74b, Wil71c, dR72]. used [ME72, Roa70]. useful [Bai78, Bar72, Cam72, CJ74, NG72]. using [AMS73, And75, BH78, BM79, Cav75, CO75, CN75, Cho78a, Cla77, CH73, CG76, DL72, DGGJ75, FPW73, FB79, Gra76, GJ75, Gru79, GM79b, HRY73, HS76, HB77, KD76, KK75, Kus73, MH79, MS77, NR79a, NR79b, PR72, PL75, PB76a, RB71, Slo77, Ste79c, Sto77, TTMW77, VT77, Whi79]. Utility [BS79].

V [vL79]. vacancies [LM73]. vacuum [AHMB74]. valid [GJA79]. Validity [Mil76]. value [AW74, BS79, Bol78, Cel76, CP73b, Chi75, CN75, Dav73, DL72, Det79, Dic73, EC72, FPM76, GH77, HD79, Jon73, KR71, KS73, KH75b, MAT74, Moc76, Mor70a, NG72, NR79a, NR79b, RS72a, Sas76b, Sco73, Tro76, Tru72, WD78, Wat71, WM72, WH75]. valued [RS72b]. valves [Pes72]. vapor [WCK71]. vapor-liquid [WCK71]. Variable [Van70, Fle79, HD74, Hud72, Kau72, MB77, Mil70a, Sal77, Sam71]. variable-step [Hud72]. variables [GM70, Mor78, Tal78, Wil78]. Variance [MH79]. variation [TI73]. variation-of-parameters [TI73]. Variational [FM78, Lew72, MPS75, PS73, Sas76b, Bar78, LNRY73, OCS72, ON73, PS72, RG75]. vector [AB77, Ben79, CM76a, Mor79, OS74]. vectors [JMW77, MC78]. velocity [PT75]. very [SBPH73]. via [Cam71, RB77]. vibrational [Båt78, Bas72, EJK^{+}73]. vibrational-rotational [Bas72]. vibrationally [Ale76]. vicinity [NH73]. view [Gel72]. viscoelasticity [Lon72, Voj73]. viscosity [DR79, HW75, HD74, Lut71, Roa72, She79, Str78, Whi73a, Whi73b, Whi74]. Viscous [Hof74, BE79, BD73, Bri74, Cro70a, Cro70b, Eis78, GM79a, Hai79, HLB79, MB75, MS70b, MC73, MY71, RL72, RK77, TNM72, TTMW77, TR72]. Vlasov [CGF74, CK76, Che77, EJ73, GS77, Gaz75, JKM71, Lew70, SK74a, SG76, SG77a, SG77b, SG78]. Volterra [Cha73, Kni76]. Volume [Ano72d, Ano73a, Ano73b, Ano73c, Ano74e, Ano70a, Ano70b, Ano71a, Ano71b, Ano72c, Ano75c, Ano75d, Ano76b, Ano76c, Ano76d, Ano77b, Ano77c, Ano77d, Ano78c, Ano78d, Ano78e, Ano78f, Ano79b, Ano79c, Ano79d, Ano79e, Hol77]. volumes [EF77, MR71]. Voronoi [BDF78, Fin79]. Vortex [Cho78b, Bak79, Bos70, CB73, Pat78, She79, ZG73]. vortex-ADI [She79]. vortices [Cho78a, Chr73, MS77]. vorticity [Bar78, Bea71]. Vries [CG77, GM76].

wake [Bos70]. walk [AF79, Cho78a, MS77]. walks [Gil78]. wall [KC72]. walls [Vie71]. water [CS70, Fin78, Gus71, SW70, Vre79, ZG73]. Watson [AL73]. Watson-like [AL73]. Wave [GW78, Blo73, Blu71a, CC79a, Cha74, CC71, CC70b, CG76, EM75, EM77, Fle74, Fle73, Flo77, GL77, HPS74, HK76, Ila78, Kaz70, KM78, Lin75b, dAPM77, PS76, RT79a, Sam71, SE76, Sch77, Smi74, SW79, Ste76b, Ste77a, Str77]. wavefunctions [Bot70, Emp70]. wavelength [Bal76]. Waves [GJY75, AM79, BV76, CS70, Chi70, Dal77, EM77, Quo76b, RVD78, SV76, WO75, ZG73]. weak [ZR73]. weakly [SK77]. wedge [GW78]. weights [BNP78]. Wendroff [GM70, Hun78]. Weyl [BK71]. which [JB75, Mar75, PKP79]. Whistler [Den74, RVD78]. width [Mil70a]. Wiener [FJ71b]. wind [Cro70a, Cro70b, WL71]. wind-driven [Cro70b]. wind-tree [WL71]. wire [MPB73, PMB73]. within [Bar73]. without [GL77]. works [BB73]. Wouthuysen [Ros71]. Wyk [Hud72].

x [BS73, Blu71b, JCM74, WGS72, WS73]. x-ray [BS73, Blu71b, WGS72, WS73]. XII [CS73b].

yield [Dic72].

Zachariasen [War79]. Zare [Dic73]. Zeros [Tem79a]. Zipora [BL78]. zone [Hol77, LH72, Sin72]. zones [HMB73].

References

Agmon:1979:AFD

[AAL79] N. Agmon, Y. Alhassid, and R. D. Levine. An algorithm for finding the distribution of maximal entropy. Journal of Computational Physics, 30(2):250–258, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

           sciencedirect.com/science/article/pii/0021999179901025.

Aregbesola:1977:VSP

[AB77] Y. A. S. Aregbesola and D. M. Burley. The vector and scalar potential method for the numerical solution of two- and threedimensional Navier–Stokes equations. Journal of Computational Physics, 24(4):398–415, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

           sciencedirect.com/science/article/pii/0021999177900304.

Appert:1975:NFE

[ABGR75] K. Appert, D. Berger, R. Gruber, and J. Rappaz. A new finite element approach to the normal mode analysis in magnetohydrodynamics. Journal of Computational Physics, 18(3):284– 299, July 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900030.

Aboudi:1978:NSP

[Abo78] Jacob Aboudi. A numerical solution to the problem of dynamic indentation of an elastic plate by a rigid punch. Journal of Computational Physics, 29(3):318–327, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901365.

Abramson:1972:IMC

[Abr72] Harvey D. Abramson. Intrinsic momentum computation of a charged particle in a magnetic field. Journal of Computational Physics, 9(2):372–374, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900290.

Abramovici:1973:ACF

[Abr73] Flavian Abramovici. The accurate calculation of Fourier integrals by the Fast Fourier Transform technique. Journal of Computational Physics, 11(1):28–37, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901460.

Abramovici:1975:AFF

[Abr75] Flavian Abramovici. The accuracy of finite Fourier transforms. Journal of Computational Physics, 17(4):446–449, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900492.

Abramovici:1978:GCM

[Abr78] F. Abramovici. A generalization of the Cagniard method. Journal of Computational Physics, 29(3):328–343, December 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178901377.

Ahmad:1973:NIS

[AC73] A. Ahmad and L. Cohen. A numerical integration scheme for the N-body gravitational problem. Journal of Computational Physics, 12(3):389–402, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901605.

Adam:1977:HAC

[Ada77] Yves Adam. Highly accurate compact implicit methods and boundary conditions. Journal of Computational Physics, 24(1): 10–22, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177901061.

Adams:1978:FSF

[Ada78] John C. Adams. FORTRAN subprograms for finite-difference formulas. Journal of Computational Physics, 26(1):113–116, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901031.

Alaylioglu:1973:EOI

[AEH73] Ayse Alaylioğlu, G. A. Evans, and J. Hyslop. The evaluation of oscillatory integrals with infinite limits. Journal of Computational Physics, 13(3):433–438, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900478.

Anderson:1979:QCR

[AF79] James B. Anderson and Bart H. Freihaut. Quantum chemistry by random walk: Method of successive corrections. Journal of Computational Physics, 31(3):425–437, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917990055X.

Abarbanel:1972:NSQ

[AG72] S. Abarbanel and M. Goldberg. Numerical solution of quasiconservative hyperbolic systems-the cylindrical shock problem. Journal of Computational Physics, 10(1):1–21, August 1972. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999172900873.

Alterovitz:1975:MMA

[AG75] S. Alterovitz and M. Gershenson. A moments method applied to Laplace transform technique for experimental physics. Journal of Computational Physics, 19(2):121–133, October 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900844.

Abarbanel:1976:NLF

[AG76] Saul Abarbanel and David Gottlieb. A note on the leap-frog scheme in two and three dimensions. Journal of Computational Physics, 21(3):351–355, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900334.

Alsop:1979:NSM

[AGGM79] L. E. Alsop, A. S. Goodman, F. G. Gustavson, and W. L. Miranker. A numerical solution of a model for a superconductor field problem. Journal of Computational Physics, 31(2):216– 239, May 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900706.

Amsden:1970:SMT

[AH70] Anthony A. Amsden and Francis H. Harlow. A simplified MAC technique for incompressible fluid flow calculations. Journal of Computational Physics, 6(2):322–325, October 1970. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917090029X.

Amsden:1973:SSG

[AH73a] A. A. Amsden and C. W. Hirt. A simple scheme for generating general curvilinear grids. Journal of Computational Physics, 11(3): 348–359, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900788.

Andrews:1973:RMS

[AH73b] D. J. Andrews and Steven L. Hancock. A relaxation method for solving nonlinear stress equilibrium problems. Journal of Computational Physics, 12(2):202–209, February 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999173800117.

Atherton:1973:USC

[AH73c] Robert W. Atherton and George M. Homsy. Use of symbolic computation to generate evolution equations and asymptotic solutions to elliptic equations. Journal of Computational Physics, 13(1):45– 58, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901241.

Aharony:1972:CPM

[Aha72] Amnon Aharony. A computer program for molecular dynamics of dilute gases. Journal of Computational Physics, 10(2):341– 353, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917290071X.

Auerbach:1974:ASS

[AHMB74] T. Auerbach, W. Hälg, J. Mennig, and H. Beuchat. Analytical solution in spherical geometry of the monoenergetic S_{2} and S_{4} differential equations, and of the S_{n} equations in vacuum. Journal of Computational Physics, 15(2):168–187, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900825.

Adam:1976:FOP

[AIC76] Gh. Adam, L. Gr. Ixaru, and A. Corciovei. A first-order perturbative numerical method for the solution of the radial Schrödinger equation. Journal of Computational Physics, 22(1):1–33, September 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176901078.

Anderson:1972:CSE

[AK72] David V. Anderson and John Killeen. Computation of stable equilibria in minimum-B mirror devices. Journal of Computational Physics, 10(1):133–150, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900952.

Alspaugh:1970:AII

[AKK70] D. W. Alspaugh, H. H. Kagiwada, and R. Kalaba. Application of invariant imbedding to the buckling of columns. Journal of Computational Physics, 5(1):56–69, February 1970. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900525.

Austen:1973:ELG

[AL73] D. J. Austen and P. D. Loly. On the evaluation of lattice Green functions and Watson-like integrals. Journal of Computational Physics, 11(3):315–325, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900752.

Antal:1976:CPM

[AL76] Michael J. Antal and Clarence E. Lee. Charged particle mass and energy transport in a thermonuclear plasma. Journal of Computational Physics, 20(3):298–312, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900838.

Alexander:1976:SSC

[Ale76] Millard H. Alexander. Stable solutions of the close-coupled equations for ro-vibrationally inelastic collisions. Journal of Computational Physics, 20(2):248–250, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900693.

Allison:1970:NSC

[All70] A. C. Allison. The numerical solution of coupled differential equations arising from the Schrödinger equation. Journal of Computational Physics, 6(3):378–391, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900379.

Alexander:1979:GMA

[AM79] M. E. Alexander and J. Ll Morris. Galerkin methods applied to some model equations for non-linear dispersive waves. Journal of Computational Physics, 30(3):428–451, March 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999179901244.

Amadori:1977:CDS

[Ama77] Rüdiger Amadori. Correlational defects in the standard IBM 360 random number generator and the classical ideal gas correlation function. Journal of Computational Physics, 24(4):450– 454, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917790033X. See [Col74].

Allard:1973:ENP

[AMHH73] J. F. Allard, D. E. Medjadi, B. Houssais, and M. Hadj Hassan. Evaluation of nuclear potential matrix elements in the deformed Cartesian basis. Journal of Computational Physics, 13(2):260– 265, October 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900259.

Abe:1975:GEP

[AMI75] Hirotada Abe, Junichi Miyamoto, and Ryohei Itatani. Grid effects on the plasma simulation by the finite-sized particle. Journal of Computational Physics, 19(2):134–149, October 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900856.

Afshar:1973:HTD

[AMS73] R. Afshar, F. M. Mueller, and J. C. Shaffer. Hilbert transformation of densities of states using Hermite functions. Journal of Computational Physics, 11(2):190–209, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917390003X.

Andrews:1971:CMP

[And71] D. J. Andrews. Calculation of mixed phases in continuum mechanics. Journal of Computational Physics, 7(2):310–326, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900921.

Andrews:1973:NMC

[And73] D. J. Andrews. A numerical method for creep deformation of solids. Journal of Computational Physics, 12(2):275–279, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999173800166.

Anderson:1974:CNS

[And74] D. A. Anderson. A comparison of numerical solutions to the inviscid equations of fluid motion. Journal of Computational Physics, 15(1):1–20, May 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900655.

Anderson:1975:AMS

[And75] D. V. Anderson. Axisymmetric multifluid simulation of high beta plasmas with anisotropic transport using a moving flux coordinate grid. Journal of Computational Physics, 17(3):246– 275, March 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900534.

Andrew:1978:CIM

[And78] Alan L. Andrew. Convergence of an iterative method for derivatives of eigensystems. Journal of Computational Physics, 26(1):107– 112, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917890102X.

Anonymous:1970:AIVa

[Ano70a] Anonymous. Author index of volume 5. Journal of Computational Physics, 5(3):612–613, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900835.

Anonymous:1970:AIVb

[Ano70b] Anonymous. Author index of volume 6. Journal of Computational Physics, 6(3):533–534, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900471.

Anonymous:1970:E

[Ano70c] Anonymous. Erratum. Journal of Computational Physics, 6(1): 155, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900173.

Anonymous:1971:AIVa

[Ano71a] Anonymous. Author index of volume 7. Journal of Computational Physics, 7(3):646–647, June 1971. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171901197.

Anonymous:1971:AIVb

[Ano71b] Anonymous. Author index of volume 8. Journal of Computational Physics, 8(3):506–507, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900301.

Anonymous:1972:Aa

[Ano72a] Anonymous. Addendum. Journal of Computational Physics, 9 (3):606, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900150.

Anonymous:1972:Ab

[Ano72b] Anonymous. Announcement. Journal of Computational Physics, 10 (1):167, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900988.

Anonymous:1972:AIVa

[Ano72c] Anonymous. Author index of volume 9. Journal of Computational Physics, 9(3):607–608, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900162.

Anonymous:1972:AIVb

[Ano72d] Anonymous. Author index: Volume 10. Journal of Computational Physics, 10(3):621–622, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900575.

Anonymous:1972:E

[Ano72e] Anonymous. Errata. Journal of Computational Physics, 10(3): 620, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900563.

Anonymous:1973:AIVa

[Ano73a] Anonymous. Author index: Volume 11. Journal of Computational Physics, 11(4):621–622, April 1973. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901435.

Anonymous:1973:AIVb

[Ano73b] Anonymous. Author index: Volume 12. Journal of Computational Physics, 12(4):567–568, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901083.

Anonymous:1973:AIVc

[Ano73c] Anonymous. Author index: Volume 13. Journal of Computational Physics, 13(4):613–614, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900375.

Anonymous:1974:Aa

[Ano74a] Anonymous. Announcement. Journal of Computational Physics, 15(3):427, July 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901223.

Anonymous:1974:Ab

[Ano74b] Anonymous. Announcement. Journal of Computational Physics, 16 (2):193, October 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901132.

Anonymous:1974:AIa

[Ano74c] Anonymous. Author index. Journal of Computational Physics, 15(4):533–534, August 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999174900795.

Anonymous:1974:AIb

[Ano74d] Anonymous. Author index. Journal of Computational Physics, 16(4):401, December 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999174900503.

Anonymous:1974:AIV

[Ano74e] Anonymous. Author index: Volume 14. Journal of Computational Physics, 14(4):420–421, March 1974. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900229.

Anonymous:1975:A

[Ano75a] Anonymous. Announcement. Journal of Computational Physics, 17(3):347, March 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900583.

Anonymous:1975:AI

[Ano75b] Anonymous. Author index. Journal of Computational Physics, 17 (4):455, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900510.

Anonymous:1975:AIVa

[Ano75c] Anonymous. Author index for volume 18. Journal of Computational Physics, 18(4):471, August 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900984.

Anonymous:1975:AIVb

[Ano75d] Anonymous. Author index for volume 19. Journal of Computational Physics, 19(4):435, December 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900741.

Anonymous:1975:E

[Ano75e] Anonymous. Erratum. Journal of Computational Physics, 19(3): 338, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900832.

Anonymous:1976:A

[Ano76a] Anonymous. Announcement. Journal of Computational Physics, 20 (1):117, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176901066.

Anonymous:1976:AIVa

[Ano76b] Anonymous. Author index for volume 20. Journal of Computational Physics, 20(4):502, April 1976. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900978.

Anonymous:1976:AIVb

[Ano76c] Anonymous. Author index for volume 21. Journal of Computational Physics, 21(4):454, August 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900401.

Anonymous:1976:AIVc

[Ano76d] Anonymous. Author index for volume 22. Journal of Computational Physics, 22(4):543, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900498.

Anonymous:1976:MN

[Ano76e] Anonymous. Meeting notice. Journal of Computational Physics, 22 (2):269, October 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900796.

Anonymous:1977:A

[Ano77a] Anonymous. Announcement. Journal of Computational Physics, 24(1):115, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177901140.

Anonymous:1977:AIVa

[Ano77b] Anonymous. Author index for volume 23. Journal of Computational Physics, 23(4):453, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900754.

Anonymous:1977:AIVb

[Ano77c] Anonymous. Author index for volume 24. Journal of Computational Physics, 24(4):456, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900353.

Anonymous:1977:AIVc

[Ano77d] Anonymous. Author index for volume 25. Journal of Computational Physics, 25(4):409, December 1977. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900067.

Anonymous:1977:E

[Ano77e] Anonymous. Erratum. Journal of Computational Physics, 23(4): 452, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900742.

Anonymous:1978:Aa

[Ano78a] Anonymous. Announcement. Journal of Computational Physics, 26 (1):117, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901043.

Anonymous:1978:Ab

[Ano78b] Anonymous. Announcement. Journal of Computational Physics, 28(2):294–295, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999178900414.

Anonymous:1978:AIVa

[Ano78c] Anonymous. Author index for volume 26. Journal of Computational Physics, 26(3):457, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900840.

Anonymous:1978:AIVb

[Ano78d] Anonymous. Author index for volume 27. Journal of Computational Physics, 27(3):453, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900220.

Anonymous:1978:AIVc

[Ano78e] Anonymous. Author index for volume 28. Journal of Computational Physics, 28(3):442, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900682.

Anonymous:1978:AIVd

[Ano78f] Anonymous. Author index for volume 29. Journal of Computational Physics, 29(3):447, December 1978. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917890147X.

Anonymous:1978:Ea

[Ano78g] Anonymous. Errata. Journal of Computational Physics, 28(3):439– 440, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900669.

Anonymous:1978:Eb

[Ano78h] Anonymous. Erratum. Journal of Computational Physics, 29(1): 145, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917890116X.

Anonymous:1978:LFAa

[Ano78i] Anonymous. List of forthcoming articles. Journal of Computational Physics, 28(1):133–134, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900542.

Anonymous:1978:LFAb

[Ano78j] Anonymous. List of forthcoming articles. Journal of Computational Physics, 28(2):296, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900426.

Anonymous:1978:LFAc

[Ano78k] Anonymous. List of forthcoming articles. Journal of Computational Physics, 28(3):441, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900670.

Anonymous:1978:LFAd

[Ano78l] Anonymous. List of forthcoming articles. Journal of Computational Physics, 29(1):146, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901171.

Anonymous:1978:LFAe

[Ano78m] Anonymous. List of forthcoming articles. Journal of Computational Physics, 29(2):311–312, November 1978. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917890164X.

Anonymous:1978:LFAf

[Ano78n] Anonymous. List of forthcoming articles. Journal of Computational Physics, 29(3):446, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901468.

Anonymous:1979:A

[Ano79a] Anonymous. Announcements. Journal of Computational Physics, 30(1):149–150, January 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900949.

Anonymous:1979:AIVa

[Ano79b] Anonymous. Author index for volume 30. Journal of Computational Physics, 30(3):453, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901268.

Anonymous:1979:AIVb

[Ano79c] Anonymous. Author index for volume 31. Journal of Computational Physics, 31(3):451, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900585.

Anonymous:1979:AIVc

[Ano79d] Anonymous. Author index for volume 32. Journal of Computational Physics, 32(3):448, September 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901566.

Anonymous:1979:AIVd

[Ano79e] Anonymous. Author index for volume 33. Journal of Computational Physics, 33(3):445, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901724.

Anonymous:1979:LFAa

[Ano79f] Anonymous. List of forthcoming articles. Journal of Computational Physics, 30(1):151, January 1979. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900950.

Anonymous:1979:LFAb

[Ano79g] Anonymous. List of forthcoming articles. Journal of Computational Physics, 30(2):306–307, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901062.

Anonymous:1979:LFAc

[Ano79h] Anonymous. List of forthcoming articles. Journal of Computational Physics, 30(3):452, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901256.

Anonymous:1979:LFAd

[Ano79i] Anonymous. List of forthcoming articles. Journal of Computational Physics, 31(1):152, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900676.

Anonymous:1979:LFAe

[Ano79j] Anonymous. List of forthcoming articles. Journal of Computational Physics, 31(2):300, May 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900755.

Anonymous:1979:LFAf

[Ano79k] Anonymous. List of forthcoming articles. Journal of Computational Physics, 31(3):450, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900573.

Anonymous:1979:LFAg

[Ano79l] Anonymous. List of forthcoming articles. Journal of Computational Physics, 32(1):144, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901475.

Anonymous:1979:LFAh

[Ano79m] Anonymous. List of forthcoming articles. Journal of Computational Physics, 32(2):288, August 1979. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901360.

Anonymous:1979:LFAi

[Ano79n] Anonymous. List of forthcoming articles. Journal of Computational Physics, 32(3):447, September 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901554.

Anonymous:1979:LFAj

[Ano79o] Anonymous. List of forthcoming articles. Journal of Computational Physics, 33(1):151–152, October 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900329.

Anonymous:1979:LFAk

[Ano79p] Anonymous. List of forthcoming articles. Journal of Computational Physics, 33(2):300, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917990024X.

Anonymous:1979:LFAl

[Ano79q] Anonymous. List of forthcoming articles. Journal of Computational Physics, 33(3):444, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901712.

Antar:1976:STP

[Ant76] Basil N. Antar. On the solution of two-point linear differential eigenvalue problems. Journal of Computational Physics, 20(2):208– 219, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900656.

Antia:1979:FDM

[Ant79] H. M. Antia. Finite-difference method for generalized eigenvalue problem in ordinary differential equations. Journal of Computational Physics, 30(2):283–295, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901049.

Adelberg:1972:NAM

[AS72] Michael L. Adelberg and Alvin M. Saperstein. Numerical aspects of momentum-space scattering equations for peaked potentials. Journal of Computational Physics, 9(3):398–411, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900022.

Ablow:1978:CC

[AS78] C. M. Ablow and S. Schechter. Campylotropic coordinates. Journal of Computational Physics, 27(3):351–362, June 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178900153.

Abu-Shumays:1971:SSU

[ASM71] I. K. Abu-Shumays and L. D. Marinelli. A smoothing solution (unfolding) of a two dimensional density function from its measured spectrum. Journal of Computational Physics, 7(2):219– 238, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900866.

Allen:1974:IIA

[AW74] Richard C. Allen, Jr. and G. Milton Wing. An invariant imbedding algorithm for the solution of inhomogeneous linear two-point boundary value problems. Journal of Computational Physics, 14(1):40–58, January 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900047.

Alder:1971:HSE

[AYMS71] B. J. Alder, D. A. Young, M. R. Mansigh, and Z. W. Salsburg. Hard sphere equation of state in the close-packed limit. Journal of Computational Physics, 7(3):361–366, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900970.

Bailey:1978:SMP

[Bai78] Paul B. Bailey. A slightly modified Prüfer transformation useful for calculating Sturm–Liouville eigenvalues. Journal of Computational Physics, 29(2):306–310, November 1978. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901638.

Baker:1979:CCT

[Bak79] Gregory R. Baker. The “cloud in cell” technique applied to the roll up of vortex sheets. Journal of Computational Physics, 31(1): 76–95, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900639.

Balanis:1976:LWA

[Bal76] George N. Balanis. Long wavelength approximations to periodic elastic media and numerical applications. Journal of Computational Physics, 20(3):279–297, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900826.

Bandopadhyay:1973:AEF

[Ban73] P. K. Bandopadhyay. Average electronic form factors for neutral atoms with 40 < Z < 90. Journal of Computational Physics, 12(2):280–284, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999173800178.

Barfield:1970:OMG

[Bar70a] W. D. Barfield. An optimal mesh generator for Lagrangian hydrodynamic calculations in two space dimensions. Journal of Computational Physics, 6(3):417–429, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900409.

Barreld:1970:NMG

[Bar70b] W. D. Barreld. Numerical method for generating orthogonal curvilinear meshes. Journal of Computational Physics, 5(1):23– 33, February 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900501.

Barfield:1972:TRT

[Bar72] W. D. Barfield. A transformation of the radiative transfer equation useful for problems with steep source gradients. Journal of Computational Physics, 10(3):420–431, December 1972. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999172900459.

Barker:1973:CPG

[Bar73] Derek C. Barker. The calculation of potential gradients in a moving normal region within a superconducting thin film. Journal of Computational Physics, 11(3):333–347, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900776.

Barrett:1978:VPS

[Bar78] K. E. Barrett. A variational principle for the stream functionvorticity formulation of the Navier–Stokes equations incorporating no-slip conditions. Journal of Computational Physics, 26(2):153– 161, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900876.

Bass:1972:OCV

[Bas72] James N. Bass. Optimized calculations of vibrational-rotational states of diatomic molecules. Journal of Computational Physics, 9(3):555–567, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999172900113.

Baath:1978:SVB

[Båt78] Markus Båth. Structural and vibrational bedrock properties in Sweden. Journal of Computational Physics, 29(3):344–356, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901389.

Bunge:1971:ESO

[BB71] Carlos F. Bunge and Annik Bunge. Eigenfunctions of spin and orbital angular momentum by the projection operator technique. Journal of Computational Physics, 8(3):409–416, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900209.

Boris:1973:FCT

[BB73] Jay P. Boris and David L. Book. Flux-corrected transport. I. SHASTA, a fluid transport algorithm that works. Journal of Computational Physics, 11(1):38–69, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901472.

Boris:1976:FCT

[BB76] J. P. Boris and D. L. Book. Flux-corrected transport. III. Minimalerror FCT algorithms. Journal of Computational Physics, 20(4): 397–431, April 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900917.

Barthes-Biesel:1973:CGA

[BBA73] Dominique Barthès-Biesel and Andreas Acrivos. On computer generated analytic solutions to the equations of fluid mechanics. The case of creeping flows. Journal of Computational Physics, 12(3): 403–411, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901617.

Book:1975:FCT

[BBH75] D. L. Book, J. P. Boris, and K. Hain. Flux-corrected transport II: Generalizations of the method. Journal of Computational Physics, 18(3):248–283, July 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900029.

Byers:1978:HSQ

[BCCH78] J. A. Byers, B. I. Cohen, W. C. Condit, and J. D. Hanson. Hybrid simulations of quasineutral phenomena in magnetized plasma. Journal of Computational Physics, 27(3):363– 396, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900165.

Bozeman:1973:NSV

[BD73] James D. Bozeman and Charles Dalton. Numerical study of viscous flow in a cavity. Journal of Computational Physics, 12(3): 348–363, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901575.

Benjamin:1979:CNS

[BD79] A. S. Benjamin and V. E. Denny. On the convergence of numerical solutions for 2-D flows in a cavity at large Re. Journal of Computational Physics, 33(3):340–358, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901608.

Brostow:1978:CVP

[BDF78] Witold Brostow, Jean-Pierre Dussault, and Bennett L. Fox. Construction of Voronoi polyhedra. Journal of Computational Physics, 29(1):81–92, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901109.

Brusset:1976:CSD

[BDPH76] Henry Brusset, Domnaique Depeyre, Jean-Pierre Petit, and François Haffner. On the convergence of standard and damped least squares methods. Journal of Computational Physics, 22(4):534– 542, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900486.

Benson:1976:SPR

[BE76] A. Benson and D. J. Evans. Successive peripheral over-relaxation and other block methods. Journal of Computational Physics, 21 (1):1–19, May 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900176.

Bercovier:1979:FEN

[BE79] Michel Bercovier and Michael Engelman. A finite element for the numerical solution of viscous incompressible flows. Journal of Computational Physics, 30(2):181–201, February 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999179900986.

Beardsley:1971:IPV

[Bea71] R. C. Beardsley. Integration of the planetary vorticity equation on a parabolic circular grid. Journal of Computational Physics, 7(2): 273–288, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900891.

Beeman:1976:SMM

[Bee76] D. Beeman. Some multistep methods for use in molecular dynamics calculations. Journal of Computational Physics, 20(2):130– 139, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900590.

Blakemore:1976:CSM

[BEH76] M. Blakemore, G. A. Evans, and J. Hyslop. Comparison of some methods for evaluating infinite range oscillatory integrals. Journal of Computational Physics, 22(3):352–376, November 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900541. See comment [Win78].

Belotserkovskii:1970:NMS

[Bel70] O. M. Belotserkovskii. Numerical methods of some transsonic aerodynamics problems. Journal of Computational Physics, 5(3): 587–611, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900823.

Bellomonte:1975:CCC

[Bel75] L. Bellomonte. Calculation of the Coulomb coefficients and the Madelung constant in a face-centered cubic crystal with additional charges located along the (100) direction. Journal of Computational Physics, 17(3):340–346, March 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900571.

Bender:1972:ITB

[Ben72] C. F. Bender. Integral transformations. A bottleneck in molecular quantum mechanical calculations. Journal of Computational Physics, 9(3):547–554, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900101.

Bennett:1975:MTM

[Ben75] Charles H. Bennett. Mass tensor molecular dynamics. Journal of Computational Physics, 19(3):267–279, November 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900777.

Bennett:1976:EEF

[Ben76] Charles H. Bennett. Efficient estimation of free energy differences from Monte Carlo data. Journal of Computational Physics, 22(2):245–268, October 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900784.

Bender:1979:SAV

[Ben79] C. F. Bender. Spin-adapted vector method: An alternative to the conventional configuration interaction approach. Journal of Computational Physics, 30(3):324–332, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901189.

Beyerlein:1971:HDH

[Bey71] Adolph Beyerlein. High density hard square particles approaching a close packed “brickwall” configuration. Journal of Computational Physics, 7(3):403–408, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917190101X.

Blue:1970:RAM

[BG70] James L. Blue and Hermann K. Gummel. Rational approximations to matrix exponential for systems of stiff differential equations. Journal of Computational Physics, 5(1):70–83, February 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900537.

Brown:1974:TSI

[BGL74] D. I. Brown, S. J. Gitomer, and H. R. Lewis. The two-stream instability studied with four one-dimensional plasma simulation models. Journal of Computational Physics, 14(2):193–199, February 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900138.

Boyko:1973:ISS

[BGP73] E. R. Boyko, B. J. Gilbert, and S. H. Puon. An iterative solution to the second order eigenvalue equation with periodic boundary conditions. Journal of Computational Physics, 12(2):171– 178, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999173800087.

Baker:1972:DSM

[BH72] George A. Baker, Jr. and D. L. Hunter. A diverse strategy method of minimization. Journal of Computational Physics, 9(2):273– 289, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917290023X.

Birge:1978:EET

[BH78] Robert R. Birge and Lynn M. Hubbard. Eigenvector eigenvalue theory: The diagonalization of symmetric matrices using a least-squares optimized threshold method. Journal of Computational Physics, 29(2):199–207, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901511.

Bird:1971:CSR

[Bir71] G. A. Bird. Comments on “Simulation of relaxation processes by Monte Carlo”. Journal of Computational Physics, 8(2):309– 310, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900131. See [Mat70].

Bird:1977:DMS

[Bir77] G. A. Bird. Direct molecular simulation of a dissociating diatomic gas. Journal of Computational Physics, 25(4):353–365, December 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900031.

Bonnerot:1977:NCF

[BJ77] R. Bonnerot and P. Jamet. Numerical computation of the free boundary for the two-dimensional Stefan problem by space-time finite elements. Journal of Computational Physics, 25(2):163– 181, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900195.

Bonnerot:1979:TOA

[BJ79] R. Bonnerot and P. Jamet. A third order accurate discontinuous finite element method for the one-dimensional Stefan problem. Journal of Computational Physics, 32(2):145–167, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917990127X.

Beck:1971:GWG

[BK71] Robert E. Beck and Bernard Kolman. Generation of the Weyl group on a computer. Journal of Computational Physics, 7(2): 346–353, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900945.

Bomberger:1973:NMC

[BK73] D. C. Bomberger and M. D. Kostin. A numerical method for calculating the spatial distribution function of quantum statistical thermodynamics. Journal of Computational Physics, 12(3): 308–314, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917390154X.

Butscher:1976:MDM

[BK76] W. Butscher and W. E. Kammer. Modification of Davidson’s method for the calculation of eigenvalues and eigenvectors of large real-symmetric matrices: “root homing procedure”. Journal of Computational Physics, 20(3):313–325, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917690084X.

Byrne:1978:GTD

[BK78] R. N. Byrne and H. H. Klein. G2M, a two-dimensional diffusion time scale tokamak code. Journal of Computational Physics, 26(3): 352–378, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917890075X.

Bellman:1972:DQT

[BKC72] Richard Bellman, B. G. Kashef, and J. Casti. Differential quadrature: A technique for the rapid solution of nonlinear partial differential equations. Journal of Computational Physics, 10(1):40– 52, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900897.

Bortz:1975:NAM

[BKL75] A. B. Bortz, M. H. Kalos, and J. L. Lebowitz. A new algorithm for Monte Carlo simulation of Ising spin systems. Journal of Computational Physics, 17(1):10–18, January 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900601.

Brettschneider:1973:NSR

[BKM73] M. Brettschneider, J. Killeen, and A. A. Mirin. Numerical simulation of relativistic electrons confined in an axisymmetric mirror field. Journal of Computational Physics, 11(3):360–399, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917390079X.

Brown:1972:ASM

[BL72] R. A. Brown and F. Lee. An application of the shooting method to the stability problem for a stratified, rotating boundary layer. Journal of Computational Physics, 10(1):107–122, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900939.

Boris:1973:NSF

[BL73] Jay P. Boris and Roswell Lee. Nonphysical self forces in some electromagnetic plasma-simulation algorithms. Journal of Computational Physics, 12(1):131–136, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901745.

Borgnakke:1975:SCM

[BL75] Claus Borgnakke and Poul S. Larsen. Statistical collision model for Monte Carlo simulation of polyatomic gas mixture. Journal of Computational Physics, 18(4):405–420, August 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900947.

Bolt:1978:MBP

[BL78] Bruce A. Bolt and Dan Loewenthal. Memorial biography [for Professor Zipora Alterman]. Journal of Computational Physics, 29(3):ii + 313–317, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901353.

Blokker:1973:SPC

[Blo73] Esko Blokker. Symmetry projection of crystal wave functions by means of a computer. Journal of Computational Physics, 12(4): 471–490, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900995.

Blood:1976:CPR

[Blo76] F. A. Blood, Jr. Correlations in power residue generated random numbers. Journal of Computational Physics, 20(3):372– 379, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900899.

Blue:1971:FAM

[Blu71a] James L. Blue. On finding the admittance matrix of a thinfilm network by solving the reduced wave equation in two dimensions. Journal of Computational Physics, 7(2):327–345, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900933.

Blum:1971:XRS

[Blu71b] L. Blum. X-ray scattering from liquids with nearly spherical molecules. Journal of Computational Physics, 7(3):592–602, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171901148.

Bassett:1976:DNA

[BLV76] I. M. Bassett, B. R. E. Lederer, and G. C. Vorlicek. A direct numerical approach to three nucleon bound state problems. Journal of Computational Physics, 22(1):34–73, September 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917690108X.

Burstein:1970:TOD

[BM70] Samuel Z. Burstein and Arthur A. Mirin. Third order difference methods for hyperbolic equations. Journal of Computational Physics, 5(3):547–571, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917090080X.

Briley:1977:SMC

[BM77] W. R. Briley and H. McDonald. Solution of the multidimensional compressible Navier–Stokes equations by a generalized implicit method. Journal of Computational Physics, 24(4):372– 397, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900298.

Blatt:1979:CSO

[BM79] David W. E. Blatt and Bruce H. J. McKellar. The computation of second order energies using a reference spectrum Euler function. Journal of Computational Physics, 32(1):89–100, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917990144X.

Burke:1972:IDE

[BMP^{+}72] G. J. Burke, E. K. Miller, A. J. Poggio, G. M. Pjerrou, B. J. Maxum, and W. Meecham. An integro-differential equation approach to acoustic scattering from fluid-immersed elastic bodies. Journal of Computational Physics, 10(1):22–39, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900885.

Bivins:1973:NCO

[BMP73] R. L. Bivins, N. Metropolis, and John R. Pasta. Nonlinear coupled oscillators: Modal equation approach. Journal of Computational Physics, 12(1):65–87, May 1973. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.osti.gov/energycitations/product.biblio.jsp?
osti_id=4457833&query_id=0;   http://www.sciencedirect.
com/science/article/pii/0021999173901691.

Bitoun:1973:CPH

[BNGF73] J. Bitoun, A. Nadeau, J. Guyard, and M. R. Feix. A computational procedure for high-order adiabatic invariants. Journal of Computational Physics, 12(3):315–333, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901551.

Boyko:1978:IWE

[BNP78] E. R. Boyko, Peter J. Nassiff, and J. C. Pezzullo. Integration weights for experimental data. Journal of Computational Physics, 27(3):443–446, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900207.

Busnardo-Neto:1977:SCM

[BNPLD77] J. Busnardo-Neto, P. L. Pritchett, A. T. Lin, and J. M. Dawson. A self-consistent magnetostatic particle code for numerical simulation of plasmas. Journal of Computational Physics, 23(3):300– 312, March 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900961.

Boris:1975:NEO

[BO75] Jay P. Boris and Elaine S. Oran. Numerical evaluation of oscillatory integrals such as the modified Bessel function K_{iζ}(x). Journal of Computational Physics, 17(4):425–433, April 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900455.

Bolt:1978:SVS

[Bol78] Bruce A. Bolt. Summary value smoothing of physical time series with unequal intervals. Journal of Computational Physics, 29(3):357–369, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901390.

Bossel:1970:UEI

[Bos70] Hartmut H. Bossel. Use of exponentials in the integral solution of the parabolic equations of boundary layer, wake, jet, and vortex flows. Journal of Computational Physics, 5(3):359– 382, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900690.

Bottcher:1970:IIC

[Bot70] C. Bottcher. Integrals involving continuum wavefunctions I: One-center Coulomb integrals. Journal of Computational Physics, 6(2):237–246, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900239.

Boyer:1974:CGT

[Boy74] L. L. Boyer. Computerized group theory for lattice dynamical problems. Journal of Computational Physics, 16(2):167–185, October 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901119.

Brackbill:1973:IAL

[BP73] J. U. Brackbill and W. E. Pracht. An implicit, almost-Lagrangian algorithm for magnetohydrodynamics. Journal of Computational Physics, 13(4):455–482, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900284.

Bardo:1971:NAE

[BR71a] Richard D. Bardo and Klaus Ruedenberg. Numerical analysis and evaluation of normalized repeated integrals of the error function and related functions. Journal of Computational Physics, 8(1):167– 174, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900441.

Blottner:1971:NMS

[BR71b] F. G. Blottner and P. J. Roache. Nonuniform mesh systems. Journal of Computational Physics, 8(3):498–499, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900283.

Bauer:1974:NST

[BR74] Frances Bauer and Edward L. Reiss. On the numerical solution of two-dimensional elasticity problems. Journal of Computational Physics, 15(1):21–33, May 1974. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900667.

Brahic:1976:NSS

[Bra76] André Brahic. Numerical simulation of a system of colliding bodies in a gravitational field. Journal of Computational Physics, 22(2):171–188, October 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900747.

Briley:1974:NMP

[Bri74] W. Roger Briley. Numerical method for predicting threedimensional steady viscous flow in ducts. Journal of Computational Physics, 14(1):8–28, January 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900023.

Brown:1979:FEM

[Bro79] Robert A. Brown. Finite-element methods for the calculation of capillary surfaces. Journal of Computational Physics, 33(2):217– 235, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900172.

Brunner:1973:NSC

[Bru73] H. Brunner. The numerical solution of a class of Abel integral equations by piecewise polynomials. Journal of Computational Physics, 12(3):412–416, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901629.

Brumer:1974:SCN

[Bru74] Paul Brumer. Stability concepts in the numerical solution of classical atomic and molecular scattering problems. Journal of Computational Physics, 14(4):391–419, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900217.

Bender:1970:IPC

[BS70] C. F. Bender and I. Shavitt. An iterative procedure for the calculation of the lowest real eigenvalue and eigenvector of a nonsymmetric matrix. Journal of Computational Physics, 6(1):146– 149, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900148.

Bentley:1973:TCC

[BS73] John Bentley and Robert F. Stewart. Two-center calculations for x-ray scattering. Journal of Computational Physics, 11(1):127– 145, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901514.

Baker:1979:UFE

[BS79] A. J. Baker and M. O. Soliman. Utility of a finite element solution algorithm for initial-value problems. Journal of Computational Physics, 32(3):289–324, September 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901487.

Burridge:1979:FPS

[BT79] David M. Burridge and Clive Temperton. A fast Poisson-solver for large grids. Journal of Computational Physics, 30(1):145– 148, January 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900937.

Book:1979:FCE

[BTS79] D. L. Book, P. J. Turchi, and D. L. Stein. A finite-circuit-element code for modeling the compression of a gyrating chargedparticle beam. Journal of Computational Physics, 33(2):271–289, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900214.

Budgor:1979:MDA

[Bud79] Aaron B. Budgor. The method of differential areas for computing crystal symmetry independent density of states spectra. Journal of Computational Physics, 32(2):254–262, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901323.

Buneman:1971:AIF

[Bun71] O. Buneman. Analytic inversion of the five-point Poisson operator. Journal of Computational Physics, 8(3):500–505, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900295.

Buneman:1973:AIN

[Bun73a] O. Buneman. Analytic inversion of nine-point Poisson operator. Journal of Computational Physics, 11(3):447–448, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173900843.

Buneman:1973:ILP

[Bun73b] O. Buneman. Inversion of Lewis’ Poisson operator. Journal of Computational Physics, 11(3):307–314, March 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900740.

Buneman:1973:IHL

[Bun73c] O. Buneman. Inversion of the Helmholtz (or Laplace–Poisson) operator for slab geometry. Journal of Computational Physics, 12(1): 124–130, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901733.

Buneman:1973:SRF

[Bun73d] O. Buneman. Subgrid resolution of flow and force fields. Journal of Computational Physics, 11(2):250–268, February 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900065.

Buneman:1978:DOG

[Bun78] Oscar Buneman. Diagnosing oscillatory growth or decay. Journal of Computational Physics, 29(2):295–296, November 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178901602.

Burrell:1978:NTC

[Bur78] Keith H. Burrell. NEUCG: A transport code for hydrogen atoms in cylindrical hydrogenic plasmas. Journal of Computational Physics, 27(1):88–102, April 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917890027X.

Bushard:1975:MNA

[Bus75] Louis B. Bushard. Mathematical and numerical analysis of a differential equation model for a gas glow discharge. Journal of Computational Physics, 18(4):360–375, August 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900923.

Berezin:1976:FIA

[BV76] Yu. Berezin and V. A. Vshivkov. On the firehose instability of Alfvén waves. Journal of Computational Physics, 20(1):81– 96, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176901030.

Beam:1976:IFD

[BW76] Richard M. Beam and R. F. Warming. An implicit finitedifference algorithm for hyperbolic systems in conservation-law form. Journal of Computational Physics, 22(1):87–110, September 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176901108.

Braaten:1979:ILD

[BW79] Eric Braaten and George Weller. An improved low-discrepancy sequence for multidimensional quasi-Monte Carlo integration. Journal of Computational Physics, 33(2):249–258, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900196.

Belyaev:1976:SEP

[BWSE76] V. B. Belyaev, J. Wrzecionko, M. I. Sakvarelidze, and A. T. Eliava. Separable expansion of potential in multidimensional space. Three-body problem. Journal of Computational Physics, 20(4): 495–501, April 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900966.

Camp:1971:KFS

[Cam71] Jerome A. Camp. Kalman filtering simulation via numerical solution of the associated matrix differential equations. Journal of Computational Physics, 8(2):175–196, October 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999171900015.

Campbell:1972:CCF

[Cam72] J. A. Campbell. Computation of a class of functions useful in the phase-integral approximation. I. Results. Journal of Computational Physics, 10(2):308–315, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917290068X.

Canosa:1971:NSM

[Can71] José Canosa. Numerical solution of Mathieu’s equation. Journal of Computational Physics, 7(2):255–272, April 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917190088X.

Canosa:1973:NSL

[Can73] José Canosa. Numerical solution of Landau’s dispersion equation. Journal of Computational Physics, 13(1):158–160, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901319.

Cavendish:1975:LMR

[Cav75] James C. Cavendish. Local mesh refinement using rectangular blended finite elements. Journal of Computational Physics, 19(2):211–228, October 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900881.

Crowley:1971:TPB

[CB71] B. K. Crowley and L. K. Barr. TENSOR–PUFL: Boundary condition linked codes. Journal of Computational Physics, 7(1):167– 171, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900568.

Chorin:1973:DVS

[CB73] Alexandre Joel Chorin and Peter S. Bernard. Discretization of a vortex sheet, with an example of roll-up. Journal of Computational Physics, 13(3):423–429, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900454.

Chang:1970:PDS

[CC70a] J. S. Chang and G. Cooper. A practical difference scheme for Fokker–Planck equations. Journal of Computational Physics, 6(1): 1–16, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917090001X.

Collins:1970:PSW

[CC70b] Richard Collins and Hsiang-Teh Chen. Propagation of a shock wave of arbitrary strength in two half planes containing a free surface. Journal of Computational Physics, 5(3):415–442, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900720.

Chen:1971:SWP

[CC71] H.-T Chen and R. Collins. Shock wave propagation past an ocean surface. Journal of Computational Physics, 7(1):89–101, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900519.

Cambiaggio:1979:FDA

[CC79a] E. Cambiaggio and F. Cuozzo. Finite difference analysis of surface acoustic wave propagation and scattering in piezoelectric crystals. Journal of Computational Physics, 33(2):153–168, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900135.

Craxton:1979:SRT

[CC79b] R. S. Craxton and R. L. Mc Crory. A simple rezoning technique for use with the flux-corrected transport algorithm. Journal of Computational Physics, 33(3):432–440, December 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999179901694.

Car:1978:HRE

[CCQ78] R. Car, G. Ciucci, and L. Quartapelle. High-resolution and efficient calculation of photoelectric spectra of crystals. Journal of Computational Physics, 27(2):272–283, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900098.

Cavendish:1972:CCN

[CCV72] J. C. Cavendish, W. E. Culham, and R. S. Varga. A comparison of Crank–Nicolson and Chebyshev rational methods for numerically solving linear parabolic equations. Journal of Computational Physics, 10(2):354–368, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900721.

Crowder:1971:EUN

[CD71a] H. J. Crowder and C. Dalton. Errors in the use of nonuniform mesh systems. Journal of Computational Physics, 7(1):32– 45, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900477.

Crowder:1971:SPF

[CD71b] H. J. Crowder and C. Dalton. On the stability of Poiseuille flow in a pipe. Journal of Computational Physics, 7(1):12–31, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900465.

Carey:1978:MCP

[CD78a] David C. Carey and Daniel Druard. Monte Carlo phase space with limited transverse momentum. Journal of Computational Physics, 28(3):327–356, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900578.

Chang:1978:NSC

[CD78b] S. H. Chang and J. T. Day. On the numerical solution of a certain nonlinear integro-differential equation. Journal of Computational Physics, 26(2):162–168, February 1978. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900888.

Cebeci:1979:LBL

[Ceb79] Tuncer Cebeci. The laminar boundary layer on a circular cylinder started impulsively from rest. Journal of Computational Physics, 31(2):153–172, May 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900688.

Celmins:1976:CEN

[Cel76] Aivars Celmins. Comment on “The equivalence and numerical solution of regression optimal policy and boundary value problems involving differential equations”. Journal of Computational Physics, 21(2):243–244, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900152.

See [WH75].

Casati:1976:CSE

[CF76] Giulio Casati and Joseph Ford. Computer study of ergodicity and mixing in a two-particle, hard point gas system. Journal of Computational Physics, 20(1):97–109, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176901042.

Cox:1977:CRL

[CF77] S. B. Cox and J. L. Fry. Convergence of reciprocal-lattice sums: Exact formulas for the Ewald method. Journal of Computational Physics, 23(1):42–52, January 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900869.

Cloutman:1978:ACM

[CF78a] L. D. Cloutman and L. W. Fullerton. Automated computation of modified equations. Journal of Computational Physics, 29(1):141– 144, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901158.

Crank:1978:NSE

[CF78b] J. Crank and R. M. Furzeland. The numerical solution of elliptic and parabolic partial differential equations with boundary singularities. Journal of Computational Physics, 26(3):285– 296, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900712.

Cayford:1974:FDNa

[CFU74] J. K. Cayford, W. R. Fimple, and D. G. Unger. A finite difference Newton–Raphson solution of the atomic Hartree–Fock problem. Journal of Computational Physics, 15(1):81–97, May 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900709.

Cayford:1974:FDNb

[CFUW74] J. K. Cayford, W. R. Fimple, D. G. Unger, and S. P. White. A finite-difference Newton–Raphson solution of the two-center electronic Schrödinger equation. Journal of Computational Physics, 16(3):259–270, November 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900941.

Canosa:1970:NMS

[CG70] José Canosa and Roberto Gomes De Oliveira. A new method for the solution of the Schrödinger equation. Journal of Computational Physics, 5(2):188–207, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900598.

Colomhant:1976:SWP

[CG76] D. G. Colomhant and J. H. Gardner. Shock wave propagation in an inhomogeneous medium using finite differences. Journal of Computational Physics, 22(3):389–395, November 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900565.

Canosa:1977:KVB

[CG77] José Canosa and Jenö Gazdag. The Korteweg–de Vries–Burgers equation. Journal of Computational Physics, 23(4):393–403, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900705.

Canosa:1974:RIS

[CGF74] José Canosa, Jenö Gazdag, and J. E. Fromm. The recurrence of the initial state in the numerical solution of the Vlasov equation. Journal of Computational Physics, 15(1):34–45, May 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900679.

Chance:1978:CNS

[CGG^{+}78] M. S. Chance, J. M. Greene, R. C. Grimm, J. L. Johnson, J. Manickam, W. Kerner, D. Berger, L. C. Bernard, R. Gruber, and F. Troyon. Comparative numerical studies of ideal magnetohydrodynamic instabilities. Journal of Computational Physics, 28 (1):1–13, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900438.

Chapin:1970:UMF

[CGN70] C. E. Chapin, R. Goulard, and H. F. Nelson. The use of multiplying factors in radiation transfer iterations. Journal of Computational Physics, 5(3):455–462, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900744.

Chisholm:1972:ACS

[CGR72] J. S. R. Chisholm, A. Genz, and Glenys E. Rowlands. Accelerated convergence of sequences of quadrature approximations. Journal of Computational Physics, 10(2):284–307, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900678.

Campbell:1970:SAF

[CH70] J. A. Campbell and Anthony C. Hearn. Symbolic analysis of Feynman diagrams by computer. Journal of Computational Physics, 5(2):280–327, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900641.

Cline:1973:ANC

[CH73] Michael C. Cline and Toe D. Hoffman. The analysis of nonequilibrium, chemically reacting, supersonic flow in three dimensions using a bicharacteristics method. Journal of Computational Physics, 12(1):1–23, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901666.

Chamayou:1973:VFI

[Cha73] J. M. F. Chamayou. Volterra’s functional integral equations of the statistical theory of damage. Journal of Computational Physics, 13(1):70–93, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901265.

Chan:1974:DSS

[Cha74] Robert K.-C Chan. A discretized solution for the solitary wave. Journal of Computational Physics, 16(1):32–48, September 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174901028.

Chan:1975:GAL

[Cha75] Robert K.-C. Chan. A generalized arbitrary Lagrangian–Eulerian method for incompressible flows with sharp interfaces. Journal of Computational Physics, 17(3):311–331, March 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900558.

Chan:1976:SOT

[Cha76a] Robert K.-C Chan. A second-order, time integration scheme for calculating stratified incompressible flows. Journal of Computational Physics, 22(1):74–86, September 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176901091.

Chang:1976:AET

[Cha76b] Shih-Yung Chang. An algorithm for evaluation of two-electron integrals by numerical integration method. Journal of Computational Physics, 20(2):243–247, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900681.

Chen:1970:CGI

[Che70] Reuven Chen. On the computation of the generalized integral in glow curve theory. Journal of Computational Physics, 6(2):314– 316, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900276.

Chen:1971:RTA

[Che71a] Reuven Chen. On the remainder of truncated asymptotic series. Journal of Computational Physics, 8(1):156–161, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900428.

Chesnut:1971:MCS

[Che71b] Dwayne A. Chesnut. A Monte Carlo study of the thermodynamic properties of “hard hexagons” on the triangular lattice. Journal of Computational Physics, 7(3):409–434, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171901021.

Chesnut:1971:RBH

[Che71c] Dwayne A. Chesnut. The relation between hard-core lattice gases and hard discs. Journal of Computational Physics, 7(3):435– 444, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171901033.

Chen:1973:NBC

[Che73] J. H. Chen. Numerical boundary conditions and computational modes. Journal of Computational Physics, 13(4):522–535, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173900314.

Cheng:1977:IVE

[Che77] C. Z. Cheng. The integration of the Vlasov equation for a magnetized plasma. Journal of Computational Physics, 24(4):348– 360, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900274.

Chiu:1970:DMM

[Chi70] S. S. Chiu. Difference method for multiple reflection of elastic stress waves. Journal of Computational Physics, 6(1):17– 28, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900021.

Chisholm:1971:AAE

[Chi71] J. S. R. Chisholm. Algorithms for the γ-algebra of electromagnetic form factors. Journal of Computational Physics, 8(1):1– 18, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900313.

Chin:1974:NAL

[Chi74] Raymond C. Y. Chin. Notes on the article “Local errors of difference approximations to hyperbolic equations”. Journal of Computational Physics, 16(3):311–313, November 1974. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999174900990. See [OJ74].

Chin:1975:DGP

[Chi75] Raymond C. Y. Chin. Dispersion and Gibbs phenomenon associated with difference approximations to initial boundary-value problems for hyperbolic equations. Journal of Computational Physics, 18(3):233–247, July 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900017.

Chien:1976:GFD

[Chi76] John C. Chien. A general finite-difference formulation with application to Navier–Stokes equations. Journal of Computational Physics, 20(3):268–278, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900814.

Choong:1971:MMC

[Cho71a] Philip T. Choong. Modified Monte Carlo technique for clausing factor calculations. Journal of Computational Physics, 7(2):358– 360, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900969.

Chorin:1971:HEM

[Cho71b] Alexandre Joel Chorin. Hermite expansions in Monte–Carlo computation. Journal of Computational Physics, 8(3):472–482, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900258.

Chow:1973:SCF

[Cho73] T. L. Chow. Solution by characteristics at fixed time interval of the equations of one dimensional unsteady flow. Journal of Computational Physics, 12(2):153–170, February 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999173800075.

Chorin:1976:RCS

[Cho76] Alexandre Joel Chorin. Random choice solution of hyperbolic systems. Journal of Computational Physics, 22(4):517–533, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900474.

Chorin:1977:RCM

[Cho77] Alexandre Joel Chorin. Random choice methods with applications to reacting gas flow. Journal of Computational Physics, 25(3):253– 272, November 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177901012.

Chorin:1978:NCP

[Cho78a] Alexandre J. Chorin. Note: A comment on the paper “The calculation of large Reynolds number flow using discrete vortices with random walk” by F. Milinazzo and P. G. Saffman. Journal of Computational Physics, 26(3):453–454, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900827.

See [MS77] and reply [Saf78].

Chorin:1978:VSA

[Cho78b] Alexandre Joel Chorin. Vortex sheet approximation of boundary layers. Journal of Computational Physics, 27(3):428–442, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900190.

Christiansen:1973:NSH

[Chr73] I. P. Christiansen. Numerical simulation of hydrodynamics by the method of point vortices. Journal of Computational Physics, 13(3):363–379, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900429.

Carter:1978:AMC

[CHS78] L. L. Carter, H. G. Horak, and M. T. Sandford II. An adjoint Monte Carlo treatment of the equations of radiative transfer for polarized light. Journal of Computational Physics, 26(2):119– 138, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900852.

Chushkin:1970:SFA

[Chu70] P. I. Chushkin. Supersonic flows about conical bodies. Journal of Computational Physics, 5(3):572–586, June 1970. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999170900811.

Chu:1971:DGF

[Chu71] Wen-Hwa Chu. Development of a general finite difference approximation for a general domain part I: Machine transformation. Journal of Computational Physics, 8(3):392–408, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900192.

Cohen:1974:NTS

[CI74] H. Cohen and J. Ickovic. Numerical treatment of singular integral equations. Journal of Computational Physics, 16(4):371– 382, December 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900461.

Campbell:1974:CCF

[CJ74] J. A. Campbell and William H. Jefferys. Computation of a class of functions useful in the phase-integral approximation II. Performance of TRIGMAN, SYMBAL, and LISP. Journal of Computational Physics, 15(3):413–420, July 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917490120X.

Cebeci:1971:SPS

[CK71] Tuncer Cebeci and Herbert B. Keller. Shooting and parallel shooting methods for solving the Falkner–Skan boundary-layer equation. Journal of Computational Physics, 7(2):289–300, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900908.

Cheng:1976:IVE

[CK76] C. Z. Cheng and Georg Knorr. The integration of the Vlasov equation in configuration space. Journal of Computational Physics, 22(3):330–351, November 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917690053X.

Cebeci:1979:SBL

[CKW79] T. Cebeci, H. B. Keller, and P. G. Williams. Separating boundarylayer flow calculations. Journal of Computational Physics, 31(3): 363–378, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900524.

Clauser:1973:ECN

[Cla73] M. J. Clauser. Ellipsoidal coordinates — a natural coordinate system for calculations of laser irradiations of slabs. Journal of Computational Physics, 12(3):369–388, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901599.

Clark:1977:SSD

[Cla77] Terry L. Clark. A small-scale dynamic model using a terrainfollowing coordinate transformation. Journal of Computational Physics, 24(2):186–215, June 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900572.

Chen:1974:RGE

[CLB74] Liu Chen, A. Bruce Langdon, and C. K. Birdsall. Reduction of the grid effects in simulation plasmas. Journal of Computational Physics, 14(2):200–222, February 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917490014X.

Cukier:1978:NSA

[CLS78] R. I. Cukier, H. B. Levine, and K. E. Shuler. Nonlinear sensitivity analysis of multiparameter model systems. Journal of Computational Physics, 26(1):1–42, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900979.

Ciment:1978:OCI

[CLW78] Melvyn Ciment, Stephen H. Leventhal, and Bernard C. Weinberg. The operator compact implicit method for parabolic equations. Journal of Computational Physics, 28(2):135–166, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900311.

Campbell:1973:CCR

[CM73] J. B. Campbell and D. Maxow. Calculation of the capacitance of a ring capacitor of rectangular cross section with no insulating gaps. Journal of Computational Physics, 12(1):137– 142, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901757.

Cowan:1974:SSH

[CM74] Robert D. Cowan and Joseph B. Mann. Stabilization of solution of the Hartree–Fock equations. Journal of Computational Physics, 16(2):160–166, October 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174901107.

Campbell:1976:ECD

[CM76a] Edwin S. Campbell and Mihaly Mezei. Efficient construction of directional derivatives of a function of a vector magnitude and Maxwell’s invariant multipole form. Journal of Computational Physics, 21(1):114–122, May 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917690022X.

Cunningham:1976:CTM

[CM76b] G. W. Cunningham and P. H. E. Meuer. A comparison of two Monte Carlo methods for computations in statistical mechanics. Journal of Computational Physics, 20(1):50–63, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176901017.

Cordey:1978:NEM

[CMM^{+}78] J. G. Cordey, K. D. Marx, M. G. McCoy, A. A. Mirin, and M. E. Rensink. A new expansion method for computing σ̄ν for reactant distribution functions. Journal of Computational Physics, 28(1): 115–121, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900517.

Chiou:1975:STN

[CN75] J. P. Chiou and T. Y. Na. On the solution of Troesch’s nonlinear two-point boundary value problem using an initial value method. Journal of Computational Physics, 19(3):311–316, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900807.

Chen:1975:TPS

[CO75] Liu Chen and Hideo Okuda. Theory of plasma simulation using multipole-expansion scheme. Journal of Computational Physics, 19(4):339–352, December 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900674.

Cheng:1977:NTD

[CO77] C. Z. Cheng and H. Okuda. New three-dimensional simulation models for cylindrical and toroidal plasmas. Journal of Computational Physics, 25(2):133–150, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900171.

Cohen:1976:SPA

[Coh76] I. Cohen. Slowly pulsating axisymmetric systems in general relativity. Journal of Computational Physics, 22(3):396–402, November 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900577.

Cohen:1978:ANS

[Coh78] Harold Cohen. Accurate numerical solutions of integral equations with kernels containing poles. Journal of Computational Physics, 26(3):257–276, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900694.

Coldwell:1974:CDS

[Col74] Robert Lynn Coldwell. Correlational defects in the standard IBM 360 random number generator and the classical ideal gas correlation function. Journal of Computational Physics, 14(2):223–226, February 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900151. See comment [Ama77].

Cornille:1974:NSC

[Cor74] P. Cornille. Numerical study of a cylindrical plasma expansion in a magnetic field. Journal of Computational Physics, 14(4):371– 390, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900205.

Corradi:1975:SAC

[Cor75] C. Corradi. On the solution of the Ambartsumian–Chandrasekhar equation by monotone iteration processes. Journal of Computational Physics, 17(4):440–445, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900480.

Calmet:1971:AER

[CP71] J. Calmet and M. Perrottet. An attempt to evaluate renormalized radiative corrections by computer. Journal of Computational Physics, 7(2):191–200, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900842.

Canosa:1973:PSS

[CP73a] José Canosa and H. R. Penafiel. Parallel shooting solution of the neutron transport equation in spherical geometry. Journal of Computational Physics, 13(3):380–397, November 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900430.

Chang:1973:EPV

[CP73b] T. N. Chang and Robert T. Poe. On the evaluation of the principal value integral in the scattering problems. Journal of Computational Physics, 12(4):557–560, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917390106X.

Chen:1974:MTE

[CP74] M-S Chen and R. F. Peierls. A mapping technique for efficient random event generation with constraints. Journal of Computational Physics, 16(3):195–220, November 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900916.

Carlson:1975:EGN

[CP75] Gerald L. Carlson and Lee G. Pedersen. Extension of GAUS-SIAN70 to any number of atomic orbitals. Journal of Computational Physics, 17(4):415–419, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900431.

Coldwell:1976:LSR

[CP76] R. L. Coldwell and M. A. Pokrant. Least square regression with multinomials of restricted amplitudes. Journal of Computational Physics, 20(3):365–371, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900887.

Cole:1979:EI

[CP79a] R. J. Cole and C. Pescatore. Evaluation of the integral ∫_{0}^{∞} t^{n} exp(−t^{2} − x/t) dt. Journal of Computational Physics, 32(2): 280–287, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901359.

Concus:1979:NSN

[CP79b] Paul Concus and Włodzimierz Proskurowski. Numerical solution of a nonlinear hyperbolic equation by the random choice method. Journal of Computational Physics, 30(2):153–166, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900962.

Christiansen:1975:NTR

[CR75a] J. P. Christiansen and K. V. Roberts. Numerical treatment of rapid equipartition rates. Journal of Computational Physics, 17(3):332– 339, March 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917590056X.

Crater:1975:EFD

[CR75b] H. W. Crater and G. W. Reddien. Extrapolation of finite difference approximations for bound state equations. Journal of Computational Physics, 19(2):236–240, October 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917590090X.

Christiansen:1978:EME

[CR78] J. P. Christiansen and K. V. Roberts. Exact magnetic energy conservation on a nonuniform mesh. Journal of Computational Physics, 28(2):279–286, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900396.

Crippen:1975:GOP

[Cri75] Gordon M. Crippen. Global optimization and polypeptide conformation. Journal of Computational Physics, 18(2):224–231, June 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900303.

Crippen:1977:NAC

[Cri77] Gordon M. Crippen. A novel approach to calculation of conformation: Distance geometry. Journal of Computational Physics, 24(1): 96–107, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177901127.

Crippen:1978:NRC

[Cri78] G. M. Crippen. Note: Rapid calculation of coordinates from distance matrices. Journal of Computational Physics, 26(3):449– 452, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900815.

Crowley:1970:NMVa

[Cro70a] W. P. Crowley. A numerical model for viscous, free-surface, barotropic wind driven ocean circulations. Journal of Computational Physics, 5(1):139–168, February 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900574.

Crowley:1970:NMVb

[Cro70b] W. P. Crowley. A numerical model for viscous, nondivergent, barotropic, wind-driven, ocean circulations. Journal of Computational Physics, 6(2):183–199, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900203.

Chan:1970:CSF

[CS70] Robert K.-C Chan and Robert L. Street. A computer study of finite-amplitude water waves. Journal of Computational Physics, 6(1):68–94, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900057.

Ciment:1973:MRP

[CS73a] Melvyn Ciment and Roland A. Sweet. Mesh refinements for parabolic equations. Journal of Computational Physics, 12(4):513– 525, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901022.

Crippen:1973:MPE

[CS73b] G. M. Crippen and H. A. Scheraga. Minimization of polypeptide energy XII. The methods of partial energies and cubic subdivision. Journal of Computational Physics, 12(4):491–497, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901009.

Chu:1974:BCF

[CS74] C. K. Chu and Aron Sereny. Boundary conditions in finite difference fluid dynamic codes. Journal of Computational Physics, 15(4): 476–491, August 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900746. See note [Sun75].

Chang:1976:CAD

[CS76] Sin-Chung Chang and J. Kenneth Shultis. Calculation of the asymptotic diffusion lengths for highly anisotropic particle transport. Journal of Computational Physics, 20(3):336–345, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900863.

Coope:1977:NAD

[CS77] J. A. R. Coope and D. W. Sabo. A new approach to the determination of several eigenvectors of a large Hermitian matrix. Journal of Computational Physics, 23(4):404–424, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900717.

Cohen:1978:SPS

[CS78] I. Cohen and S. Yu Slavyanov. Smooth perturbations of the Schrödinger equation related to the charmonium models. Journal of Computational Physics, 29(2):289–294, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901596.

Chen:1977:NAR

[CT77] Y. M. Chen and D. S. Tsien. A numerical algorithm for remote sensing of density profiles of a simple ocean model by acoustic pulses. Journal of Computational Physics, 25(4):366–385, December 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900043.

Cummins:1975:CPP

[Cum75] W. F. Cummins. Computer processing of plasma interferometer signals. Journal of Computational Physics, 18(1):106–113, May 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175901059.

Curl:1970:DLS

[Cur70] R. F. Curl, Jr. Diagnostic least squares. Journal of Computational Physics, 6(3):367–377, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900367.

Chu:1976:IEM

[CW76] C.-M Chu and Herschel Weil. Integral equation method for scattering and absorption of electromagnetic radiation by thin lossy dielectric discs. Journal of Computational Physics, 22(1):111–124, September 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917690111X.

Came:1979:CRM

[CZ79] A. O. Came and S. I. Zanette. On the calculation of rotation matrices. Journal of Computational Physics, 33(3):441–443, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901700.

Caride:1978:RRG

[CZF78] A. O. Caride, S. I. Zanette, and H. A. Farach. Recursion relations for the group SU (2). Journal of Computational Physics, 29(2):278– 282, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901572.

Dalrymple:1977:NMP

[Dal77] Robert A. Dalrymple. A numerical model for periodic finite amplitude waves on a rotational fluid. Journal of Computational Physics, 24(1):29–42, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177901085.

Martins:1977:RFP

[dAPM77] Pedro de A. P. Martins. Recurrence formulas for phases and amplitudes of spherical function of a free wave. Journal of Computational Physics, 25(2):182–193, October 1977. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900201.

Martins:1979:DLO

[dAPM79] Pedro de A. P. Martins. Determination of large-order spherical Coulomb functions with an argument lying between the origin and the common point of inflection. Journal of Computational Physics, 31(2):293–299, May 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900743.

Davies:1973:IBV

[Dav73] Huw C. Davies. On the initial-boundary value problem of some geophysical fluid flows. Journal of Computational Physics, 13(3):398– 422, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900442.

Davidson:1975:NIC

[Dav75] Ernest R. Davidson. Note: The iterative calculation of a few of the lowest eigenvalues and corresponding eigenvectors of large real-symmetric matrices. Journal of Computational Physics, 17(1):87–94, January 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://

comptop.stanford.edu/u/references/da.pdf;   http://www.
sciencedirect.com/science/article/pii/0021999175900650.

This paper generalizes the work of Lanczos [?] by a modification that improves the convergence. See also later generalizations [?, ?, ?].

Davey:1977:NSD

[Dav77] A. Davey. On the numerical solution of difficult eigenvalue problems. Journal of Computational Physics, 24(3):331–338, July 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900419.

Davies:1978:UGM

[Dav78] A. M. Davies. The use of the Galerkin method with a basis of B-splines for the solution of the one-dimensional primitive equations. Journal of Computational Physics, 27(1):123– 137, April 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900293.

Davey:1979:RSR

[Dav79a] A. Davey. On the removal of the singularities from the Riccati method. Journal of Computational Physics, 30(1):137–144, January 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900925.

Davidson:1979:PAD

[Dav79b] G. T. Davidson. Pitch angle diffusion of trapped particles in the presence of a loss cone: Calculating the distribution of particles precipitating from the Earth’s radiation belts. Journal of Computational Physics, 31(3):301–312, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900494.

Denny:1974:CGF

[DC74] V. E. Denny and R. M. Clever. Comparisons of Galerkin and finite difference methods for solving highly nonlinear thermally driven flows. Journal of Computational Physics, 16(3):271–284, November 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900953.

Decyk:1979:CMB

[DD79] Viktor K. Decyk and John M. Dawson. Computer model for bounded plasma. Journal of Computational Physics, 30(3):407– 427, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901232.

deAPMartins:1977:DSB

[de 77] Pedro de A. P. Martins. Determination of spherical Bessel functions of an order larger than the argument. Journal of Computational Physics, 25(2):194–198, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900213.

Drake:1970:ALS

[DE70] Ronald L. Drake and Michael B. Ellingson. The application of a local similarity concept in solving the radial flow problem. Journal of Computational Physics, 6(2):200–218, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900215.

Deardorff:1971:MSS

[Dea71] J. W. Deardorff. On the magnitude of the subgrid scale eddy coefficient. Journal of Computational Physics, 7(1):120–133, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900532.

Delic:1974:LSQ

[Del74] George Delic. The Legendre series and a quadrature formula for its coefficients. Journal of Computational Physics, 14(3):254– 268, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900527.

Denavit:1972:NSP

[Den72] J. Denavit. Numerical simulation of plasmas with periodic smoothing in phase space. Journal of Computational Physics, 9(1):75– 98, February 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917290037X.

Denavit:1974:DPE

[Den74] J. Denavit. Discrete particle effects in Whistler simulation. Journal of Computational Physics, 15(4):449–475, August 1974. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999174900734.

Detyna:1979:PCR

[Det79] E. Detyna. Point cyclic reductions for elliptic boundary-value problems. I. The constant-coefficient case. Journal of Computational Physics, 33(2):204–216, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900160.

Deville:1974:NEM

[Dev74] Michel O. Deville. Numerical experiments on the MAC code for a slow flow. Journal of Computational Physics, 15(3):362– 374, July 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917490117X.

Dickinson:1976:SAO

[DG76] Robert P. Dickinson and Robert J. Gelinas. Sensitivity analysis of ordinary differential equation systems — a direct method. Journal of Computational Physics, 21(2):123–143, June 1976. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999176900073.

Degani:1977:TDL

[DG77] David Degani and Cham Gutfinger. Three-dimensional low Reynolds number flows with a free surface. Journal of Computational Physics, 24(1):108–114, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177901139.

Dewar:1975:NSM

[DGGJ75] R. L. Dewar, J. M. Greene, R. C. Grimm, and J. L. Johnson. Numerical study of the magnetohydrodynamic spectra in tokamaks using Galerkin’s method. Journal of Computational Physics, 18(2): 132–153, June 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917590025X.

Doering:1973:NAO

[DGH73] R. R. Doering, A. I. Galonsky, and R. A. Hinrichs. The numerical accuracy of ^{3} he optical-model calculations at 70 MeV. Journal of Computational Physics, 12(4):498–512, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901010.

Dickinson:1972:DAP

[Dic72] A. S. Dickinson. Determination of accurate potential strengths to yield specified eigenvalues of the radial Schrödinger equation. Journal of Computational Physics, 10(1):162–166, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900976.

Dickinson:1973:CFD

[Dic73] A. S. Dickinson. Comparison of the finite difference boundary value method and the Cooley–Cashion–Zare method for solving one-dimensional eigenvalue equations. Journal of Computational Physics, 11(3):458–462, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900879.

Dennis:1979:FDM

[DIC79] S. C. R. Dennis, D. B. Ingham, and R. N. Cook. Finitedifference methods for calculating steady incompressible flows in three dimensions. Journal of Computational Physics, 33(3):325– 339, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901591.

Dibiase:1977:NMR

[DK77] James A. Dibiase and John Killeen. A numerical model for resistive magnetohydrodynamic instabilities. Journal of Computational Physics, 24(2):158–185, June 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900560.

Denny:1972:NMS

[DL72] V. E. Denny and R. B. Landis. A new method for solving two-point boundary value problems using optimal node distribution. Journal of Computational Physics, 9(1):120–137, February 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900393.

Darby:1973:NSN

[DM73] M. I. Darby and N. Morton. Numerical solutions for the nonlinear penetration of magnetic flux into type II superconductors. Journal of Computational Physics, 13(1):35–44, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917390123X.

Davies:1979:NIL

[DM79] Brian Davies and Brian Martin. Numerical inversion of the Laplace transform: a survey and comparison of methods. Journal of Computational Physics, 33(1):1–32, October 1979. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900251.

Dietrich:1975:OBI

[DMWV75] D. Dietrich, B. E. McDonald, and A. Warn-Varnas. Optimized block-implicit relaxation. Journal of Computational Physics, 18(4): 421–439, August 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900959.

Dowsing:1970:UGS

[Dow70] R. D. Dowsing. Use of grid search techniques to extend the use of a least squares program for analysis of electron spin resonance spectra. Journal of Computational Physics, 6(2):326– 328, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900306.

deRivas:1972:UNG

[dR72] Eugenia Kálnay de Rivas. On the use of nonuniform grids in finitedifference equations. Journal of Computational Physics, 10(2):202– 210, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900605.

Dukowicz:1979:TVM

[DR79] J. K. Dukowicz and J. D. Ramshaw. Tensor viscosity method for convection in numerical fluid dynamics. Journal of Computational Physics, 32(1):71–79, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901426.

DeSanto:1974:NSS

[DS74] John A. DeSanto and Oved Shisha. Numerical solution of a singular integral equation in random rough surface scattering theory. Journal of Computational Physics, 15(2):286–292, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900898.

Davies:1975:CCE

[DS75] K. T. R. Davies and A. J. Sierk. Calculation of Coulomb energies for uniform charge distributions of arbitrary shape. Journal of Computational Physics, 18(3):311–325, July 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900054.

Dennis:1978:CFB

[DS78] S. C. R. Dennis and S. N. Singh. Calculation of the flow between two rotating spheres by the method of series truncation. Journal of Computational Physics, 28(3):297–314, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900554.

Desideri:1977:RAM

[DT77] J.-A Désidéri and J. C. Tannehill. Over-relaxation applied to the MacCormack finite-difference scheme. Journal of Computational Physics, 23(3):313–326, March 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900973.

Dishon:1978:NIM

[DW78] Menachem Dishon and George H. Weiss. Numerical inversion of Mellin and two-sided Laplace transforms. Journal of Computational Physics, 28(1):129–132, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900530.

Easton:1972:HBC

[Eas72] C. R. Easton. Homogeneous boundary conditions for pressure in the Mac method. Journal of Computational Physics, 9(2):375– 379, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900307.

Eastwood:1975:OPM

[Eas75] J. W. Eastwood. Optimal particle-mesh algorithms. Journal of Computational Physics, 18(1):1–20, May 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900996.

Eastwood:1979:RSP

[EB79] J. W. Eastwood and D. R. K. Brownrigg. Remarks on the solution of Poisson’s equation for isolated systems. Journal of Computational Physics, 32(1):24–38, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901396.

Easton:1972:IVT

[EC72] C. R. Easton and Ivan Catton. Initial value techniques in freesurface hydrodynamics. Journal of Computational Physics, 9(3): 424–439, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900046.

Elbert:1974:EER

[ED74] Stephen T. Elbert and Ernest R. Davidson. Evaluation of electron repulsion integrals over Gaussian lobe basis functions. Journal of Computational Physics, 16(4):391–395, December 1974. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999174900485.

Edelson:1973:SDE

[Ede73] D. Edelson. On the solution of differential equations arising in chemical kinetics. Journal of Computational Physics, 11(3):455– 457, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900867.

Erber:1979:SRP

[EEJ79] T. Erber, P. Everett, and P. W. Johnson. The simulation of random processes on digital computers with Čebyšev mixing transformations. Journal of Computational Physics, 32(2):168– 211, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901281.

Engels:1977:EPS

[EF77] J. Engels and J. Fleischer. Evaluation of phase-space volumes and inclusive spectra in the uncorrelated jet model. Journal of Computational Physics, 23(2):200–217, February 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917790122X.

Engquist:1978:NSP

[EGV78] Björn Engquist, Bertil Gustafsson, and Joop Vreeburg. Numerical solution of a PDE system describing a catalytic converter. Journal of Computational Physics, 27(3):295–314, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900128.

Eilon:1972:NSC

[EGZ72] B. Eilon, D. Gottlieb, and G. Zwas. Numerical stabilizers and computing time for second-order accurate schemes. Journal of Computational Physics, 9(3):387–397, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900010.

Ewig:1973:MSP

[EH73] Carl S. Ewig and David O. Harris. Matrix solution of periodic Mathieu equations. Journal of Computational Physics, 11(4):606– 611, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917390140X.

Eastwood:1974:SFL

[EH74] J. W. Eastwood and R. W. Hockney. Shaping the force law in two-dimensional particle-mesh models. Journal of Computational Physics, 16(4):342–359, December 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900448.

Erdos:1970:NMS

[EHMM70] P. Erdős, S. B. Haley, J. T. Marti, and J. Mennig. A new method for the solution of the transport equation in slab geometry. Journal of Computational Physics, 6(1):29–55, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900033.

Einarsson:1976:URE

[Ein76] Bo Einarsson. Use of Richardson extrapolation for the numerical calculation of Fourier transforms. Journal of Computational Physics, 21(3):365–370, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917690036X.

Eiseman:1978:CSV

[Eis78] Peter R. Eiseman. A coordinate system for a viscous transonic cascade analysis. Journal of Computational Physics, 26(3):307– 338, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900736.

Eiseman:1979:MSM

[Eis79] Peter R. Eiseman. A multi-surface method of coordinate generation. Journal of Computational Physics, 33(1):118–150, October 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900317.

Emery:1973:CFH

[EJ73] Mark H. Emery and Glenn Joyce. A comparison of the Fourier– Hermite and Power transform solutions of the nonlinear Vlasov equation. Journal of Computational Physics, 11(4):493–506, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901344.

Esbjorn:1973:EFH

[EJK^{+}73] P. O. Esbjørn, E. J. Jensen, W. D. Kristensen, J. W. Martin, and L. B. Pedersen. Entropy of fcc and hcp lattices: Comparison of H theorem method with vibrational analysis. Journal of Computational Physics, 12(3):289–307, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901538.

Eilbeck:1975:NSR

[EM75] J. C. Eilbeck and G. R. McGuire. Numerical study of the regularized long-wave equation I: Numerical methods. Journal of Computational Physics, 19(1):43–57, September 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175901151.

Eilbeck:1977:NSR

[EM77] J. C. Eilbeck and G. R. McGuire. Numerical study of the regularized long-wave equation. II: Interaction of solitary waves. Journal of Computational Physics, 23(1):63–73, January 1977. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999177900882.

Empedocles:1970:GSE

[Emp70] Philip Empedocles. Ground-state eigenvectors for many configuration wavefunctions. Journal of Computational Physics, 6(1):113– 119, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900082.

Ekstrom:1974:AEE

[ER74] M. P. Ekstrom and R. L. Rhoads. On the application of eigenvector expansions to numerical deconvolution. Journal of Computational Physics, 14(4):319–340, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900163.

Evans:1979:ND

[Eva79] G. A. Evans. A numerical differentiator. Journal of Computational Physics, 32(2):247–253, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901311.

Fox:1976:TOH

[FB76] Herbert L. Fox and Richard Barakat. Trace orthogonal Hermitian basis matrices of arbitrary dimension. Journal of Computational Physics, 21(3):326–332, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900292.

Fritts:1979:LST

[FB79] M. J. Fritts and J. P. Boris. The Lagrangian solution of transient problems in hydrodynamics using a triangular mesh. Journal of Computational Physics, 31(2):173–215, May 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917990069X.

Fleck:1971:IMC

[FC71] J. A. Fleck, Jr. and J. D. Cummings, Jr. An implicit Monte Carlo scheme for calculating time and frequency dependent nonlinear radiation transport. Journal of Computational Physics, 8(3):313– 342, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900155.

Fripiat:1979:FRC

[FD79] J. G. Fripiat and J. Delhalle. Fourier representation of the Coulombic contributions to polymer chains. Journal of Computational Physics, 33(3):425–431, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901682.

Forester:1972:CML

[FE72] C. K. Forester and A. F. Emery. A computational method for low Mach number unsteady compressible free convective flows. Journal of Computational Physics, 10(3):487–502, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900496.

Feler:1974:CEL

[Fel74] M. Guy Feler. Calculation of eigenvectors of large matrices. Journal of Computational Physics, 14(4):341–349, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900175.

Fix:1973:USF

[FGW73] G. J. Fix, S. Gulati, and G. I. Wakoff. On the use of singular functions with finite element approximations. Journal of Computational Physics, 13(2):209–228, October 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900235.

Fox:1974:ACS

[FH74] John A. Fox and Anthony C. Hearn. Analytic computation of some integrals in fourth order quantum electrodynamics. Journal of Computational Physics, 14(3):301–317, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900552.

Fletcher:1975:IMI

[FH75] Clive A. J. Fletcher and Maurice Holt. An improvement to the method of integral relations. Journal of Computational Physics, 18(2):154–164, June 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900261.

Finney:1978:LRF

[Fin78] J. L. Finney. Long-range forces in molecular dynamics calculations on water. Journal of Computational Physics, 28(1):92– 102, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900499.

Finney:1979:PCV

[Fin79] J. L. Finney. A procedure for the construction of Voronoi polyhedra. Journal of Computational Physics, 32(1):137–143, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901463.

Fast:1971:DDS

[FJ71a] G. Fast and T. Janssen. Determination of n-dimensional space groups by means of an electronic computer. Journal of Computational Physics, 7(1):1–11, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900453.

Fosdick:1971:ESA

[FJ71b] Lloyd D. Fosdick and Robert C. Jacobson. Experiments on a superposition approximation for Wiener path integrals. Journal of Computational Physics, 7(1):157–166, February 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999171900556.

Fox:1977:CCH

[FK77] Kenneth Fox and Burton J. Krohn. Computation of cubic harmonics. Journal of Computational Physics, 25(4):386–408, December 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900055.

Fleischer:1973:PWA

[Fle73] J. Fleischer. Partial wave analysis of nucleon-nucleon Bethe– Salpeter equation on the computer. Journal of Computational Physics, 12(1):112–123, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901721.

Fleck:1974:CSM

[Fle74] J. A. Fleck, Jr. A cubic spline method for solving the wave equation of nonlinear optics. Journal of Computational Physics, 16(4):324– 341, December 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900436.

Fletcher:1979:PVF

[Fle79] Clive A. J. Fletcher. A primitive variable finite element formulation for inviscid, compressible flow. Journal of Computational Physics, 33(3):301–312, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901578.

Flodmark:1977:STC

[Flo77] Stig Flodmark. Symmetry transformation of crystal wave functions. Journal of Computational Physics, 25(3):314–318, November 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177901048.

Friedman:1971:EMC

[FLRZ71] Jerome H. Friedman, Gerald R. Lynch, Clifford G. Risk, and Thomas A. Zang, Jr. An efficient Monte Carlo event generation method for multiperipheral models. Journal of Computational Physics, 8(1):144–152, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900404.

Fanning:1973:SNP

[FM73] A. E. Fanning and T. J. Mueller. On the solution of a Neumann problem for an inhomogeneous Laplace equation. Journal of Computational Physics, 13(3):450–454, November 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900508.

Fix:1978:VMU

[FM78] George J. Fix and Samuel P. Marin. Variational methods for underwater acoustic problems. Journal of Computational Physics, 28(2): 253–270, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900372.

Fimple:1976:FDN

[FMW76] W. R. Fimple, B. J. McKenzie, and S. P. White. A finite-difference Newton–Raphson solution of the multiconfiguration Hartree–Fock problem. Journal of Computational Physics, 22(4):435–450, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900425.

Fox:1973:PAT

[FO73] Douglas G. Fox and Steven A. Orszag. Pseudospectral approximation to two-dimensional turbulence. Journal of Computational Physics, 11(4):612–619, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901411.

Foote:1973:NMS

[Foo73] G. Brant Foote. A numerical method for studying liquid drop behavior: Simple oscillation. Journal of Computational Physics, 11(4):507–530, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901356.

Forester:1973:MTC

[For73] C. K. Forester. A method of time-centering the Lagrangian Marker particle computation. Journal of Computational Physics, 12(2):269–274, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999173800154.

Forester:1977:HOM

[For77a] C. K. Forester. Higher order monotonic convective difference schemes. Journal of Computational Physics, 23(1):1–22, January 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900845.

Fornberg:1977:NSD

[For77b] Bengt Fornberg. A numerical study of 2-D turbulence. Journal of Computational Physics, 25(1):1–31, September 1977. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999177900237.

Foster:1973:NMR

[Fos73] P. M. Foster. A nonfunctional method for reducing cumulative or histogram data to a smooth distributional form or for constructing a smooth approximation to experimental data. Journal of Computational Physics, 12(1):143–149, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901769.

Fox:1977:RRR

[Fox77] Kenneth Fox. On recursion relations for rapid computation of rotation functions. Journal of Computational Physics, 24(4): 455, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900341.

Filippov:1976:NMS

[FPM76] A. T. Filippov, I. V. Puzynin, and Dmitrij P. Mavlo. Numerical methods for solving some singular boundary value problems of quantum mechanics and quantum field theory. I. Journal of Computational Physics, 22(2):150–170, October 1976. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999176900735.

Ferris-Prabhu:1973:NAC

[FPW73] A. V. Ferris-Prabhu and D. H. Withers. Numerical analytic continuation using Padé approximants. Journal of Computational Physics, 13(1):94–99, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901277.

Farmer:1977:MRA

[FR77] Otis A. Farmer and John D. Ramshaw. A method for rapid attainment of the steady state in two-dimensional time-marching supersonic flow calculations. Journal of Computational Physics, 24(1): 23–28, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177901073.

Freeman:1973:MCT

[Fre73] J. R. Freeman. A method for computing the transition from ambipolar to free diffusion in a decaying plasma. Journal of Computational Physics, 13(1):59–69, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901253.

Friedman:1971:REG

[Fri71] Jerome H. Friedman. Random event generation with preferred frequency distributions. Journal of Computational Physics, 7(2): 201–218, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900854.

Friar:1978:NRE

[Fri78] J. L. Friar. A note on the roundoff error in the Numerov algorithm. Journal of Computational Physics, 28(3):426–432, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900633.

FroeseFischer:1972:ISH

[Fro72] Charlotte Froese Fischer. Iterative solution of the Hartree–Fock equations with improved stability. Journal of Computational Physics, 10(2):211–223, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900617.

FroeseFischer:1973:SSE

[Fro73] Charlotte Froese Fischer. The solution of Schrödinger’s equation for two-electron systems by an MCHF procedure. Journal of Computational Physics, 13(4):502–521, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900302.

FroeseFischer:1978:NSG

[Fro78] Charlotte Froese Fischer. Numerical solution of general Hartree– Fock equations for atoms. Journal of Computational Physics, 27(2): 221–240, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900062.

Friedman:1978:FEM

[FRRT78] M. Friedman, Y. Rosenfeld, A. Rabinovitch, and R. Thieberger. Finite element method for solving the two-dimensional Schrödinger equation. Journal of Computational Physics, 26(2):169–180, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917890089X.

Fairweather:1979:NST

[FRSW79] Graeme Fairweather, Frank J. Rizzo, David J. Shippy, and Yensen S. Wu. On the numerical solution of two-dimensional potential problems by an improved boundary integral equation method. Journal of Computational Physics, 31(1):96–112, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900640.

Friedman:1979:AFE

[FRT79] M. Friedman, A. Rabinovitch, and R. Thieberger. Application of the finite-element method to the hydrogen atom in a box in an electric field. Journal of Computational Physics, 33(3):359– 368, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917990161X.

Froeschle:1973:NSS

[FS73] Claude Froeschle and Jean-Paul Scheidecker. Numerical study of the stochasticity of dynamical systems with more than two degrees of freedom. Journal of Computational Physics, 11(3):423– 439, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917390082X.

Freeman:1977:NMS

[FT77] John R. Freeman and Samuel L. Thompson. Numerical methods for studying compressed magnetic field generators. Journal of Computational Physics, 25(4):332–352, December 1977. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917790002X.

Fortin:1979:MFE

[FT79] M. Fortin and F. Thomasset. Mixed finite-element methods for incompressible flow problems. Journal of Computational Physics, 31(1):113–145, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900652.

Galloway:1972:FDA

[GA72] F. M. Galloway and R. J. Adler. On a finite-difference approach to turbulence problems. Journal of Computational Physics, 10(3):379–399, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900435.

Gary:1974:OTS

[Gar74] John M. Gary. On the optimal time step and computational efficiency of difference schemes for PDE. Journal of Computational Physics, 16(3):298–303, November 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900977.

Gary:1976:MLA

[Gar76] John Gary. The method of lines applied to a simple hyperbolic equation. Journal of Computational Physics, 22(2):131– 149, October 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900723.

Gary:1978:BCH

[Gar78] John Gary. On boundary conditions for hyperbolic difference schemes. Journal of Computational Physics, 26(3):339–351, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900748. See comment [Gus80].

Gaster:1978:SRE

[Gas78] M. Gaster. Series representation of the eigenvalues of the Orr–Sommerfeld equation. Journal of Computational Physics, 29(2):147–162, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901481.

Gatski:1975:NMO

[Gat75] Thomas B. Gatski. On the numerical modeling of an Oldroydtype constitutive equation. Journal of Computational Physics, 19(4):376–389, December 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900698.

Gazdag:1970:BCP

[Gaz70] Jeno Gazdag. Bubble-chamber picture processing through digital filtering. Journal of Computational Physics, 5(1):34–55, February 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900513.

Gazdag:1973:NCS

[Gaz73] Jenö Gazdag. Numerical convective schemes based on accurate computation of space derivatives. Journal of Computational Physics, 13(1):100–113, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901289.

Gazdag:1975:NSV

[Gaz75] Jeno Gazdag. Numerical solution of the Vlasov equation with the accurate space derivative method. Journal of Computational Physics, 19(1):77–89, September 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175901175.

Gazdag:1976:TDS

[Gaz76] Jenö Gazdag. Time-differencing schemes and transform methods. Journal of Computational Physics, 20(2):196–207, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900644.

Gal-Chen:1975:NSN

[GCS75a] Tzvi Gal-Chen and Richard C. J. Somerville. Numerical solution of the Navier–Stokes equations with topography. Journal of Computational Physics, 17(3):276–310, March 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900546.

Gal-Chen:1975:UCT

[GCS75b] Tzvi Gal-Chen and Richard C. J. Somerville. On the use of a coordinate transformation for the solution of the Navier–Stokes equations. Journal of Computational Physics, 17(2):209–228, February 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900376.

Godunov:1970:HEC

[GDZK70] S. K. Godunov, A. A. Deribas, A. V. Zabrodin, and N. S. Kozin. Hydrodynamic effects in colliding solids. Journal of Computational Physics, 5(3):517–539, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900781.

Gelinas:1972:SSK

[Gel72] Robert J. Gelinas. Stiff systems of kinetic equations — a practitioner’s view. Journal of Computational Physics, 9(2):222– 236, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900204.

Gregory:1978:LMR

[GFSW78] J. A. Gregory, D. Fishelov, B. Schiff, and J. R. Whiteman. Local mesh refinement with finite elements for elliptic problems. Journal of Computational Physics, 29(1):133–140, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901146.

Gary:1970:MMO

[GH70] John Gary and Richard Helgason. A matrix method for ordinary differential eigenvalue problems. Journal of Computational Physics, 5(2):169–187, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900586.

Guderley:1973:SFG

[GH73] Karl G. Guderley and Chen-Chi Hsu. A special form of Galerkin’s method applied to heat transfer in plane Couette–Poiseuille flows. Journal of Computational Physics, 11(1):70–89, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901484.

Gordon:1977:EMB

[GH77] William J. Gordon and Charles A. Hall. Exact matching of boundary conditions and incorporation of semiquantitative solution characteristics in initial approximations to boundary value problems. Journal of Computational Physics, 25(2):151–162, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900183.

Gilat:1972:AMC

[Gil72] G. Gilat. Analysis of methods for calculating spectral properties in solids. Journal of Computational Physics, 10(3):432–465, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900460.

Gillespie:1976:GMN

[Gil76] Daniel T. Gillespie. A general method for numerically simulating the stochastic time evolution of coupled chemical reactions. Journal of Computational Physics, 22(4):403–434, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900413.

Gillespie:1978:MCS

[Gil78] Daniel T. Gillespie. Monte Carlo simulation of random walks with residence time dependent transition probability rates. Journal of Computational Physics, 28(3):395–407, September 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178900608.

Grimm:1973:SAL

[GJ73] Ray C. Grimm and John L. Johnson. Stability analysis of low-pressure axisymmetric toroidal plasma simulations. Journal of Computational Physics, 11(4):591–605, April 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173901393.

Grimm:1975:FSL

[GJ75] R. C. Grimm and J. L. Johnson. Fluid simulation of a low-pressure axisymmetric toroidal plasma using natural coordinates. Journal of Computational Physics, 17(2):192–208, February 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900364.

Griffin:1979:CFD

[GJA79] Michael D. Griffin, Everett Jones, and John D. Anderson, Jr. A computational fluid dynamic technique valid at the centerline for non-axisymmetric problems in cylindrical coordinates. Journal of Computational Physics, 30(3):352–360, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901207.

Golub:1975:WPS

[GJY75] G. H. Golub, L. Jenning, and W. H. Yang. Waves in periodically structured media. Journal of Computational Physics, 17(4): 349–357, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917590039X.

Gustafsson:1979:BCT

[GK79] Bertil Gustafsson and Heinz-Otto Kreiss. Boundary conditions for time dependent problems with an artificial boundary. Journal of Computational Physics, 30(3):333–351, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901190.

Godfrey:1976:SLD

[GL76] Brendan B. Godfrey and A. Bruce Langdon. Stability of the Langdon–Dawson advective algorithm. Journal of Computational Physics, 20(2):251–255, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917690070X.

Gustafson:1977:NCI

[GL77] Sven-Åke Gustafson and Staffan Lindahl. Numerical computation of an integral appearing in the Fröman–Fröman phase-integral formula for calculation of quantal matrix elements without the use of wave functions. Journal of Computational Physics, 24(1): 81–95, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177901115.

Gatski:1978:SFN

[GL78] Thomas B. Gatski and John L. Lumley. Steady flow of a non-Newtonian fluid through a contraction. Journal of Computational Physics, 27(1):42–70, April 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900256.

Gourlay:1970:CMF

[GM70] A. R. Gourlay and J. Ll Morris. On the comparison of multistep formulations of the optimized Lax–Wendroff method for nonlinear hyperbolic systems in two space variables. Journal of Computational Physics, 5(2):229–243, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900616.

Greig:1976:HMK

[GM76] I. S. Greig and J. Ll Morris. A hopscotch method for the Korteweg– de-Vries equation. Journal of Computational Physics, 20(1):64– 80, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176901029.

Gupta:1979:BAA

[GM79a] Murli M. Gupta and Ram P. Manohar. Boundary approximations and accuracy in viscous flow computations. Journal of Computational Physics, 31(2):265–288, May 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917990072X.

Gupta:1979:DSB

[GM79b] Murli M. Gupta and Ram P. Manohar. Direct solution of the biharmonic equation using noncoupled approach. Journal of Computational Physics, 33(2):236–248, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900184.

Grosch:1977:NSP

[GO77] Chester E. Grosch and Steven A. Orszag. Numerical solution of problems in unbounded regions: Coordinate transforms. Journal of Computational Physics, 25(3):273–295, November 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177901024.

Godfrey:1974:NCI

[God74] Brendan B. Godfrey. Numerical Cherenkov instabilities in electromagnetic particle codes. Journal of Computational Physics, 15(4): 504–521, August 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917490076X.

Godfrey:1975:CMN

[God75] Brendan B. Godfrey. Canonical momenta and numerical instabilities in particle codes. Journal of Computational Physics, 19(1):58– 76, September 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175901163.

Goda:1979:MTI

[God79] Katuhiko Goda. A multistep technique with implicit difference schemes for calculating two- or three-dimensional cavity flows. Journal of Computational Physics, 30(1):76–95, January 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900883.

Goldman:1972:NML

[Gol72] Edward B. Goldman. Numerical modeling of laser produced plasmas: a second-order integration scheme. Journal of Computational Physics, 9(2):380–384, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900319.

Gottlieb:1974:SRT

[Got74] David Gottlieb. On the stability of Rusanov’s third-order scheme. Journal of Computational Physics, 15(3):421–426, July 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174901211.

Gawain:1970:UHM

[GP70] T. H. Gawain and J. W. Pritchett. A unified heuristic model of fluid turbulence. Journal of Computational Physics, 5(3):383– 405, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900707.

Graves:1973:PI

[Gra73] Randolph A. Graves, Jr. Partial implicitization. Journal of Computational Physics, 13(3):439–444, November 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917390048X.

Graves:1976:SDF

[Gra76] Randolph A. Graves, Jr. Solutions to divergence form equations using the method of partial implicitization. Journal of Computational Physics, 21(3):340–342, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900310.

Gruber:1978:FHE

[Gru78] R. Gruber. Finite hybrid elements to compute the ideal magnetohydrodynamic spectrum of an axisymmetric plasma. Journal of Computational Physics, 26(3):379–389, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900761.

Grundy:1979:BTS

[Gru79] R. E. Grundy. Borel-type sums using two-point rational approximants. Journal of Computational Physics, 33(2):290–293, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900226.

Grimm:1970:NSS

[GS70] R. Grimm and R. G. Storer. The numerical solution of Schrödinger’s equation with a nonlocal potential. Journal of Computational Physics, 5(2):350–354, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900677.

Grimm:1971:MCS

[GS71] R. C. Grimm and R. G. Storer. Monte–Carlo solution of Schrödinger’s equation. Journal of Computational Physics, 7(1):134–156, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900544.

Gagne:1977:SSN

[GS77] Réal R. J. Gagné and Magdi M. Shoucri. A splitting scheme for the numerical solution of a one-dimensional Vlasov equation. Journal of Computational Physics, 24(4):445–449, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900328.

Gelbard:1978:NSD

[GS78] Fred Gelbard and John H. Seinfeld. Numerical solution of the dynamic equation for particulate systems. Journal of Computational Physics, 28(3):357–375, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917890058X.

Grimm:1972:INM

[GSD72] R. C. Grimm, R. G. Storer, and B. Davies. Improvements on a new method for the numerical solution of the Schrödinger equation. Journal of Computational Physics, 9(3):538–546, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900095.

Gottlieb:1974:PES

[GT74] David Gottlieb and Eli Turkel. Phase error and stability of second order methods for hyperbolic problems. II. Journal of Computational Physics, 15(2):251–265, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900874.

Gottlieb:1978:BCM

[GT78a] David Gottlieb and Eli Turkel. Boundary conditions for multistep finite-difference methods for time-dependent equations. Journal of Computational Physics, 26(2):181–196, February 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178900906.

Gottlieb:1978:AMS

[GT78b] David Gottlieb and Eli Turkel. On acceleration of MacCormack’s scheme. Journal of Computational Physics, 26(2):252– 256, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900967.

Gustafsson:1971:ADI

[Gus71] Bertil Gustafsson. An alternating direction implicit method for solving the shallow water equations. Journal of Computational Physics, 7(2):239–254, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900878.

Gustafsson:1980:CBC

[Gus80] Bertil Gustafsson. Comment on “On boundary conditions for hyperbolic difference schemes”. Journal of Computational Physics, 34(1):108–110, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180901151.

See [Gar78].

Gammel:1972:SOE

[GW72] J. L. Gammel and W. R. Wortman. Sixth-order effects in nucleonnucleon scattering. Journal of Computational Physics, 10(3):568– 606, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900538.

Guderley:1973:EAC

[GW73] K. G. Guderley and R. H. Warren. An economical approximation for the coefficients in the development of a function with respect to an orthogonal system. Journal of Computational Physics, 11(1):146–151, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901526.

Gangi:1978:WRW

[GW78] Anthony F. Gangi and Robert L. Wesson. P -wave to Rayleighwave conversion coefficients for wedge corners; model experiments. Journal of Computational Physics, 29(3):370–388, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901407.

Harlow:1971:NFD

[HA71] Francis H. Harlow and Anthony A. Amsden. A numerical fluid dynamics calculation method for all flow speeds. Journal of Computational Physics, 8(2):197–213, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900027.

Harlow:1974:MFC

[HA74] Francis H. Harlow and Anthony A. Amsden. Multifluid flow calculations at all Mach numbers. Journal of Computational Physics, 16(1):1–19, September 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174901004.

Harlow:1975:FIM

[HA75a] Francis H. Harlow and Anthony A. Amsden. Flow of interpenetrating material phases. Journal of Computational Physics, 18(4): 440–464, August 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900960.

Harlow:1975:NCM

[HA75b] Francis H. Harlow and Anthony A. Amsden. Numerical calculation of multiphase fluid flow. Journal of Computational Physics, 17(1): 19–52, January 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900613.

Hirt:1974:ALE

[HAC74] C. W. Hirt, A. A. Amsden, and J. L. Cook. An arbitrary Lagrangian–Eulerian computing method for all flow speeds. Journal of Computational Physics, 14(3):227–253, March 1974. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999174900515.

Haidvogel:1979:RDB

[Hai79] Dale B. Haidvogel. Resolution of downstream boundary layers in the Chebyshev approximation to viscous flow problems. Journal of Computational Physics, 33(3):313–324, December 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917990158X.

Hanson:1972:EES

[Han72] James N. Hanson. Experiments with equation solutions by functional analysis algorithms and formula manipulation. Journal of Computational Physics, 9(1):26–52, February 1972. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999172900356.

Harlow:1976:RFD

[HAN76] F. H. Harlow, A. A. Amsden, and J. R. Nix. Relativistic fluid dynamics calculations with the particle-in-cell technique. Journal of Computational Physics, 20(2):119–129, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900589.

Harlow:1971:CDN

[Har71] Francis H. Harlow. Contour dynamics for numerical fluid-flow calculations. Journal of Computational Physics, 8(2):214–229, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900039.

Hargrave:1976:NAE

[Har76] B. A. Hargrave. Numerical approximation of eigenvalues of Sturm– Liouville systems. Journal of Computational Physics, 20(4):381– 396, April 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900905.

Haridoss:1978:STT

[Har78] S. Haridoss. Simulation of thermoluminescence and thermally stimulated currents. Journal of Computational Physics, 26(2):232– 235, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900931.

Hategan:1970:ISS

[Hat70] C. Hategan. Irregular solution of a system of coupled Schrödinger equations. Journal of Computational Physics, 6(2):333–335, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917090032X.

Haussling:1979:BFC

[Hau79] H. J. Haussling. Boundary-fitted coordinates for accurate numerical solution of multibody flow problems. Journal of Computational Physics, 30(1):107–124, January 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900901.

Hung:1977:IFD

[HB77] Tin-Kan Hung and Thomas D. Brown. An implicit finite-difference method for solving the Navier–Stokes equation using orthogonal curvilinear coordinates. Journal of Computational Physics, 23(4): 343–363, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900675.

Hitchcock:1978:NIP

[HBH78] D. A. Hitchcock, S. H. Brecht, and W. Horton, Jr. Note: an improved polynomial representation for the delta function, δ(μ − μ_{0}). Journal of Computational Physics, 26(3):443–446, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900797.

Hirt:1972:CTD

[HC72] C. W. Hirt and J. L. Cook. Calculating three-dimensional flows around structures and over rough terrain. Journal of Computational Physics, 10(2):324–340, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900708.

Hirt:1970:LMC

[HCB70] C. W. Hirt, J. L. Cook, and T. D. Butler. A Lagrangian method for calculating the dynamics of an incompressible fluid with free surface. Journal of Computational Physics, 5(1):103– 124, February 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900550.

Hung:1972:NSR

[HCH72] W. S. Y. Hung, C. S. Chen, and J. K. Haviland. Numerical solutions of reactive fluid flows during postignition transients in hybrid rocket systems. Journal of Computational Physics, 9(2): 167–193, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900174.

Houston:1974:ASF

[HD74] M. H. Houston, Jr. and J. Cl. De Bremaecker. ADI solution of free convection in a variable viscosity fluid. Journal of Computational Physics, 16(3):221–239, November 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900928.

Hendry:1979:GEM

[HD79] J. A. Hendry and L. M. Delves. The global element method applied to a harmonic mixed boundary value problem. Journal of Computational Physics, 33(1):33–44, October 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900263.

Harten:1978:PTS

[HE78] Amiram Harten and Shamir Efrony. A partition technique for the solution of potential flow problems by integral equation methods. Journal of Computational Physics, 27(1):71–87, April 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900268.

Hedstrom:1979:NBC

[Hed79] G. W. Hedstrom. Nonreflecting boundary conditions for nonlinear hyperbolic systems. Journal of Computational Physics, 30(2):222– 237, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901001.

Hewett:1974:NSR

[Hew74] D. W. Hewett. Numerical simulation of a radially inhomogeneous cylindrical plasma. Journal of Computational Physics, 16(1):49– 65, September 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917490103X.

Householder:1971:DES

[HF71] A. S. Householder and Kenneth Fox. Determination of eigenvectors of symmetric idempotent matrices. Journal of Computational Physics, 8(2):292–294, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917190009X.

Hager:1979:RGC

[HG79] Douglas F. Hager and Roy G. Gordon. The reduction of a general complex matrix to a condensed form by bounded single element transformations. Journal of Computational Physics, 32(3):377– 402, September 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901517.

Hockney:1974:QHR

[HGE74] R. W. Hockney, S. P. Goel, and J. W. Eastwood. Quiet highresolution computer models of a plasma. Journal of Computational Physics, 14(2):148–158, February 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900102.

Hawk:1976:FDS

[HH76] I. L. Hawk and D. L. Hardcastle. Finite-difference solution to the Schrödinger equation for the ground state and first-excited state of helium. Journal of Computational Physics, 21(2):197– 207, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900115.

Harrington:1977:ANF

[HH77] Steven Jay Harrington and Alistair D. C. Holden. Automation of the nonrelativistic ϕ^{3} field theory in two spatial dimensions. Journal of Computational Physics, 25(3):296–313, November 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177901036.

Hendry:1978:GEM

[HH78] J. A. Hendry and M. A. Hennell. The global element method applied to a discontinuous nuclear potential. Journal of Computational Physics, 28(3):376–394, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900591.

Hillion:1978:NSH

[Hil78] Pierre Hillion. Numerical solution of the Helmholtz equation in optics. Journal of Computational Physics, 28(2):232–252, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900360.

Hillion:1979:NSS

[Hil79] Pierre Hillion. Numerical solution of the spinor Helmholtz equation in optics. Journal of Computational Physics, 32(1):56– 70, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901414.

Hirsh:1975:HOA

[Hir75] Richard S. Hirsh. Higher order accurate difference solutions of fluid mechanics problems by a compact differencing technique. Journal of Computational Physics, 19(1):90–109, September 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175901187.

Horak:1978:ADR

[HJKS78] H. G. Horak, E. M. Jones, J. W. Kodis, and M. T. Sandford II. An algorithm for the discrete rezoning of Lagrangian meshes. Journal of Computational Physics, 26(3):277–284, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900700.

Hembd:1972:STI

[HK72] H. Hembd and H. Kschwendt. Solution of the transport integral equation with anisotropic scattering. Journal of Computational Physics, 10(3):534–553, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900514.

Hunter:1976:AEM

[HK76] G. Hunter and Mary Kuriyan. Asymptotic expansions of Mathieu functions in wave mechanics. Journal of Computational Physics, 21(3):319–325, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900280.

Hajj:1974:NSS

[HKN74] F. Y. Hajj, H. Kobeisse, and N. R. Nassif. On the numerical solution of Schrödinger’s radial equation. Journal of Computational Physics, 16(2):150–159, October 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174901090.

Hall:1975:NSA

[HL75] C. A. Hall and J. J. Lowke. A non-self-adjoint finite difference model of electron diffusion parallel to electric fields. Journal of Computational Physics, 19(3):297–310, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900790.

Howell:1976:CMQ

[HL76] T. Howell and J. N. Lyness. On a conjecture in multidimensional quadrature. Journal of Computational Physics, 21(3):361– 364, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900358.

Hughes:1979:FEA

[HLB79] Thomas J. R. Hughes, Wing Kam Liu, and Alec Brooks. Finite element analysis of incompressible viscous flows by the penalty function formulation. Journal of Computational Physics, 30(1): 1–60, January 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917990086X.

Haile:1976:MCS

[HLM^{+}76] J. M. Haile, D. Litchinsky, R. McPherson, C. G. Gray, and K. E. Gubbins. Monte Carlo simulation of molecular fluids on a minicomputer. Journal of Computational Physics, 21(2):227– 237, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900139.

Houstis:1978:ENM

[HLRP78] E. N. Houstis, R. E. Lynch, J. R. Rice, and T. S. Papatheodorou. Evaluation of numerical methods for elliptic partial differential equations. Journal of Computational Physics, 27(3):323–350, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900141.

Hall:1975:TIS

[HM75] Charles A. Hall and Boris A. Mutafelija. Transfinite interpolation of steam tables. Journal of Computational Physics, 18(1): 79–91, May 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175901035.

Hardy:1973:NME

[HMB73] Robert J. Hardy, Ira W. Morrison, and Sudhir Bijanki. Numerical methods for evaluating integrals over first Brillouin zones of cubic lattices. Journal of Computational Physics, 13(4):591–610, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900351.

Helton:1977:TDM

[HMR77] F. J. Helton, R. L. Miller, and John M. Rawls. Two-dimensional multi-fluid tokamak transport code. Journal of Computational Physics, 24(2):117–132, June 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900523.

Holt:1970:NMC

[HN70] Maurice Holt and D. Ejike Ndefo. A numerical method for calculating steady unsymmetrical supersonic flow past cones. Journal of Computational Physics, 5(3):463–486, June 1970. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999170900756.

Hillion:1977:IHO

[HN77] P. Hillion and G. Nurdin. Integration of highly oscillatory functions. Journal of Computational Physics, 23(1):74–81, January 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900894.

Hewett:1978:MQP

[HN78] D. W. Hewett and C. W. Nielson. A multidimensional quasineutral plasma simulation model. Journal of Computational Physics, 29(2):219–236, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901535.

Hockney:1971:MCH

[Hoc71] R. W. Hockney. Measurements of collision and heating times in a two-dimensional thermal computer plasma. Journal of Computational Physics, 8(1):19–44, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900325.

Hoffman:1973:ASN

[Hof73] Joe D. Hoffman. Accuracy studies of the numerical method of characteristics for axisymmetric, steady supersonic flows. Journal of Computational Physics, 11(2):210–239, February 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900041.

Hoffman:1974:EPS

[Hof74] Gilbert H. Hoffman. Extension of perturbation series by computer: Viscous flow between two infinite rotating disks. Journal of Computational Physics, 16(3):240–258, November 1974. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917490093X.

Hohl:1972:ESD

[Hoh72] Frank Hohl. Evolution of a stationary disk of stars. Journal of Computational Physics, 9(1):10–25, February 1972. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900344.

Holas:1977:NIS

[Hol77] A. Holas. Numerical integration over the solid angle and volume of the Brillouin zone. Journal of Computational Physics, 23(2):150– 166, February 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177901188.

Ho:1975:ESI

[HP75] Y. K. Ho and B. A. P. Page. Evaluation of some integrals required in low-energy electron or positron-atom scattering. Journal of Computational Physics, 17(2):122–131, February 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900327.

Hanson:1976:BCS

[HP76] M. E. Hanson and A. G. Petschek. A boundary condition for significantly reducing boundary reflections with a Lagrangian mesh. Journal of Computational Physics, 21(3):333–339, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900309.

Hughes:1978:MCA

[HP78] M. H. Hughes and D. E. Post. A Monte Carlo algorithm for calculating neutral gas transport in cylindrical plasmas. Journal of Computational Physics, 28(1):43–55, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900451.

Howarth:1974:GRI

[HPS74] B. A. Howarth, T. J. F. Pavlasek, and P. Silvester. A graphical representation for interpreting scalar wave multiple-scattering phenomena. Journal of Computational Physics, 15(2):266–285, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900886.

Huntley:1978:NSL

[HPZ78] E. Huntley, W. M. Pickering, and A. S. I. Zinober. The numerical solution of linear time-dependent partial differential equations by the Laplace and fast Fourier transforms. Journal of Computational Physics, 27(2):256–271, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900086.

Hirsh:1974:RDD

[HR74] Richard S. Hirsh and David H. Rudy. The role of diagonal dominance and cell Reynolds number in implicit difference methods for fluid mechanics problems. Journal of Computational Physics, 16(3):304–310, November 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900989.

Holmlid:1978:UCP

[HR78] Leif Holmlid and Kjell Rynefors. Uniformity of congruential pseudorandom number generators. Dependence on length of number sequence and resolution. Journal of Computational Physics, 26(3): 297–306, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900724.

Heller:1973:EQC

[HRY73] Eric J. Heller, William P. Reinhardt, and Hashim A. Yamani. On an “equivalent quadrature” calculation of matrix elements of (z − p^{2}/2m)^{−1} using an L^{2} expansion technique. Journal of Computational Physics, 13(4):536–550, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900326.

Heydweiller:1976:SDS

[HS76] J. C. Heydweiller and R. F. Sincovec. A stable difference scheme for the solution of hyperbolic equations using the method of lines. Journal of Computational Physics, 22(3):377–388, November 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900553.

Hill:1974:NEA

[HSM74] T. R. Hill, J. Kenneth Shultis, and J. O. Mingle. Numerical evaluation of the azimuthally dependent albedo problem in slab geometry. Journal of Computational Physics, 15(2):200– 211, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900849.

Hall:1974:CSI

[HU74] David L. Hall and Peter D. Usher. Coordinate stretching and interface location III. A general relaxation method with application to composite polytropes. Journal of Computational Physics, 15(3): 345–361, July 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901168.

Hubbard:1971:SMS

[Hub71] J. B. Hubbard. Statistical mechanics of small systems. Journal of Computational Physics, 7(3):502–514, June 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999171901070.

Huddleston:1972:SSV

[Hud72] Robert E. Huddleston. Selection of stepsize in the variablestep predictor-corrector method of van Wyk. Journal of Computational Physics, 9(3):528–537, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900083.

Hunter:1978:SEL

[Hun78] C. Hunter. Some extensions of the Lax–Wendroff method. Journal of Computational Physics, 27(3):447–452, June 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178900219.

Haney:1975:RBG

[HW75] Robert L. Haney and Julian M. Wright, Jr. The relationship between the grid size and the coefficient of nonlinear lateral eddy viscosity in numerical ocean circulation models. Journal of Computational Physics, 19(3):257–266, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900765.

Hyman:1977:MRL

[Hym77] James M. Hyman. Mesh refinement and local inversion of elliptic partial differential equations. Journal of Computational Physics, 23(2):124–134, February 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177901164.

Harten:1972:SAH

[HZ72] A. Harten and G. Zwas. Self-adjusting hybrid schemes for shock computations. Journal of Computational Physics, 9(3):568–583, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900125.

Haidvogel:1979:ASP

[HZ79] Dale B. Haidvogel and Thomas Zang. The accurate solution of Poisson’s equation by expansion in Chebyshev polynomials. Journal of Computational Physics, 30(2):167–180, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900974.

Icschwendt:1970:LEI

[Ics70] H. Icschwendt. Legendre expansion and integral equations of displacement type. Journal of Computational Physics, 5(1):84– 102, February 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900549.

Cheng:1978:CAM

[iCS78] Sin i Cheng and Gregory Shubin. Computational accuracy and mesh Reynolds number. Journal of Computational Physics, 28(3):315–326, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900566.

Ilan:1978:SFD

[Ila78] Almoga Ilan. Stability of finite difference schemes for the problem of elastic wave propagation in a quarter plane. Journal of Computational Physics, 29(3):389–403, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901419.

Isenberg:1973:CCN

[Ise73] C. Isenberg. A conjecture on the convergence of numerical estimates for multidimensional integrals with singular integrands. Journal of Computational Physics, 12(3):417–421, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901630.

Ixaru:1972:EAA

[Ixa72] L. Gr Ixaru. The error analysis of the algebraic method for solving the Schrödinger equation. Journal of Computational Physics, 9(1):159–163, February 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900411.

Jacobs:1973:SIP

[Jac73] D. A. H. Jacobs. The strongly implicit procedure for biharmonic problems. Journal of Computational Physics, 13(3):303– 315, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900387.

James:1977:SPE

[Jam77] R. A. James. The solution of Poisson’s equation for isolated source distributions. Journal of Computational Physics, 25(2):71– 93, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900134.

Jonas:1972:NSP

[JB72] P. R. Jonas and J. T. Bartlett. The numerical simulation of particle motion in a homogeneous field of turbulence. Journal of Computational Physics, 9(2):290–302, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900241.

Jamet:1975:NSE

[JB75] P. Jamet and R. Bonnerot. Numerical solution of the Eulerian equations of compressible flow by a finite element method which follows the free boundary and the interfaces. Journal of Computational Physics, 18(1):21–45, May 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917590100X.

Jaffe:1974:UCM

[JCM74] H. H. Jaffé, H. M. Chang, and C. A. Masmanidis. The use of the CNDO method in spectroscopy. X. The calculation of the self-consistent doublet and triplet states. Journal of Computational Physics, 14(2):180–192, February 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900126.

Junior:1970:LMF

[JD70] P. H. Junior and H. Deitinghoff. Low magnetic-field optimisation of focusing device in a helix linear accelerator. Journal of Computational Physics, 5(1):1–12, February 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900483.

Janssens:1974:GTC

[JD74] M. Janssens and E. De Vos. Generation technique for chains on lattices in computer experiments. Journal of Computational Physics, 15(4):529–532, August 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900783.

Johnson:1979:NDA

[JDG^{+}79] J. L. Johnson, H. E. Dalhed, J. M. Greene, R. C. Grimm, Y. Y. Hsieh, S. C. Jardin, J. Manickam, M. Okabayashi, R. G. Storer, A. M. M. Todd, D. E. Voss, and K. E. Weimer. Numerical determination of axisymmetric toroidal magnetohydrodynamic equilibria. Journal of Computational Physics, 32(2):212– 234, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901293.

Jaffe:1974:TSM

[JE74] H. H. Jaffé and R. L. Ellis. A treatment of symmetry in MO calculations. Journal of Computational Physics, 16(1):20–31, September 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901016.

Jenkins:1978:CIG

[Jen78] T. R. Jenkins. On computing the integral of glow curve theory. Journal of Computational Physics, 29(2):302–305, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901626.

Jespersen:1974:AMF

[Jes74] Dennis C. Jespersen. Arakawa’s method is a finite-element method. Journal of Computational Physics, 16(4):383–390, December 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900473.

Janin:1974:NEE

[JH74] G. Janin and M. J. Haggerty. Numerical experiments on expanding gravitational systems. Journal of Computational Physics, 16(1):76– 92, September 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901053.

Jardin:1978:DGM

[JJGG78] S. C. Jardin, J. L. Johnson, J. M. Greene, and R. C. Grimm. Dynamical grid method for time-dependent simulations of axisymmetric instabilities in tokamaks. Journal of Computational Physics, 29(1):101–126, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901122.

Johnson:1977:GCS

[JKAR77] Thomas H. Johnson, John Killeen, Oscar A. Anderson, and Marvin E. Rensink. Guiding-center simulation of toroidal plasmas. Journal of Computational Physics, 23(3):219–241, March 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900924.

Joyce:1971:NIM

[JKM71] Glenn Joyce, Georg Knorr, and Homer K. Meier. Numerical integration methods of the Vlasov equation. Journal of Computational Physics, 8(1):53–63, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900349.

Jones:1977:TKS

[JMW77] W. E. Jones, G. Matinopoulos, and J. S. Wasson. The truncated Kullback statistic as an objective function for fitting partially known probability vectors. Journal of Computational Physics, 23(4):446–451, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900730.

James:1971:CAF

[JN71] P. B. James and G. R. North. Computer analysis of Feynman diagrams for a static model. Journal of Computational Physics, 7(2): 354–357, April 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900957.

Jobert:1978:CSP

[Job78] N. Jobert. Contribution of some particularities in the dispersion curves to numerical seismograms computed by normal modes. Journal of Computational Physics, 29(3):404–411, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901420.

Johnson:1973:MLD

[Joh73] B. R. Johnson. The multichannel log-derivative method for scattering calculations. Journal of Computational Physics, 13(3):445– 449, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900491.

Johnson:1976:FTP

[Joh76] Thomas H. Johnson. Failure testing: A proposal for increasing confidence in the results of numerical simulations. Journal of Computational Physics, 21(2):245–250, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900164.

Jones:1973:STO

[Jon73] D. J. Jones. Solution of Troesch’s, and other, two point boundary value problems by shooting techniques. Journal of Computational Physics, 12(3):429–434, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901654.

Jones:1979:SAM

[Jon79] Everett Jones. Stability analysis for MacCormack’s time-splitting technique applied to a model conduction problem. Journal of Computational Physics, 30(3):389–406, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901220.

Joshi:1971:GAT

[Jos71] Bhairav D. Joshi. Generalized from of atomic two electron integrals over hydrogenic functions. Journal of Computational Physics, 8(2):300–303, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900118.

Jain:1978:FRF

[JP78] Anil K. Jain and Robert A. Padgug. Fast restoration of finite objects degraded by finite PSF. Journal of Computational Physics, 28(2):167–180, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900323.

Jurak:1972:LSS

[JPC72] K. Jurak, G. J. Pert, and C. E. Capjack. Least-squares solution for the blunt body hypersonic flow problem. Journal of Computational Physics, 10(2):369–373, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900733.

Jones:1972:NSE

[JSK72] D. J. Jones, J. C. South, Jr., and E. B. Klunker. On the numerical solution of elliptic partial differential equations by the method of lines. Journal of Computational Physics, 9(3):496– 527, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900071.

Kaldor:1976:AGG

[Kal76] Uzi Kaldor. An algorithm for generating Goldstone and Bloch– Brandow diagrams. Journal of Computational Physics, 20(4):432– 441, April 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900929.

Kane:1976:NIS

[Kan76] Ian David Kane. A numerical investigation into the solution of the equations describing baroclinic–barotropic instability in the atmosphere. Journal of Computational Physics, 20(1):1– 21, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917690098X.

Kaup:1972:MVP

[Kau72] David J. Kaup. A modification of the variable phase method for optimum numerical calculation of phase shifts. Journal of Computational Physics, 9(2):254–272, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900228.

Kaup:1973:ONC

[Kau73] David J. Kaup. Optimum numerical calculation of JWKB phase shifts. Journal of Computational Physics, 12(2):285–287, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002199917380018X.

Kazhdan:1970:ABC

[Kaz70] Ja. M. Kazhdan. Asymptotic behavior of a convergent spherical rarefaction wave. Journal of Computational Physics, 5(3):540– 546, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900793.

Kestemont:1971:MDR

[KB71] E. Kestemont and A. Bellemans. Molecular dynamics of rough disks assemblies and dielectric relaxation. Journal of Computational Physics, 7(3):515–527, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171901082.

Kolman:1973:SMR

[KB73] Bernard Kolman and Robert E. Beck. A stopping method for Racah’s formulas. Journal of Computational Physics, 13(1):161– 163, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901320.

Krishnaprasad:1977:DAC

[KB77] P. S. Krishnaprasad and Richard Barakat. A descent approach to a class of inverse problems. Journal of Computational Physics, 24(4): 339–347, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900262.

Kerr:1979:ALS

[KB79a] Christopher L. Kerr and Alan F. Blumberg. An analysis of a local second-moment conserving quasi-Lagrangian scheme for solving the advection equation. Journal of Computational Physics, 32 (1):1–9, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901372.

Klein:1979:TSN

[KB79b] H. H. Klein and R. N. Byrne. A technique for solving the non-linear partial differential equations of Tokamak transport. Journal of Computational Physics, 33(2):294–299, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900238.

Keller:1972:IPB

[KC72] Herbert B. Keller and Tuncer Cebeci. An inverse problem in boundary-layer flows: Numerical determination of pressure gradient for a given wall shear. Journal of Computational Physics, 10(1): 151–161, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900964.

King:1976:NIU

[KD76] Harry F. King and Michel Dupuis. Numerical integration using Rys polynomials. Journal of Computational Physics, 21(2):144– 165, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900085.

Kruer:1973:DEM

[KDR73] W. L. Kruer, J. M. Dawson, and B. Rosen. The dipole expansion method for plasma simulation. Journal of Computational Physics, 13(1):114–129, September 1973. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901290.

Koda:1979:SAP

[KDS79] Masato Koda, Ali H. Dogru, and John H. Seinfeld. Sensitivity analysis of partial differential equations with application to reaction and diffusion processes. Journal of Computational Physics, 30(2):259– 282, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901037.

Kentzer:1970:TFP

[Ken70] Czeslaw P. Kentzer. Transonic flows past a circular cylinder. Journal of Computational Physics, 6(2):168–182, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900197.

Kershaw:1978:ICC

[Ker78] David S. Kershaw. The incomplete Cholesky-conjugate gradient method for the iterative solution of systems of linear equations. Journal of Computational Physics, 26(1):43–65, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900980.

Kubicek:1975:DEB

[KH75a] Milan Kubícek and Vladimír Hlavácek. Direct evaluation of branching points for equations arising in the theory of explosions of solid explosives. Journal of Computational Physics, 17(1):79– 86, January 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900649.

Kubicek:1975:STT

[KH75b] Milan Kubicek and Vladimír Hlavácek. Solution of Troesch’s two-point boundary value problem by shooting technique. Journal of Computational Physics, 17(1):95–101, January 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900662.

Klenk:1975:EPP

[KK75] K. F. Klenk and A. S. Kanofsky. Evaluation of proton-proton elastic scattering with quark hard cores using Glauber and Monte Carlo methods. Journal of Computational Physics, 19(4):353– 375, December 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900686.

Koehler:1976:EAM

[KL76] T. R. Koehler and P. A. Lee. Efficient analysis of molecular dynamics data. Journal of Computational Physics, 22(3):319–329, November 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900528.

Kriegsmann:1978:NSE

[KM78] Gregory Kriegsmann and Catiileen S. Morawetz. Numerical solutions of exterior problems with the reduced wave equation. Journal of Computational Physics, 28(2):181–197, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900335.

Knirk:1976:AEQ

[Kni76] Dwayne L. Knirk. Accurate and efficient quadrature for Volterra integral equations. Journal of Computational Physics, 21(4):371– 399, August 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900371.

Knorr:1973:PSF

[Kno73] Georg Knorr. Plasma simulation with few particles. Journal of Computational Physics, 13(2):165–180, October 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900211.

Knop:1975:MCG

[Kno75] R. E. Knop. Monte Carlo generation of two body resonant states. Journal of Computational Physics, 18(2):115–118, June 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900236.

Korn:1978:NDT

[Kor78] David G. Korn. Numerical design of transonic cascades. Journal of Computational Physics, 29(1):20–34, October 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178901067.

Kovahic:1970:ICM

[Kov70a] P. Kovahic. Inversion of a covariance matrix. Journal of Computational Physics, 5(2):355–357, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900689.

Kovanic:1970:ALU

[Kov70b] Pavel Kovanic. Accuracy limitations in an unbiased optimum data treatment. Journal of Computational Physics, 6(3):473–482, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900434.

Klink:1976:ANB

[KP76] W. H. Klink and G. L. Payne. Approximating with nonorthogonal basis functions. Journal of Computational Physics, 21(2): 208–226, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900127.

Kalaba:1971:RDP

[KR71] Robert Kalaba and Enrique H. Ruspine. Reduction of the Dirichlet problem to an initial value problem. Journal of Computational Physics, 8(3):489–497, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900271.

Kalaba:1973:IVM

[KS73] Robert E. Kalaba and Melvin R. Scott. An initial value method for integral operators: I-complex eigenvalues. Journal of Computational Physics, 12(3):364–368, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901587.

Kenan:1974:NEF

[KS74a] Richard P. Kenan and Paul R. Sievert. Numerical evaluation of the free-electron lattice Green’s function. Journal of Computational Physics, 14(4):350–360, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900187.

Knorr:1974:PSE

[KS74b] Georg Knorr and Magdi Shoucri. Plasma simulation as eigenvalue problem. Journal of Computational Physics, 14(1):1–7, January 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900011.

Kaper:1970:NES

[KSV70] Hans G. Kaper, J. Kenneth Shultis, and Jaap G. Veninga. Numerical evaluation of the slab albedo problem solution in onespeed anisotropic transport theory. Journal of Computational Physics, 6(2):288–313, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900264.

Kusic:1973:SCS

[Kus73] G. L. Kusic. Self-consistent space-charge-flow using integral equation field methods. Journal of Computational Physics, 13(4):551– 568, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900338.

Kestemont:1976:CCF

[KV76] E. Kestemont and J. Van Craen. On the computation of correlation functions in molecular dynamics experiments. Journal of Computational Physics, 22(4):451–458, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900437.

Kirkpatrick:1972:NMD

[KW72] John R. Kirkpatrick and William F. Walker. Numerical method for describing turbulent, compressible, subsonic separated jet flows. Journal of Computational Physics, 10(2):185–201, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900599.

Lado:1971:NFT

[Lad71] F. Lado. Numerical Fourier transforms in one, two, and three dimensions for liquid state calculations. Journal of Computational Physics, 8(3):417–433, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900210.

Langdon:1970:ESG

[Lan70] A. Bruce Langdon. Effects of the spatial grid in simulation plasmas. Journal of Computational Physics, 6(2):247–267, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900240.

Langdon:1973:ECP

[Lan73] A. Bruce Langdon. ”energy-conserving” plasma simulation algorithms. Journal of Computational Physics, 12(2):247–268, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999173800142.

Langdon:1979:ATI

[Lan79] A. Bruce Langdon. Analysis of the time integration in plasma simulation. Journal of Computational Physics, 30(2):202–221, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900998.

Lapidus:1971:CRS

[Lap71] A. Lapidus. Computation of radially symmetric shocked flows. Journal of Computational Physics, 8(1):106–118, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999171900386.

Larson:1978:FPS

[Lar78] Richard B. Larson. A finite-particle scheme for three-dimensional gas dynamics. Journal of Computational Physics, 27(3):397– 409, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900177.

Laura:1971:CMS

[Lau71] Patricio A. Laura. A collocation method for solving Laplace’s equation. Journal of Computational Physics, 8(2):311–312, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900143.

Lau:1979:CFD

[Lau79] P. C. M. Lau. Curvilinear finite difference method for threedimensional potential problems. Journal of Computational Physics, 32(3):325–344, September 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901499.

Lavine:1973:NAL

[Lav73] J. P. Lavine. A numerical aide in the location of few-body bound states. Journal of Computational Physics, 12(4):561– 566, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901071.

Lucchese:1978:BOE

[LBM^{+}78] Robert R. Lucchese, Bernard R. Brooks, James H. Meadows, William C. Swope, and Henry F. Schaefer III. BERKELEY: An “open ended” Configuration Interaction (CI) program designed for minicomputers. Journal of Computational Physics, 26(2):243– 251, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900955.

Lilley:1972:ISP

[LC72] D. G. Lilley and N. A. Chigier. An inverse solution procedure for turbulent swirling boundary layer combustion flow. Journal of Computational Physics, 9(2):237–253, April 1972. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999172900216.

LeBail:1972:UFF

[Le 72] Roland C. Le Bail. Use of fast Fourier transforms for solving partial differential equations in physics. Journal of Computational Physics, 9(3):440–465, June 1972. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900058.

Lepage:1978:NAA

[Lep78] G. Peter Lepage. A new algorithm for adaptive multidimensional integration. Journal of Computational Physics, 27(2):192– 203, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900049.

Lesk:1973:FPS

[Les73] A. M. Lesk. A FORTRAN program for the solution of simultaneous linear Boolean inequalities by the algorithm of Hammer and Rudeanu. Journal of Computational Physics, 12(1):150– 152, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901770.

Lewis:1970:ECN

[Lew70] H. Ralph Lewis. Energy-conserving numerical approximations for Vlasov plasmas. Journal of Computational Physics, 6(1):136– 141, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900124.

Lewis:1972:VAN

[Lew72] H. Ralph Lewis. Variational algorithms for numerical simulation of collisionless plasma with point particles including electromagnetic interactions. Journal of Computational Physics, 10(3):400– 419, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900447.

Levin:1978:GME

[LFMM78] Deborah A. Levin, Arne W. Fluflet, Michael Ma, and Vincent McKoy. Gaussian matrix elements of the free-particle Green’s function. Journal of Computational Physics, 28(3):416–425, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900621.

LaBudde:1974:DMG

[LG74] Robert A. LaBudde and Donald Greenspan. Discrete mechanics — a general treatment. Journal of Computational Physics, 15(2): 134–167, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900813.

Loly:1972:EBZ

[LH72] P. D. Loly and B. J. S. Huett. Evaluation of Brillouin-zone sums on the cubic lattices. Journal of Computational Physics, 9(3): 592–605, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900149.

Liebermann:1978:THE

[Lie78] Raanan Liebermann. Traces of high-energy processes in strong magnetic fields. Journal of Computational Physics, 29(2):208– 218, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901523.

Lindman:1970:DRC

[Lin70] E. L. Lindman. Dispersion relation for computer-simulated plasmas. Journal of Computational Physics, 5(1):13–22, February 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900495.

Lindemuth:1975:CFC

[Lin75a] Irvin R. Lindemuth. Conservation form in computational magnetohydrodynamics. Journal of Computational Physics, 18(2):119– 131, June 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900248.

Lindman:1975:FSB

[Lin75b] E. L. Lindman. “Free-space” boundary conditions for the time dependent wave equation. Journal of Computational Physics, 18(1): 66–78, May 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175901023.

Lindemuth:1977:BCC

[Lin77] Irvin R. Lindemuth. A boundary condition for computational magnetohydrodynamics. Journal of Computational Physics, 25(2):104– 117, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900158.

Lipps:1977:IOD

[Lip77] Frank B. Lipps. Inertial oscillations and the DuFort–Frankel difference scheme. Journal of Computational Physics, 24(1):1– 9, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917790105X.

Lindemuth:1973:ADI

[LK73] Irvin Lindemuth and John Killeen. Alternating direction implicit techniques for two-dimensional magnetohydrodynamic calculations. Journal of Computational Physics, 13(2):181–208, October 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173900223.

Longman:1975:DDS

[LL75] I. M. Longman and D. Levin. On the density distribution in a self-gravitating liquid sphere of constant compressibility. Journal of Computational Physics, 18(4):347–359, August 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900911.

Langeberg:1973:IBV

[LM73] John C. Langeberg and Rex B. McLellan. Interactions between vacancies in rigid disk solids. Journal of Computational Physics, 11(3):400–408, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900806.

Lomax:1974:FDN

[LM74] Harvard Lomax and E. Dale Martin. Fast direct numerical solution of the nonhomogeneous Cauchy–Riemann equations. Journal of Computational Physics, 15(1):55–80, May 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900692.

Lyons:1973:EIR

[LN73] J. D. Lyons and R. K. Nesbet. Evaluation of integrals required in scattering theory: II. exchange type integrals. Journal of Computational Physics, 11(2):166–189, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900028.

Lewis:1975:CTT

[LN75] H. R. Lewis and C. W. Nielson. A comparison of three two-dimensional electrostatic plasma simulation models. Journal of Computational Physics, 17(1):1–9, January 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900595.

Lyons:1973:MVM

[LNRY73] J. D. Lyons, R. K. Nesbet, C. C. Rankini, and A. C. Yates. Matrix variational method for electron-atom scattering. Journal of Computational Physics, 13(2):229–259, October 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900247.

Lee:1978:SMS

[LO78] W. W. Lee and H. Okuda. A simulation model for studying low-frequency microinstabilities. Journal of Computational Physics, 26(2):139–152, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900864.

LoDato:1970:NSN

[LoD70] V. A. LoDato. The numerical solution of a nonlinear master equation. Journal of Computational Physics, 6(1):105–112, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900070.

Longman:1972:NLT

[Lon72] I. M. Longman. Numerical Laplace transform inversion of a function arising in viscoelasticity. Journal of Computational Physics, 10(2):224–231, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900629.

See notes [Voj73].

Latham:1974:CPC

[LR74] W. P. Latham and Rogers W. Redding. On the calculation of the parabolic cylinder functions. Journal of Computational Physics, 16(1):66–75, September 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174901041.

Liles:1978:SIM

[LR78] D. R. Liles and Wm. H. Reed. A semi-implicit method for twophase fluid dynamics. Journal of Computational Physics, 26(3): 390–407, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900773.

Lam:1976:CDB

[LS76] D. C. L. Lam and R. B. Simpson. Centered differencing and the box scheme for diffusion convection problems. Journal of Computational Physics, 22(4):486–500, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900450.

Lapwood:1978:PEO

[LS78] E. R. Lapwood and Ryosuke Sato. The pattern of eigenfrequencies of overtones of torsional oscillations of a layered spherical shell. Journal of Computational Physics, 29(3):412–420, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901432.

Langeberg:1971:EME

[LSM71] John C. Langeberg, Zevi W. Salsburg, and Rex B. McLellan. The effect of monovacancies on the elastic properties of rigid disk solids. Journal of Computational Physics, 7(3):445–464, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171901045.

Lewis:1972:CSP

[LSW72] H. Ralph Lewis, A. Sykes, and J. A. Wesson. A comparison of some particle-in-cell plasma simulation methods. Journal of Computational Physics, 10(1):85–106, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900927.

Leboeuf:1979:MPC

[LTD79] J. N. Leboeuf, T. Tajima, and J. M. Dawson. A magnetohydrodynamic particle code for fluid simulation of plasmas. Journal of Computational Physics, 31(3):379–408, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900536.

Lutzky:1971:QAV

[Lut71] M. Lutzky. Quadratic artificial viscosity in numerical magnetic gas dynamics. Journal of Computational Physics, 7(1):172–177, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917190057X.

Lodhi:1978:SFS

[LW78] M. A. K. Lodhi and B. T. Waak. Semiempirical formulas for single-particle energies of neutrons and protons. Journal of Computational Physics, 28(1):65–81, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900475.

Lastman:1978:NSB

[LWH78] G. J. Lastman, R. A. Wentzell, and A. C. Hindmarsh. Numerical solution of a bubble cavitation problem. Journal of Computational Physics, 28(1):56–64, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900463.

Lyness:1976:AET

[Lyn76] J. N. Lyness. Applications of extrapolation techniques to multidimensional quadrature of some integrand functions with a singularity. Journal of Computational Physics, 20(3):346–364, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900875.

Macdonald:1973:FED

[Mac73] J. Ross Macdonald. Fitting experimental data. Journal of Computational Physics, 11(4):620, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901423.

Martin:1975:NRW

[Mar75] Brian Martin. Numerical representations which model properties of the solution to the diffusion equation. Journal of Computational Physics, 17(4):358–383, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900406.

Margolis:1978:TDS

[Mar78] Stephen B. Margolis. Time-dependent solution of a premixed laminar flame. Journal of Computational Physics, 27(3):410– 427, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900189.

Masson:1973:IFD

[Mas73] Bruce S. Masson. An implicit finite difference method for Eulerian fluid dynamics. Journal of Computational Physics, 12(1): 88–102, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901708.

Matthes:1970:SRP

[Mat70] Wilhelm Matthes. Simulation of relaxation processes by Monte Carlo. Journal of Computational Physics, 6(2):157–167, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900185. See comments [Bir71].

Miele:1974:STP

[MAT74] A. Miele, A. K. Aggarwal, and J. L. Tietze. Solution of twopoint boundary-value problems with Jacobian matrix characterized by large positive eigenvalues. Journal of Computational Physics, 15(2):117–133, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900801.

Mandel:1971:SRN

[MB71] Frederic Mandel and Richard J. Bearman. Some remarks on the numerical evaluation of Fourier and Fourier–Bessel transforms. Journal of Computational Physics, 7(3):637–645, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171901185.

McDonald:1975:TDS

[MB75] H. McDonald and W. R. Briley. Three-dimensional supersonic flow of a viscous or inviscid gas. Journal of Computational Physics, 19(2):150–178, October 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900868.

MacCracken:1977:TAF

[MB77] Michael C. MacCracken and Robert D. Bornstein. On the treatment of advection in flux formulations for variable grid models, with application to two models of the atmosphere. Journal of Computational Physics, 23(2):135–149, February 1977. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999177901176.

Mankoc-Borstnik:1976:MPA

[MBM76] N. Mankoc-Borstnik and M. V. Mihailović. A method for the projection of angular momentum. Journal of Computational Physics, 22(3):271–294, November 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900504.

Marsh:1972:USM

[MC72] Brent L. Marsh and Frederick A. Costello. Use of the simplex method in nonlinear programming for duct layout design. Journal of Computational Physics, 9(2):365–371, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900289.

Morihara:1973:NSV

[MC73] Hiroshi Morihara and Ralph Ta-Shun Cheng. Numerical solution of the viscous flow in the entrance region of parallel plates. Journal of Computational Physics, 11(4):550–572, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917390137X.

Mezei:1976:CPO

[MC76] Mihaly Mezei and Edwin S. Campbell. A computational procedure for obtaining the poles of a spherical harmonic of order N: Application to the multipole expansion of electrostatic interaction. Journal of Computational Physics, 20(1):110–116, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176901054.

Mezei:1978:CAP

[MC78] Mihaly Mezei and Edwin S. Campbell. Computer algorithms and programs for permanent multipole and induced dipole interaction energies and dipole vectors in crystals. Journal of Computational Physics, 29(2):297–301, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901614.

McDonald:1974:NCN

[McD74] B. E. McDonald. Numerical calculation of nonunique solutions of a two-dimensional sinh-Poisson equation. Journal of Computational Physics, 16(4):360–370, December 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917490045X.

McKee:1971:USC

[McK71] S. McKee. The use of the Schur–Cohn conditions for determining the stability condition of multi-level finite difference schemes. Journal of Computational Physics, 8(1):153–155, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900416.

McMurchie:1978:OTE

[MD78] Larry E. McMurchie and Ernest R. Davidson. One- and twoelectron integrals over Cartesian Gaussian functions. Journal of Computational Physics, 26(2):218–231, February 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917890092X.

Mallinson:1973:MFT

[MdVD73] G. D. Mallinson and G. de Vahl Davis. The method of the false transient for the solution of coupled elliptic equations. Journal of Computational Physics, 12(4):435–461, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173900971.

MacKnight:1972:NMU

[ME72] Allen K. MacKnight and Henry Eyring. Numerical methods used in the significant structure theory of liquids. Journal of Computational Physics, 9(3):584–591, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900137.

Meng:1974:SCH

[Men74] James C. S. Meng. A study of convective heat transfer in a supersonic recirculating flow region by Telenin’s method. Journal of Computational Physics, 15(3):320–344, July 1974. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999174901156.

Meyder:1975:SCE

[Mey75] R. Meyder. Solving the conservation equations in fuel rod bundles exposed to parallel flow by means of curvilinear-orthogonal coordinates. Journal of Computational Physics, 17(1):53–67, January 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900625.

Marchuk:1973:NMS

[MGRK73] G. I. Marchuk, R. G. Gordeev, V. Ya Rivkind, and B. A. Kagan. A numerical method for the solution of tidal dynamics equations and the results of its application. Journal of Computational Physics, 13(1):15–34, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901228.

Maltz:1979:VRM

[MH79] F. H. Maltz and D. L. Hitzl. Variance reduction in Monte Carlo computations using multi-dimensional Hermite polynomials. Journal of Computational Physics, 32(3):345–376, September 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901505.

Miller:1970:VIW

[Mil70a] Edmund K. Miller. A variable interval width quadrature technique based on Romberg’s method. Journal of Computational Physics, 5(2):265–279, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917090063X.

Miller:1970:GBC

[Mil70b] Richard H. Miller. Gravitational N-body calculation in a discrete phase space. Journal of Computational Physics, 6(3):449– 472, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900422.

Miller:1971:ETL

[Mil71a] Bruce N. Miller. Electromagnetic theory of light scattering from an inhomogeneous fluid. Journal of Computational Physics, 7(3): 576–591, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171901136.

Miller:1971:ESN

[Mil71b] R. H. Miller. Experimental studies of the numerical stability of the gravitational n-body problem. Journal of Computational Physics, 8(3):449–463, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900234.

Miller:1971:PIR

[Mil71c] R. H. Miller. Partial iterative refinements. Journal of Computational Physics, 8(3):464–471, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900246.

Miller:1974:EAS

[Mil74] Kenneth J. Miller. Evaluation of atomic scattering integrals by numerical techniques. Journal of Computational Physics, 14(3): 269–285, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900539.

Miller:1976:VDG

[Mil76] R. H. Miller. Validity of disk galaxy simulations. Journal of Computational Physics, 21(4):400–437, August 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900383.

Mirin:1977:RTF

[MKMR77] A. A. Mirin, J. Killeen, K. D. Marx, and M. E. Rensink. A radial transport/Fokker–Planck model for a Tokamak plasma. Journal of Computational Physics, 23(1):23–41, January 1977. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999177900857.

Media:1977:CNM

[MKS77] R. Media, R. B. Kellogg, and J. L. Stephenson. Comparison of numerical methods for renal network flows. Journal of Computational Physics, 23(1):53–62, January 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900870.

Miller:1971:RSF

[ML71] C. Grant Miller and Russell D. Larsen. Relative stabilities of a family of plastically-deformed lattice close packings of rigid disks. Journal of Computational Physics, 7(3):465–488, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171901057.

Moore:1973:NAS

[ML73] R. A. Moore and C. F. Liu. A numerical approach for the solution of the radial Schrödinger equation with atomic-like potentials. Journal of Computational Physics, 12(1):103–111, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917390171X.

Matte:1977:SDP

[ML77] Jean-Pierre Matte and Gaétan Lafrance. Solution of the discrete Poisson equation with complicated boundaries. Journal of Computational Physics, 23(1):86–91, January 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900912.

McGuire:1973:CSO

[MM73] G. R. McGuire and J. Ll Morris. A class of second-order accurate methods for the solution of systems of conservation laws. Journal of Computational Physics, 11(4):531–549, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901368.

McGuire:1974:CIS

[MM74] G. R. McGuire and J. Ll Morris. A class of implicit, secondorder accurate, dissipative schemes for solving systems of conservation laws. Journal of Computational Physics, 14(2):126– 147, February 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900096.

Morris:1973:HMA

[MN73] John Ll Morris and Ian F. Nicoll. Hopscotch methods for an anisotropic thermal print head problem. Journal of Computational Physics, 13(3):316–337, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900399.

Morris:1974:ECH

[MN74] John Ll Morris and Ian F. Nicoll. The efficient computation of the heat distribution in a 5 × 5 matrix thermal print head. Journal of Computational Physics, 15(2):188–199, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900837.

Mock:1976:TDN

[Moc76] M. S. Mock. Time discretization of a nonlinear initial value problem. Journal of Computational Physics, 21(1):20–37, May 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900188.

Modrak:1975:LTE

[Mod75] P. Modrak. On the low-temperature expansion of the thermodynamical integral for fcc lattice. Journal of Computational Physics, 18(3):342–346, July 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900078.

Morawrrz:1970:NAP

[Mor70a] Hans Morawrrz. A numerical approach to principal value integrals in dispersion relations. Journal of Computational Physics, 6(1): 120–123, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900094.

Morris:1970:NSH

[Mor70b] John L. L. Morris. On the numerical solution of a heat equation associated with a thermal print-head. Journal of Computational Physics, 5(2):208–228, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900604.

Morris:1971:NSH

[Mor71] John L. L. Morris. On the numerical solution of a heat equation associated with a thermal print head. II. Journal of Computational Physics, 7(1):102–119, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900520.

Morrow:1978:ANE

[Mor78] Richard A. Morrow. Algorithm for the numerical evaluation of the n-particle relativistic phase space integral in invariant variables. Journal of Computational Physics, 29(2):245–269, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901559.

Moreau:1979:VPS

[Mor79] W. R. Moreau. A vector perturbation series based upon a representation in a finite Banach space. Journal of Computational Physics, 31(2):240–264, May 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900718.

Moseley:1970:CDC

[Mos70] P. G. N. Moseley. Computation of dissociation constants and equivalent conductances at infinite dilution. Journal of Computational Physics, 6(1):150–151, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917090015X.

Madala:1977:SIN

[MP77] Rangarao V. Madala and Steve A. Piacseki. A semi-implicit numerical model for baroclinic oceans. Journal of Computational Physics, 23(2):167–178, February 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917790119X.

Miller:1973:IDE

[MPB73] E. K. Miller, A. J. Poggio, and G. J. Burke. An integrodifferential equation technique for the time-domain analysis of thin wire structures. I. The numerical method. Journal of Computational Physics, 12(1):24–48, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901678.

Marinov:1975:VAR

[MPS75] M. S. Marinov, V. S. Popov, and V. L. Stolin. Variational approach to the relativistic two-center problem: Critical internuclear distance. Journal of Computational Physics, 19(3):241–256, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900753.

Melnyk:1971:SVC

[MR71] T. W. Melnyk and J. S. Rowlinson. The statistics of the volumes covered by systems of penetrating spheres. Journal of Computational Physics, 7(3):385–393, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900994.

Metcalf:1973:SEM

[MR73] M. Metcalf and M. Regler. Solution of the equations of motion of a charged particle in a stationary magnetic field. Journal of Computational Physics, 11(2):240–249, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900053.

Minicozzi:1970:DIE

[MS70a] William P. Minicozzi and Michael T. Stroot. On the determination of interaction energy functions II. Crystalline formic acid. Journal of Computational Physics, 6(1):95–104, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900069.

Moretti:1970:NAV

[MS70b] Gino Moretti and Manuel D. Salas. Numerical analysis of viscous one-dimensional flows. Journal of Computational Physics, 5(3): 487–506, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900768.

Maria-Sube:1974:NSF

[MS74] Roland Maria-Sube. Numerical solution to free surface axisymmetric rotational flows. Journal of Computational Physics, 16(2):127– 149, October 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901089.

Milinazzo:1977:CLR

[MS77] F. Milinazzo and P. G. Saffman. The calculation of large Reynolds number two-dimensional flow using discrete vortices with random walk. Journal of Computational Physics, 23(4):380–392, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900699. See comment [Cho78a] and re-

ply [Saf78].

Mason:1978:SCM

[MS78a] P. J. Mason and R. I. Sykes. A simple Cartesian model of boundary layer flow over topography. Journal of Computational Physics, 28(2):198–210, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900347.

Moler:1978:HFA

[MS78b] Cleve B. Moler and G. W. Stewart. On the Householder–Fox algorithm for decomposing a projection. Journal of Computational Physics, 28(1):82–91, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900487.

Mejia:1979:NSM

[MS79a] R. Mejia and J. L. Stephenson. Numerical solution of multinephron kidney equations. Journal of Computational Physics, 32(2):235– 246, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917990130X.

Milgram:1979:PCP

[MS79b] Michael S. Milgram and Kenneth N. Sly. On the properties of collision probability integrals in annular geometry. II. Evaluation. Journal of Computational Physics, 33(3):417–424, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901670.

Murch:1978:PGE

[MT78] G. E. Murch and R. J. Thorn. Petit and grand ensemble Monte Carlo calculations of the thermodynamics of the lattice gas. Journal of Computational Physics, 29(2):237–244, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901547.

Mudrick:1976:USD

[Mud76] Stephen Mudrick. On the use of a scale-dependent filter in channel model integrations. Journal of Computational Physics, 20(1):33– 49, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176901005.

Maruhn:1976:RNS

[MWW76] J. A. Maruhn, T. A. Welton, and C. Y. Wong. Remarks on the numerical solution of Poisson’s equation for isolated charge distributions. Journal of Computational Physics, 20(3):326–335, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900851.

Myers:1971:SOV

[MY71] R. M. Myers and M. Yanowitch. Small oscillations of a viscous isothermal atmosphere. Journal of Computational Physics, 8(2):241–257, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900052.

Mecca:1973:MRB

[MY73] S. J. Mecca and P. F. Yergin. A method for reducing bremsstrahlung-produced photoneutron spectra to partial (γ, n) cross sections. Journal of Computational Physics, 13(3):338–347, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900405.

Nardo:1971:ADI

[NC71] C. T. Nardo and R. J. Cresci. An alternating directional implicit scheme for three-dimensional hypersonic flows. Journal of Computational Physics, 8(2):268–284, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900076.

Nesbet:1965:ADL

[Nes65] R. K. Nesbet. Algorithm for diagonalization of large matrices. Journal of Chemical Physics, 43(1):311–312, ???? 1965. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). See modification [Sha70b].

Nesbet:1971:ELH

[Nes71] R. K. Nesbet. Evaluation of m^{†}(h − ϵ)^{−1}m for a large Hermitian matrix h. Journal of Computational Physics, 8(3):483–488, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917190026X.

Nettleton:1971:OBC

[Net71] R. E. Nettleton. Optical birefringence in cubic crystals. Journal of Computational Physics, 7(3):603–605, June 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917190115X.

Nelson:1972:UDC

[NG72] Paul Nelson, Jr. and C. A. Giles. A useful device for certain boundary-value problems. Journal of Computational Physics, 10(2):374–378, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900745.

Ng:1974:ANF

[Ng74] Kin-Chue Ng. On the accuracy of numerical Fourier transforms. Journal of Computational Physics, 16(4):396–400, December 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900497.

Nichols:1971:IFS

[NH71] B. D. Nichols and C. W. Hirt. Improved free surface boundary conditions for numerical incompressible-flow calculations. Journal of Computational Physics, 8(3):434–448, December 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999171900222.

Nichols:1973:CTD

[NH73] B. D. Nichols and C. W. Hirt. Calculating three-dimensional free surface flows in the vicinity of submerged and exposed structures. Journal of Computational Physics, 12(2):234–246, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999173800130.

Nicolaides:1975:MGR

[Nic75] R. A. Nicolaides. On multiple grid and related techniques for solving discrete elliptic systems. Journal of Computational Physics, 19(4):418–431, December 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900728.

Nisbet:1972:ACN

[Nis72] Roger M. Nisbet. Acceleration of the convergence in Nesbet’s algorithm for eigenvalues and eigenvectors of large matrices. Journal of Computational Physics, 10(3):614–619, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900551.

Nallasamy:1974:NSQ

[NP74] M. Nallasamy and K. Krishna Prasad. Numerical studies on quasilinear and linear elliptic equations. Journal of Computational Physics, 15(4):429–448, August 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900722.

Nakayama:1971:NMA

[NR71] P. I. Nakayama and N. C. Romero. Numerical method for almost three-dimensional incompressible fluid flow and a simple internal obstacle treatment. Journal of Computational Physics, 8(2):230– 240, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900040.

Ng:1979:IVM

[NR79a] B. S. Ng and W. H. Reid. An initial value method for eigenvalue problems using compound matrices. Journal of Computational Physics, 30(1):125–136, January 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900913.

Ng:1979:NML

[NR79b] B. S. Ng and W. H. Reid. A numerical method for linear two-point boundary-value problems using compound matrices. Journal of Computational Physics, 33(1):70–85, October 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900287.

Neece:1971:TPC

[NRH71] G. A. Neece, F. J. Rogers, and W. G. Hoover. Thermodynamic properties of compressed solid hydrogen. Journal of Computational Physics, 7(3):621–636, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171901173.

Naitou:1979:BCS

[NTK79] H. Naitou, S. Tokuda, and T. Kamimura. On boundary conditions for a simulation plasma in a magnetic field. Journal of Computational Physics, 33(1):86–101, October 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900299.

Okamoto:1978:MSI

[OA78] Masao Okamoto and Tsuneo Amano. A method to solve the impurity diffusion equation with ionization and recombination source terms. Journal of Computational Physics, 26(1):80–86, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901006.

OBrien:1969:DNS

[O’B69] James J. O’Brien. A direct numerical solution to a onedimensional Helmholtz equation. Journal of Computational Physics, 3(4):544–546, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916990062X.

See comments [Yee70].

ORourke:1979:TST

[OB79] P. J. O’Rourke and F. V. Bracco. Two scaling transformations for the numerical computation of multidimensional unsteady laminar flames. Journal of Computational Physics, 33(2):185–203, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900159.

Oberoi:1972:AEI

[OCS72] R. S. Oberoi, J. Callaway, and G. J. Seiler. Analytic evaluation of integrals in variational calculations of scattering theory. Journal of Computational Physics, 10(3):466–474, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900472.

Ohmura:1971:SCC

[OFM71] Takashi Ohmura, Hiroshi Fujii, and Eijiro Maeda. A symbolic calculation code for nuclear three-body problem. Journal of Computational Physics, 7(1):61–82, February 1971. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900490.

Orszag:1974:LED

[OJ74] Steven A. Orszag and Lance W. Jayne. Local errors of difference approximations to hyperbolic equations. Journal of Computational Physics, 14(1):93–103, January 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900072.

See notes [Chi74].

Okuda:1972:NNI

[Oku72] Hideo Okuda. Nonphysical noises and instabilities in plasma simulation due to a spatial grid. Journal of Computational Physics, 10(3):475–486, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900484.

Oberoi:1973:NAF

[ON73] R. S. Oberoi and R. K. Nesbet. Numerical asymptotic functions in variational scattering theory. Journal of Computational Physics, 12(4):526–533, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901034.

Oreg:1973:CCE

[Ore73] J. Oreg. Calculation and checking of electrostatic energy matrices by computer. Journal of Computational Physics, 11(2):153– 165, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900016.

Orlanski:1976:SBC

[Orl76] I. Orlanski. A simple boundary condition for unbounded hyperbolic flows. Journal of Computational Physics, 21(3):251– 269, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900231.

OHara:1970:ECT

[OS70] H. O’Hara and Francis J. Smith. The efficient calculation of the transport properties of a dilute gas to a prescribed accuracy. Journal of Computational Physics, 5(2):328–344, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900653.

Overton:1974:UVD

[OS74] W. C. Overton, Jr. and A. F. Schuch. A uniform vector distribution for accurate spatial integration. Journal of Computational Physics, 14(1):59–83, January 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900059.

Odaka:1978:SPS

[OU78] T. Odaka and T. Usami. Some properties of spheroidal modes of a homogeneous elastic sphere with special reference to radial dependence of displacement. Journal of Computational Physics, 29(3):431–445, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901456.

Owen:1971:MER

[Owe71] L. W. Owen. The multipole expansion of realistic effective interactions in nuclei. Journal of Computational Physics, 7(1):182– 186, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900593.

Pachner:1974:NIE

[Pac74] J. Pachner. Numerical integration of exact time-dependent Einstein equations with axial symmetry. I. Journal of Computational Physics, 15(3):385–412, July 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174901193.

Pantis:1975:EIO

[Pan75] G. Pantis. The evaluation of integrals with oscillatory integrands. Journal of Computational Physics, 17(2):229–233, February 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900388.

Paterson:1971:CET

[Pat71] William L. Paterson. Computation of the exponential trap population integral of glow curve theory. Journal of Computational Physics, 7(1):187–190, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917190060X.

Patel:1978:KVS

[Pat78] V. A. Patel. Kármán vortex street behind a circular cylinder by the series truncation method. Journal of Computational Physics, 28 (1):14–42, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917890044X.

Poston:1975:GAC

[PB75] T. Poston and A. B. Budgor. A geometrical approach to calculating the energy and frequency spectra of crystals. Journal of Computational Physics, 19(1):1–28, September 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175901138.

Piessens:1976:NSI

[PB76a] Robert Piessens and Maria Branders. Numerical solution of integral equations of mathematical physics, using Chebyshev polynomials. Journal of Computational Physics, 21(2):178–196, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900103.

Popinski:1976:IFE

[PB76b] Z. Popinski and A. J. Baker. An implicit finite-element algorithm for the boundary layer equations. Journal of Computational Physics, 21(1):55–84, May 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900206.

Pifssens:1976:GQF

[PCJ76] R. Pifssens, M. M. Cyiawla, and N. Jayarajan. Gaussian quadrature formulas for the numerical calculation of integrals with logarithmic singularity. Journal of Computational Physics, 21(3): 356–360, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900346.

Perkins:1975:EFE

[Per75] J. F. Perkins. Evaluation of a four-electron atomic integral. Journal of Computational Physics, 17(4):434–436, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900467.

Pert:1978:ACI

[Per78] G. J. Pert. Algorithms for the calculation of ionization in rapidly changing plasmas. Journal of Computational Physics, 27(2):241– 255, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900074.

Peskin:1972:FPA

[Pes72] Charles S. Peskin. Flow patterns around heart valves: A numerical method. Journal of Computational Physics, 10(2):252– 271, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900654.

Peskin:1977:NAB

[Pes77] Charles S. Peskin. Numerical analysis of blood flow in the heart. Journal of Computational Physics, 25(3):220–252, November 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177901000.

Peverley:1971:SAE

[Pev71] J. Roger Peverley. A simplified approach to the evaluation of lattice sums. Journal of Computational Physics, 7(1):83–88, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900507.

Pexton:1976:ASI

[Pex76] Robert L. Pexton. Application of Salzer’s inverse Laplace transform algorithm. Journal of Computational Physics, 20(4):492– 494, April 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900954.

Pendergast:1974:TAE

[PF74] Phil Pendergast and William H. Fink. A thorough analysis and exposition of the four-index transformation. Journal of Computational Physics, 14(3):286–300, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900540.

Pomraning:1979:TDI

[PF79] G. C. Pomraning and G. M. Foglesong. Transport-diffusion interfaces in radiative transfer. Journal of Computational Physics, 32(3):420–436, September 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901530.

Pfeiffer:1979:GMF

[Pfe79] Wayne Pfeiffer. A graphical method for finding complex roots and its application to plasma physics problems. Journal of Computational Physics, 33(3):397–404, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901645.

Prasad:1970:NIN

[PH70] K. Krishna Prasad and R. G. Hering. Numerical integration of a nonlinear, singular integro-partial differential equation. Journal of Computational Physics, 6(3):406–416, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900392.

Pendergast:1978:PTA

[PH78a] Phil Pendergast and Edward F. Hayes. A partial transformation for application to perturbation theory configuration interaction. Journal of Computational Physics, 26(2):236–242, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900943.

Pople:1978:CER

[PH78b] John A. Pople and Warren J. Hehre. Computation of electron repulsion integrals involving contracted Gaussian basis functions. Journal of Computational Physics, 27(2):161–168, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900013.

Petravic:1979:IAW

[PKP79] M. Petravic and G. Kuo-Petravic. An ILUCG algorithm which minimizes in the Euclidean norm. Journal of Computational Physics, 32(2):263–269, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901335.

Petravic:1972:AOS

[PKPR72] M. Petravic, G. Kuo-Petravic, and K. V. Roberts. Automatic optimization of symbolic Algol programs. I. General principles. Journal of Computational Physics, 10(3):503–533, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900502.

Proriol:1970:RCA

[PL70] J. Proriol and R. Laverriere. Relativistic calculations of angular correlations of particles issued from an interaction-free resonance decay. Journal of Computational Physics, 6(2):329–332, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900318.

Panikker:1975:FPI

[PL75] P. K. G. Panikker and Z. Lavan. Flow past impulsively started bodies using Green’s functions. Journal of Computational Physics, 18 (1):46–65, May 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175901011.

Poggio:1973:IDE

[PMB73] A. J. Poggio, E. K. Miller, and G. J. Burke. An integrodifferential equation technique for the time-domain analysis of thin-wire structures. II. Numerical results. Journal of Computational Physics, 12(2):210–233, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999173800129.

Pope:1978:CTR

[Pop78] S. B. Pope. The calculation of turbulent recirculating flows in general orthogonal coordinates. Journal of Computational Physics, 26(2):197–217, February 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900918.

Posener:1973:FEM

[Pos73] D. W. Posener. Fast ESR matrix diagonalizatio. Journal of Computational Physics, 11(3):326–332, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900764.

Ponomarev:1973:CAN

[PPP73] L. I. Ponomarev, I. V. Puzynin, and T. P. Puzynina. Continuous analog of Newton’s method as applied to the calculation of the binding energy of mesic molecules. Journal of Computational Physics, 13(1):1–14, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901216.

Ponomarev:1976:CAN

[PPP76] L. I. Ponomarev, I. V. Puzynin, and T. P. Puzynina. Continuous analog of Newton’s method for the calculation of quasibound states of hydrogen μ-molecules. Journal of Computational Physics, 22(1):125–130, September 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176901121.

Pirkle:1974:NPP

[PPS74] J. C. Pirkle, Jr., T. O. Poehler, and V. G. Sigillito. Numerical procedure for a pulsed DF–CO_{2} transfer laser. Journal of Computational Physics, 15(2):293–298, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900904.

Palumbo:1972:STD

[PR72] Dominic J. Palumbo and Ephraim L. Rubin. Solution of the twodimensional, unsteady, compressible Navier–Stokes equations using a second-order accurate numerical scheme. Journal of Computational Physics, 9(3):466–495, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917290006X.

Pracht:1971:NMC

[Pra71] William E. Pracht. A numerical method for calculating transient creep flows. Journal of Computational Physics, 7(1):46– 60, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900489.

Pracht:1975:CTD

[Pra75] William E. Pracht. Calculating three-dimensional fluid flows at all speeds with an Eulerian–Lagrangian computing mesh. Journal of Computational Physics, 17(2):132–159, February 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900339.

Preussner:1978:TDI

[Pre78] Paul-Rolf Preussner. Time-dependent ionization. Journal of Computational Physics, 29(2):283–288, November 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178901584.

Palmer:1970:USD

[PS70] James T. Palmer and Donald R. Smith. On an unconditionally stable difference scheme for the multifrequency Eddington equations. Journal of Computational Physics, 6(3):356–366, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900355.

Pirkle:1972:VAL

[PS72] J. C. Pirkle, Jr. and V. G. Sigillito. A variational approach to low Peclet number heat transfer in laminar flow. Journal of Computational Physics, 9(2):207–221, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900198.

Payne:1973:VMH

[PS73] G. L. Payne and L. Schlessinger. Variational methods in a Harmonic Oscillator basis. Journal of Computational Physics, 13(2):266–287, October 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900260.

Ponomarev:1976:WFC

[PS76] L. I. Ponomarev and L. N. Somov. The wave functions of continuum for the two-center problem in quantum mechanics. Journal of Computational Physics, 20(2):183–195, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900632.

Potter:1973:CDO

[PT73] David E. Potter and G. H. Tuttle. The construction of discrete orthogonal coordinates. Journal of Computational Physics, 13(4):483–501, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900296.

Parmentier:1975:KCV

[PT75] E. M. Parmentier and K. E. Torrance. Kinematically consistent velocity fields for hydrodynamic calculations in curvilinear coordinates. Journal of Computational Physics, 19(4):404–417, December 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900716.

Piacsek:1970:CPC

[PW70] Steve A. Piacsek and Gareth P. Williams. Conservation properties of convection difference schemes. Journal of Computational Physics, 6(3):392–405, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900380.

Quentrec:1973:NMS

[QB73] B. Quentrec and C. Brot. New method for searching for neighbors in molecular dynamics computations. Journal of Computational Physics, 13(3):430–432, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900466.

Quon:1976:DIB

[Quo76a] Charles Quon. Diffusively induced boundary layers in a tilted square cavity: A numerical study. Journal of Computational Physics, 22(4):459–485, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900449.

Quon:1976:MSF

[Quo76b] Charles Quon. A mixed spectral and finite difference model to study baroclinic annulus waves. Journal of Computational Physics, 20(4):442–479, April 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900930.

Raffenetti:1979:SCR

[Raf79] Richard C. Raffenetti. A simultaneous coordinate relaxation algorithm for large, sparce matrix eigenvalue problems. Journal of Computational Physics, 32(3):403–419, September 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999179901529.

Rajan:1973:PAH

[Raj73] S. Rajan. A parallel algorithm for high subsonic compressible flow over a circular cylinder. Journal of Computational Physics, 12(4): 534–552, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901046.

Roberts:1971:SPD

[RB71] K. V. Roberts and J. P. Boris. The solution of partial differential equations using a symbolic style of Algol. Journal of Computational Physics, 8(1):83–105, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900374.

Renken:1972:UNI

[RB72] James H. Renken and Frank Biggs. Use of numerically inverted Laplace transforms in time-dependent transport calculations. Journal of Computational Physics, 9(2):318–338, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900265.

Reif:1977:NSF

[RB77] John Reif and Richard Barakat. Numerical solution of the Fokker–Planck equation via Chebyschev polynomial approximations with reference to first passage time probability density functions. Journal of Computational Physics, 23(4):425–445, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900729.

Ryckaert:1977:NIC

[RCB77] Jean-Paul Ryckaert, Giovanni Ciccotti, and Herman J. C. Berendsen. Numerical integration of the Cartesian equations of motion of a system with constraints: molecular dynamics of nalkanes. Journal of Computational Physics, 23(3):327–341, March 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900985.

Rathmann:1975:SCE

[RD75] C. E. Rathmann and J. Denavit. Simulation of collisional effects in plasmas. Journal of Computational Physics, 18(2):165– 187, June 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900273.

Riehl:1974:CPD

[RDW74] J. P. Riehl, D. J. Diestler, and Albert F. Wagner. Comparison of perturbation and direct-numerical-integration techniques for the calculation of phase shifts for elastic scattering. Journal of Computational Physics, 15(2):212–225, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900850.

Read:1970:CMS

[Rea70] F. H. Read. A collocation method for solving Laplace’s equation. Journal of Computational Physics, 6(3):527–532, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917090046X.

Reilly:1973:RSI

[Rei73] Edwin D. Reilly, Jr. Resonant scattering from inhomogeneous nonspherical targets. Journal of Computational Physics, 11(4):463– 492, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901332.

Reid:1974:IRH

[Rei74a] Charles E. Reid. Integrals of the resolvent of the hydrogenic Hamiltonian over STO basis functions. Journal of Computational Physics, 15(4):522–528, August 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900771.

Reis:1974:NIF

[Rei74b] M. F. Reis. Numerical investigation of the Friedrichs model dispersion relation. Journal of Computational Physics, 15(3):375– 384, July 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901181.

Rudd:1971:ESP

[RF71] W. G. Rudd and H. L. Frisch. The equation of state of parallel hard squares. Journal of Computational Physics, 7(3):394– 402, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171901008.

Rosina:1975:VCR

[RG75] M. Rosina and C. Garrod. The variational calculation of reduced density matrices. Journal of Computational Physics, 18(3):300– 310, July 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900042.

Riddell:1979:BDS

[Rid79a] R. J. Riddell, Jr. Boundary-distribution solution of the Helmholtz equation for a region with corners. Journal of Computational Physics, 31(1):21–41, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900603.

Riddell:1979:NSH

[Rid79b] Robert J. Riddell, Jr. Numerical solution of the Helmholtz equation for two-dimensional polygonal regions. Journal of Computational Physics, 31(1):42–59, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900615.

Rigal:1979:COS

[Rig79] Alain Rigal. Convergence and optimization of successive overrelaxation for linear systems of equations with complex eigenvalues. Journal of Computational Physics, 32(1):10–23, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901384.

Ritter:1970:LMC

[Rit70] Z. W. Ritter. A Lagrangian method for calculating problems in meteorology. Journal of Computational Physics, 6(3):510–526, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900458.

Rubin:1977:PIM

[RK77] S. G. Rubin and P. K. Khosla. Polynomial interpolation methods for viscous flow calculations. Journal of Computational Physics, 24(3):217–244, July 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900365.

Rijbin:1972:NMT

[RL72] Stanley G. Rijbin and Tony C. Lin. A numerical method for three-dimensional viscous flow: Application to the hypersonic leading edge. Journal of Computational Physics, 9(2):339–364, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900277.

Redding:1976:CPC

[RL76] R. W. Redding and W. P. Latham. On the calculation of the parabolic cylinder functions. II. The function V (a, x). Journal of Computational Physics, 20(2):256–258, February 1976. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999176900711.

Robinson:1978:NTM

[RLT78] Enders A. Robinson, Dan Loewenthal, and Sven Treitel. Numerical testing of minimum-delay, positive-real, and positive-definite digital filters. Journal of Computational Physics, 29(3):421–430, December 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901444.

Rowley:1975:MCG

[RNP75] L. A. Rowley, D. Nicholson, and N. G. Parsonage. Monte Carlo grand canonical ensemble calculation in a gas-liquid transition region for 12-6 argon. Journal of Computational Physics, 17(4): 401–414, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917590042X.

Rowley:1978:LRC

[RNP78] L. A. Rowley, D. Nicholson, and N. G. Parsonage. Long-range corrections to grand canonical ensemble Monte Carlo calculations for adsorption systems. Journal of Computational Physics, 26(1):66– 79, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900992.

Roache:1970:SCC

[Roa70] Patrick J. Roache. Sufficiency conditions for a commonly used downstream boundary condition on stream function. Journal of Computational Physics, 6(2):317–321, October 1970. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999170900288.

Roache:1972:AV

[Roa72] Patrick J. Roache. On artificial viscosity. Journal of Computational Physics, 10(2):169–184, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900587.

Roache:1978:PSF

[Roa78] Patrick J. Roache. A pseudo-spectral FFT technique for nonperiodic problems. Journal of Computational Physics, 27(2):204– 220, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900050.

Rosen:1971:CTC

[Ros71] Bernard Rosen. Canonical transformation by computer (Foldy– Wouthuysen transformation). Journal of Computational Physics, 8(3):369–381, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900179.

Rosen:1974:ADD

[Ros74] Bernard Rosen. Automation of the derivation of dispersion relations. I. Cold plasma case. Journal of Computational Physics, 15 (1):98–115, May 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900710.

Rosen:1976:ADD

[Ros76a] Bernard Rosen. Automation of the derivation of dispersion relations: II. Higher order dielectric functions in hot plasmas. Journal of Computational Physics, 20(1):22–32, January 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900991.

Rosenbaum:1976:CPN

[Ros76b] Jerrold S. Rosenbaum. Conservation properties of numerical integration methods for systems of ordinary differential equations. Journal of Computational Physics, 20(3):259–267, March 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900802.

Ross:1977:SSS

[Ros77] Robert T. Ross. Steady state of stiff linear kinetic systems by a Markov matrix method. Journal of Computational Physics, 24(1): 70–80, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177901103.

Rehm:1970:MCE

[RR70] R. G. Rehm and J. W. Rich. Monte Carlo evaluation of the temperature dependence of molecular transport integrals. Journal of Computational Physics, 6(3):339–355, December 1970. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999170900343.

Roberts:1972:STT

[RS72a] S. M. Roberts and J. S. Shipman. Solution of Troesch’s two-point boundary value problem by a combination of techniques. Journal of Computational Physics, 10(2):232–241, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900630.

Robinson:1972:CFS

[RS72b] Edward L. Robinson, Jr. and Henry A. Scarton. CONTOR: A FORTRAN subroutine to plot smooth contours of a singlevalued arbitrary three-dimensional surface. Journal of Computational Physics, 10(2):242–251, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900642.

Roberts:1974:IPT

[RS74] S. M. Roberts and J. S. Shipman. An iteration perturbation technique. Journal of Computational Physics, 16(3):285–297, November 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900965.

Roberts:1976:CFS

[RS76] S. M. Roberts and J. S. Shipman. On the closed form solution of Troesch’s problem. Journal of Computational Physics, 21(3): 291–304, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900267.

Rudd:1970:MED

[RSM70] W. G. Rudd, Z. W. Salsburg, and L. M. Masinter. Methods of evaluating N-dimensional integrals with polytope bounds. Journal of Computational Physics, 5(1):125–138, February 1970. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999170900562.

Ramshaw:1976:NTL

[RT76] John D. Ramshaw and John A. Trapp. A numerical technique for low-speed homogeneous two-phase flow with sharp interfaces. Journal of Computational Physics, 21(4):438–453, August 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900395.

Rankin:1979:CWF

[RT79a] J. Rankin and W. R. Thorson. Continuum wave functions for the two-center, one-electron system. Journal of Computational Physics, 32(3):437–446, September 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901542.

Rosen:1979:ACP

[RT79b] Bernard Rosen and Herbert Tesser. Automated calculation of parametric instabilities in fluid plasmas. Journal of Computational Physics, 30(2):238–249, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901013.

Rune:1975:IEA

[Run75] Axel Rune. Iterative eigenvalue algorithms based on convergent splittings. Journal of Computational Physics, 19(1):110–120, September 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175901199.

Rusanov:1970:DST

[Rus70] V. V. Rusanov. On difference schemes of third order accuracy for nonlinear hyperbolic systems. Journal of Computational Physics, 5(3):507–516, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917090077X.

Rathmann:1978:LTS

[RVD78] C. E. Rathmann, J. L. Vomvoridis, and J. Denavit. Longtime-scale simulation of resonant particle effects in Langmuir and Whistler waves. Journal of Computational Physics, 26(3):408– 442, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900785.

Rybicki:1970:NCD

[Ryb70] George B. Rybicki. A note on the computation of diffuse reflection functions for spherical shells. Journal of Computational Physics, 6(1):131–135, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900112.

Roseman:1975:NTN

[RZ75] Joseph Roseman and Gideon Zwas. Nonlinear transformations and the numerical treatment of shocks. Journal of Computational Physics, 19(2):229–235, October 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900893.

Sandford:1973:TDI

[SA73] M. T. Sandford II and R. C. Anderson. Two-dimensional implicit radiation hydrodynamics. Journal of Computational Physics, 13(1):130–157, September 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901307.

Saffman:1978:RCJ

[Saf78] P. G. Saffman. Reply to comment by A. J. Chorin. Journal of Computational Physics, 26(3):455–456, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900839.

See [Cho78a].

Schulten:1979:AEC

[SAG79] Z. Schulten, D. G. M. Anderson, and Roy G. Gordon. An algorithm for the evaluation of the complex Airy functions. Journal of Computational Physics, 31(1):60–75, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900627.

Sandford:1975:IIR

[SAHK75] M. T. Sandford II, R. C. Anderson, H. G. Horak, and J. W. Kodis. Improved, implicit radiation hydrodynamics. Journal of Computational Physics, 19(3):280–296, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900789.

Shankar:1975:NSS

[SAK75] V. Shankar, D. Anderson, and P. Kutler. Numerical solutions for supersonic corner flow. Journal of Computational Physics, 17(2):160–180, February 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900340.

Salzer:1975:NLT

[Sal75] Herbert E. Salzer. Note on Laplace transforms of osculatory and hyperosculatory interpolation polynomials. Journal of Computational Physics, 17(4):420–424, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900443.

Salzer:1976:ILT

[Sal76] Herbert E. Salzer. Inverse Laplace transforms of osculatory and hyperosculatory interpolation polynomials. Journal of Computational Physics, 20(4):480–491, April 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900942.

Salzer:1977:VCE

[Sal77] Herbert E. Salzer. A variable coefficient extension of a formula for series conversion. Journal of Computational Physics, 23(1):82– 85, January 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900900.

Sameh:1971:DVA

[Sam71] Ahmed H. Sameh. A discrete-variable approach for elasticplastic wave motions in layered solids. Journal of Computational Physics, 8(3):343–368, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900167.

Sasaki:1976:AGF

[Sas76a] Tateaki Sasaki. Automatic generation of Feynman graphs in QED. Journal of Computational Physics, 22(2):189–214, October 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900759.

Sasaki:1976:VDF

[Sas76b] Yoshi Kazu Sasaki. Variational design of finite-difference schemes for initial value problems with an integral invariant. Journal of Computational Physics, 21(3):270–278, July 1976. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999176900243.

Stephan:1973:CSM

[SB73] Brian G. Stephan and Charles J. Bridgman. Compton scattering from a Maxwellian distribution of electrons in the discrete S_{n} method. Journal of Computational Physics, 11(2):269–295, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900077.

Smith:1975:FES

[SB75] S. L. Smith and C. A. Brebbia. Finite-element solution of Navier–Stokes equations for transient two-dimensional incompressible flow. Journal of Computational Physics, 17(3):235–245, March 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900522.

Sloan:1979:NRM

[SB79] D. M. Sloan and J. S. Bramley. Note on the Riccati method for differential eigenvalue problems of odd order. Journal of Computational Physics, 30(2):296–305, February 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901050.

Sloan:1975:NAN

[SBD75] Ian H. Sloan, B. J. Burn, and N. Datyner. A new approach to the numerical solution of integral equations. Journal of Computational Physics, 18(1):92–105, May 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175901047.

Shavitt:1973:ICS

[SBPH73] I. Shavitt, C. F. Bender, A. Pipano, and R. P. Hosteny. The iterative calculation of several of the lowest or highest eigenvalues and corresponding eigenvectors of very large symmetric matrices. Journal of Computational Physics, 11(1):90–108, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901496.

Shenker:1972:NSG

[SC72] D. Shenker and R. Chen. Numerical solution of the glow curve differential equations. Journal of Computational Physics, 10(2):272– 283, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900666.

Surmont:1973:NSN

[SC73] J. Surmont and Y. M. Chen. Numerical solutions of a nonlinear radiative transfer equation with inadequate data. Journal of Computational Physics, 13(2):288–302, October 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900272.

Shubin:1979:GMC

[SC79] Gregory Shubin and Sin-I Cheng. Gasdynamic modeling and computational accuracy. Journal of Computational Physics, 32(1): 39–55, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901402.

Scarton:1971:DPF

[Sca71] Henry A. Scarton. Double precision FORTRAN subroutines to compute both ordinary and modified Bessel functions of the first kind and of integer order with arbitrary complex argument: J_{n}(x+ jy) and I_{n}(x + jy). Journal of Computational Physics, 8(2):295– 299, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900106.

Scarton:1973:ME

[Sca73] Henry A. Scarton. The method of eigenvalleys. Journal of Computational Physics, 11(1):1–14, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901447.

Schaefer:1970:DMS

[Sch70] Henry F. Schaefer III. Diatomic molecule symmetry eigenfunctions by direct diagonalization. Journal of Computational Physics, 6(1):

142–145, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900136.

Schumann:1975:LSF

[Sch75a] U. Schumann. Linear stability of finite difference equations for three-dimensional flow problems. Journal of Computational Physics, 18(4):465–470, August 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900972.

Schumann:1975:SSM

[Sch75b] U. Schumann. Subgrid scale model for finite difference simulations of turbulent flows in plane channels and annuli. Journal of Computational Physics, 18(4):376–404, August 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900935.

Schoenstadt:1977:ESD

[Sch77] Arthur L. Schoenstadt. The effect of spatial discretization on the steady-state and transient solutions of a dispersive wave equation. Journal of Computational Physics, 23(4):364–379, April 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900687.

Scott:1973:IVM

[Sco73] Melvin R. Scott. An initial value method for the eigenvalue problem for systems of ordinary differential equations. Journal of Computational Physics, 12(3):334–347, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901563.

Scott:1979:CSF

[Sco79] T. Scott. Computation of scattering functions in second-order many-body theory. Journal of Computational Physics, 33(3):369– 381, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901621.

Schamel:1976:ASM

[SE76] Hans Schamel and Klaus Elsässer. The application of the spectral method to nonlinear wave propagation. Journal of Computational Physics, 22(4):501–516, December 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900462.

Stellman:1971:ECP

[SFG71] Steven D. Stellman, Mark Froimowitz, and Paul J. Gans. Efficient computation of polymer conformation energy. Journal of Computational Physics, 7(1):178–181, February 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900581.

Shoucri:1976:NSV

[SG76] M. Shoucri and R. R. J. Gagne. Numerical solution of the Vlasov equation by transform methods. Journal of Computational Physics, 21(2):238–242, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900140.

Shoucri:1977:MTN

[SG77a] M. M. Shoucri and R. R. T. Gagne. A multistep technique for the numerical solution of a two-dimensional Vlasov equation. Journal of Computational Physics, 23(3):242–262, March 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900936.

Shoucri:1977:NST

[SG77b] Magdi M. Shoucri and Réal R. J. Gagné. Numerical solution of a two-dimensional Vlasov equation. Journal of Computational Physics, 25(2):94–103, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900146.

Shoucri:1978:SSN

[SG78] Magdi M. Shoucri and Real R. J. Gagné. Splitting schemes for the numerical solution of a two-dimensional Vlasov equation. Journal of Computational Physics, 27(3):315–322, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917890013X.

Srinivas:1979:IFA

[SG79] K. Srinivas and J. Gururaja. An improved form of the artificial diffusion parameter — χ. Journal of Computational Physics, 31(2): 289–292, May 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900731.

Srinivas:1976:AQS

[SGP76] K. Srinivas, J. Gururaja, and K. Krishna Prasad. An assessment of the quality of selected finite difference schemes for time dependent compressible flows. Journal of Computational Physics, 20(2):140– 159, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900607.

Salsburg:1971:FEF

[SH71a] Zevi W. Salsburg and Dale A. Huckaby. Free energy of the fcc and hcp lattices under first- and second-neighbor harmonic interactions. Journal of Computational Physics, 7(3):489–501, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171901069.

Silvester:1971:PSI

[SH71b] P. Silvester and M. S. Hsieh. Projective solution of integral equations arising in electric and magnetic field problems. Journal of Computational Physics, 8(1):73–82, August 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999171900362.

Shakeshaft:1970:ARE

[Sha70a] Robin Shakeshaft. The analytical and recursive evaluation of two-center integrals. Journal of Computational Physics, 5(2):345– 349, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900665.

Shavitt:1970:MNA

[Sha70b] Isaiah Shavitt. Modification of Nesbet’s algorithm for the iterative evaluation of eigenvalues and eigenvectors of large matrices. Journal of Computational Physics, 6(1):124–130, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900100. See [Nes65].

Shapiro:1972:ILC

[Sha72] Ralph Shapiro. Information loss and compensation in linear interpolation. Journal of Computational Physics, 10(1):65–84, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900915.

Shestakov:1979:HVA

[She79] Aleksei T. Shestakov. A hybrid vortex-ADI solution for flows of low viscosity. Journal of Computational Physics, 31(3):313– 334, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900500.

Shockley:1978:NID

[Sho78] Richard C. Shockley. Numerical instability due to the Lorentz force in a relativistic fluid, laser-plasma code. Journal of Computational Physics, 29(2):270–277, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901560.

Shultis:1973:NMC

[Shu73] J. Kenneth Shultis. A new method for the calculation of the emergent distributions for the anisotropic slab albedo problem. Journal of Computational Physics, 11(1):109–126, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901502.

Singhal:1972:EIB

[Sin72] S. P. Singhal. Evaluation of integrals over the Brillouin zone by Houston’s method. Journal of Computational Physics, 10(2):316– 323, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900691.

Sato:1972:MDE

[SK72] Yasuo Satô and Kasaburo Karumi. A method for determining the eigenfrequency of a mechanical system by simulation. Journal of Computational Physics, 9(2):194–206, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900186.

Shoucri:1974:NIV

[SK74a] Magdi Shoucri and Georg Knorr. Numerical integration of the Vlasov equation. Journal of Computational Physics, 14(1):84– 92, January 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900060.

Speziale:1974:UMQ

[SK74b] Charles G. Speziale and Robert P. Kirchner. On the use of the method of quadrature by differentiation for solving eigenvalue problems in hydrodynamic stability. Journal of Computational Physics, 16(4):315–323, December 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900424.

Salu:1975:USM

[SK75] Yehuda Salu and Georg Knorr. Use of the spectral method for twoand three-dimensional guiding center plasmas. Journal of Computational Physics, 17(1):68–78, January 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900637.

Segur:1977:NMS

[SK77] Pierre Segur and Robert Keller. A numerical method of solution of the Boltzmann equation in a weakly ionized heterogeneous gas. Journal of Computational Physics, 24(1):43–69, May 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177901097.

Shin:1978:TDF

[SK78] Yong Woo Shin and Christian A. Kot. Two-dimensional fluidtransient analysis by the method of nearcharacteristics. Journal of Computational Physics, 28(2):211–231, August 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178900359.

Sadowski:1972:UOA

[SL72] W. L. Sadowski and D. W. Lozier. Use of Olver’s algorithm to evaluate certain definite integrals of plasma physics involving Chebyshev polynomials. Journal of Computational Physics, 10(3):607– 613, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917290054X.

Sloan:1977:ESL

[Slo77] D. M. Sloan. Eigenfunctions of systems of linear ordinary differential equations with separated boundary conditions using Riccati transformations. Journal of Computational Physics, 24(3): 320–330, July 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900407.

Smith:1974:NPB

[Smi74] Warwick D. Smith. A nonreflecting plane boundary for wave propagation problems. Journal of Computational Physics, 15(4):492– 503, August 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900758.

Smith:1975:NAS

[Smi75a] Ed R. Smith. A numerical approach to the solution of coupled equations involved in electron-ion and electron-atom scattering problems. Journal of Computational Physics, 18(2):201– 223, June 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900297.

Smith:1975:EAA

[Smi75b] Jimmie H. Smith. Engineering applications of ADI methods to piecewise linear multidimensional heat transfer. Journal of Computational Physics, 17(2):181–191, February 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900352.

Snider:1972:CDP

[Sni72] Arthur D. Snider. Correction to a dimensional perturbation calculation of the contraction coefficient. Journal of Computational Physics, 9(2):385–386, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900320.

Sod:1977:EEH

[Sod77] Gary A. Sod. Eigenvalues and eigenfunctions of the Hilbert operator. Journal of Computational Physics, 23(2):179–186, February 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177901206.

Sod:1978:SSF

[Sod78] Gary A. Sod. A survey of several finite difference methods for systems of nonlinear hyperbolic conservation laws. Journal of Computational Physics, 27(1):1–31, April 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900232.

Smith:1972:AHN

[SOH72] E. R. Smith, R. S. Oberoi, and Ronald J. W. Henry. Application of the Harris–Nesbet method to a dipole coupling potential. Journal of Computational Physics, 10(1):53–64, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900903.

Squire:1970:RAI

[Squ70] William Squire. A rational approximation to an integral appearing in glow curve theory. Journal of Computational Physics, 6(1):152– 153, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900161.

Starling:1971:TCL

[SS71] Kenneth E. Starling and Zevi W. Salsburg. Transport coefficients of liquids-approximate self-diffusion coefficient in a cell model. Journal of Computational Physics, 7(3):554–565, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171901112.

Schumann:1976:DMS

[SS76] U. Schumann and Roland A. Sweet. A direct method for the solution of Poisson’s equation with Neumann boundary conditions on a staggered grid of arbitrary size. Journal of Computational Physics, 20(2):171–182, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900620.

Skollermo:1978:FAS

[SS78] Anders Sköllermo and Gunilla Sköllermo. A Fourier analysis of some difference schemes for the Laplace equation in a system of rotational symmetry. Journal of Computational Physics, 28(1): 103–114, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900505.

Steger:1979:AMP

[SS79] J. L. Steger and R. L. Sorenson. Automatic mesh-point clustering near a boundary in grid generation with elliptic partial differential equations. Journal of Computational Physics, 33(3):405– 410, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901657.

Sato:1976:FEM

[ST76] Atuo Sato and Erik G. Thompson. Finite element models for creeping convection. Journal of Computational Physics, 22(2):229– 244, October 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900772.

Stern:1970:IMC

[Ste70] Frank Stern. Iteration methods for calculating self-consistent fields in semiconductor inversion layers. Journal of Computational Physics, 6(1):56–67, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900045.

Stecki:1971:KEN

[Ste71] J. Stecki. On the kinetic equation nonlocal in time for the generalized self-diffusion process. Journal of Computational Physics, 7(3): 547–553, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171901100.

Steppeller:1975:HAF

[Ste75] J. Steppeller. On a high accuracy finite difference method. Journal of Computational Physics, 19(4):390–403, December 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900704.

Steppeler:1976:AST

[Ste76a] J. Steppeler. The application of the second and third degree methods. Journal of Computational Physics, 22(3):295–318, November 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900516.

Stern:1976:NEP

[Ste76b] M. S. Stern. Numerical evaluation of partial-wave born approximations. Journal of Computational Physics, 21(1):38–54, May 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917690019X.

Stern:1977:CNS

[Ste77a] M. S. Stern. Comparison of numerical solutions of the partial-wave Schrödinger differential and integral equations. Journal of Computational Physics, 25(1):56–70, September 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900250.

Sternlieb:1977:INQ

[Ste77b] A. Sternlieb. Initial noise and quiet start (Q.S.) in electromagnetic particle simulation codes. Journal of Computational Physics, 25(2):118–132, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917790016X.

Steffen:1978:NME

[Ste78a] Bernhard Steffen. A numerical method for the evaluation of an equilibrium configuration of a toroidal pinch. Journal of Computational Physics, 27(2):180–191, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900037.

Stern:1978:AGM

[Ste78b] M. S. Stern. Application of the EI-Gendi method to the Schrödinger integral equation. Journal of Computational Physics, 28(1):122– 128, July 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900529.

Stefpeler:1979:DSU

[Ste79a] J. Stefpeler. Difference schemes with uniform second and third order accuracy and reduced smoothing. Journal of Computational Physics, 31(3):438–449, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900561.

Stewart:1979:CTB

[Ste79b] H. Bruce Stewart. Calculation of transient boiling flow in channels. Journal of Computational Physics, 30(1):61–75, January 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900871.

Stewart:1979:STP

[Ste79c] H. Bruce Stewart. Stability of two-phase flow calculation using two-fluid models. Journal of Computational Physics, 33(2):259– 270, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900202.

Stillinger:1971:PDC

[Sti71] Frank H. Stillinger. Pair distribution in the classical rigid disk and sphere systems. Journal of Computational Physics, 7(3):367– 384, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900982.

Stoutemyer:1977:DAU

[Sto77] David R. Stoutemyer. Dimensional analysis, using computer symbolic mathematics. Journal of Computational Physics, 24(2):141– 149, June 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900547.

Stoddard:1978:ICC

[Sto78] Spotswood D. Stoddard. Identifying clusters in computer experiments on systems of particles. Journal of Computational Physics, 27(2):291–293, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178900116.

Stroud:1977:IFW

[Str77] A. H. Stroud. Integration formulas for the wave equation in n space dimensions. Journal of Computational Physics, 25(1):32– 55, September 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900249.

Strauss:1978:AVT

[Str78] H. R. Strauss. An artificial viscosity for two-dimensional hydrodynamics. Journal of Computational Physics, 28(3):437–438, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900657.

Stuart:1978:DCC

[Stu78] S. N. Stuart. Depolarization correction for Coulomb lattice sums. Journal of Computational Physics, 29(1):127–132, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178901134.

Sullivan:1972:GSG

[Sul72] J. K. Sullivan. Generalized solution of the glow curve integral applied to nonlinear heating rates. Journal of Computational Physics, 9(1):164–166, February 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900423.

Sundstrom:1975:NPB

[Sun75] Arne Sundström. Note on the paper “Boundary conditions in finite difference fluid dynamic codes” by C. K. Chu and Aron Sereny. Journal of Computational Physics, 17(4):450–454, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900509. See [CS74].

Shin:1976:NAF

[SV76] Yong Woo Shin and Richard A. Valentin. Numerical analysis of fluid-hammer waves by the method of characteristics. Journal of Computational Physics, 20(2):220–237, February 1976. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999176900668.

Strauss:1978:NSN

[SV78] Walter Strauss and Luis Vazquez. Numerical solution of a nonlinear Klein–Gordon equation. Journal of Computational Physics, 28(2): 271–278, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900384.

Sielecki:1970:NIN

[SW70] Anita Sielecki and M. G. Wurtele. The numerical integration of the nonlinear shallow-water equations with sloping boundaries. Journal of Computational Physics, 6(2):219–236, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999170900227.

Steger:1975:CCF

[SW75] J. L. Steger and R. F. Warming. On the convergence of certain finite-difference schemes by an inverse-matrix method. Journal of Computational Physics, 17(2):103–121, February 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900315.

Schoenstadt:1976:CSP

[SW76] A. L. Schoenstadt and R. T. Williams. The computational stability properties of the Shuman pressure gradient averaging technique. Journal of Computational Physics, 21(2):166–177, June 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900097.

Smith:1979:NSD

[SW79] Julius Smith and J. C. Whitson. On the numerical solution of the drift wave equations by means of invariant imbedding. Journal of Computational Physics, 33(1):102–117, October 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179900305.

Swarztrauber:1974:DSD

[Swa74] Paul N. Swarztrauber. The direct solution of the discrete Poisson equation on the surface of a sphere. Journal of Computational Physics, 15(1):46–54, May 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900680.

Sweet:1973:DMS

[Swe73] Roland A. Sweet. Direct methods for the solution of Poisson’s equation on a staggered grid. Journal of Computational Physics, 12(3): 422–428, July 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173901642.

Sharma:1977:GMA

[SZ77] R. R. Sharma and Bahman Zohuri. A general method for an accurate evaluation of exponential integrals E_{1}(x), x > 0. Journal of Computational Physics, 25(2):199–204, October 1977. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999177900225.

Takizuka:1977:BCM

[TA77] Tomonor Takizuka and Hirotada Abe. A binary collision model for plasma simulation with a particle code. Journal of Computational Physics, 25(3):205–219, November 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900997.

Turner:1973:NFC

[TAE73] Emily B. Turner, W. Dallas Adams, and George Emanuel. Numerical formulation for constant-gain chemical laser calculations. Journal of Computational Physics, 11(1):15–27, January 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173901459.

Taff:1976:NSC

[Taf76] L. G. Taff. Numerical solution of the collisionless Boltzmann equation for stellar systems. Journal of Computational Physics, 20(2):160–170, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900619.

Turkel:1976:MDS

[TAG76] E. Turkel, S. Abarbanel, and D. Gottlieb. Multidimensional difference schemes with fourth-order accuracy. Journal of Computational Physics, 21(1):85–113, May 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900218.

Takagi:1971:NDS

[Tak71] Shunsuke Takagi. Numerical differentiation by spline functions applied to a lake temperature observation. Journal of Computational Physics, 8(2):285–291, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900088.

Takeda:1976:NMS

[Tak76] Tatsuoki Takeda. Numerical method for solution of the integral equation of the first kind — application to analysis of plasma density profile. Journal of Computational Physics, 21(3):305– 318, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900279.

Talman:1978:NFB

[Tal78] James D. Talman. Numerical Fourier and Bessel transforms in logarithmic variables. Journal of Computational Physics, 29(1):35– 48, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901079.

Tang:1979:CSM

[Tan79] Cha-Mei Tang. Comparison of spectral methods for flows on spheres for flows on spheres. Journal of Computational Physics, 32 (1):80–88, July 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901438.

Taylor:1970:LSN

[Tay70] P. J. Taylor. The linear stability of no-slip boundary conditions in the numerical solution of nonsteady fluid flows. Journal of Computational Physics, 6(2):268–287, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900252.

Temme:1975:NEM

[Tem75a] Nico M. Temme. On the numerical evaluation of the modified Bessel function of the third kind. Journal of Computational Physics, 19(3):324–337, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900820.

Temperton:1975:ASC

[Tem75b] Clive Temperton. Algorithms for the solution of cyclic tridiagonal systems. Journal of Computational Physics, 19(3):317–323, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900819.

Temme:1976:NEO

[Tem76] Nico M. Temme. On the numerical evaluation of the ordinary Bessel function of the second kind. Journal of Computational Physics, 21(3):343–350, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999176900322.

Temme:1979:AAP

[Tem79a] N. M. Temme. An algorithm with ALGOL 60 program for the computation of the zeros of ordinary Bessel functions and those of their derivatives. Journal of Computational Physics, 32(2):270– 279, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901347.

Temperton:1979:DMS

[Tem79b] Clive Temperton. Direct methods for the solution of the discrete Poisson equation: Some comparisons. Journal of Computational Physics, 31(1):1–20, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900597.

Terras:1979:DIG

[Ter79a] Riho Terras. The determination of incomplete gamma functions through analytic integration. Journal of Computational Physics, 31(1):146–151, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900664.

Terras:1979:REG

[Ter79b] Riho Terras. Reliable evaluation of Gaussian integrals. Journal of Computational Physics, 33(3):411–416, December 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999179901669.

Talmi:1977:MSA

[TG77] A. Talmi and G. Gilat. Method for smooth approximation of data. Journal of Computational Physics, 23(2):93–123, February 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177901152.

Thiele:1978:ANS

[Thi78] F. Thiele. Accurate numerical solutions of boundary layer flows by the finite-difference method of Hermitian type. Journal of Computational Physics, 27(1):138–159, April 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917890030X.

Taylor:1978:FES

[THM78] C. Taylor, T. G. Hughes, and K. Morgan. Finite element solution of one-equation models of turbulent flow. Journal of Computational Physics, 29(2):163–172, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901493.

Thomas:1973:NSS

[Tho73] Lowell D. Thomas. Numerical solution of scattering equations with nonlocal potentials. Journal of Computational Physics, 13(3):348– 362, November 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999173900417.

Tani:1973:MVP

[TI73] Smio Tani and Mitio Inokuti. Modification of the variationof-parameters method and integration of the Schrödinger equation. Journal of Computational Physics, 11(3):409–422, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173900818.

Theocaris:1979:MNS

[TI79] Pericles S. Theocaris and Nikolaos I. Ioakimidis. A method of numerical solution of Cauchy-type singular integral equations with generalized kernels and arbitrary complex singularities. Journal of Computational Physics, 30(3):309–323, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901177.

Thomas:1973:CFI

[TLM73] John D. Thomas, H. Ralph Lewis, and K. J. Melendez. Computation of Fourier integrals of exponentials of truncated Fourier series. Journal of Computational Physics, 12(4):462–470, August 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999173900983.

Taylor:1970:AUN

[TM70] Thomas D. Taylor and Bruce S. Masson. Application of the unsteady numerical method of Godunov to computation of supersonic flows past bell-shaped bodies. Journal of Computational Physics, 5(3):443–454, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900732.

Taylor:1972:SNM

[TNM72] T. D. Taylor, E. Ndefo, and B. S. Masson. A study of numerical methods for solving viscous and inviscid flow problems. Journal of Computational Physics, 9(1):99–119, February 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900381.

Tuann:1978:RCM

[TO78] Shih-Yu Tuann and Mervyn D. Olson. Review of computing methods for recirculating flows. Journal of Computational Physics, 29(1):1–19, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901055.

Tomishima:1979:MCS

[TO79] Yasuo Tomishima and Jiro Ozaki. Monte Carlo solution of Schrödinger’s equation for the hydrogen atom in a magnetic field. Journal of Computational Physics, 33(3):382–396, December 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999179901633.

Tolsma:1975:FMN

[Tol75] L. D. Tolsma. A fast method for nuclear coupled-channels calculations including Coulomb excitation. Journal of Computational Physics, 17(4):384–400, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900418.

Thompson:1972:ESS

[TR72] Philip D. Thompson and Henrik Rye. Equilibrium statistics of the simplest viscous flow with random forcing. Journal of Computational Physics, 9(1):1–9, February 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900332.

Trapp:1976:SHM

[TR76] John A. Trapp and John D. Ramshaw. A simple heuristic method for analyzing the effect of boundary conditions on numerical stability. Journal of Computational Physics, 20(2):238– 242, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917690067X.

Trafton:1971:RNS

[Tra71] L. Trafton. A rapid numerical solution to the radial Schrödinger equation in the oscillatory region. Journal of Computational Physics, 8(1):64–72, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999171900350.

Troesch:1976:SAS

[Tro76] B. A. Troesch. A simple approach to a sensitive two-point boundary value problem. Journal of Computational Physics, 21(3):279– 290, July 1976. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999176900255.

Truhlar:1972:FDB

[Tru72] Donald G. Truhlar. Finite difference boundary value method for solving one-dimensional eigenvalue equations. Journal of Computational Physics, 10(1):123–132, August 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900940.

Tsunematsu:1978:NIM

[TT78] Toshhide Tsunematsu and Tatsuoki Takeda. A new iterative method for solution of a large-scale general eigenvalue problem. Journal of Computational Physics, 28(2):287–293, August 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900402.

Thompson:1974:ANG

[TTM74] Joe F. Thompson, Frank C. Thames, and C. Wayne Mastin. Automatic numerical generation of body-fitted curvilinear coordinate system for field containing any number of arbitrary twodimensional bodies. Journal of Computational Physics, 15(3): 299–319, July 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174901144.

Thompson:1977:TCN

[TTM77] Joe F. Thompson, Frank C. Thames, and C. Wayne Mastin. TOM-CAT — a code for numerical generation of boundary-fitted curvilinear coordinate systems on fields containing any number of arbitrary two-dimensional bodies. Journal of Computational Physics, 24(3): 274–302, July 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900389.

Thames:1977:NSV

[TTMW77] Frank C. Thames, Joe F. Thompson, C. Wayne Mastin, and Ray L. Walker. Numerical solutions for viscous and potential flow about arbitrary two-dimensional bodies using body-fitted coordinate systems. Journal of Computational Physics, 24(3):245– 273, July 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900377.

Turkel:1974:PES

[Tur74] Eli Turkel. Phase error and stability of second order methods for hyperbolic problems. I. Journal of Computational Physics, 15(2): 226–250, June 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900862.

Torrie:1977:NSD

[TV77] G. M. Torrie and J. P. Valleau. Nonphysical sampling distributions in Monte Carlo free-energy estimation: Umbrella sampling. Journal of Computational Physics, 23(2):187–199, February 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177901218.

Twomey:1975:CCL

[Two75] S. Twomey. Comparison of constrained linear inversion and an iterative nonlinear algorithm applied to the indirect estimation of particle size distributions. Journal of Computational Physics, 18(2): 188–200, June 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999175900285.

Uncles:1974:NAE

[Unc74] R. J. Uncles. Numerical approximation of the equations describing a partially ionized plasma. Journal of Computational Physics, 16(2):93–118, October 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174901065.

VanWyk:1970:VMM

[Van70] Rodney Van Wyk. Variable mesh multistep methods for ordinary differential equations. Journal of Computational Physics, 5(2):244– 264, April 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900628.

VanderWerff:1973:PMC

[Van73] Terry J. Vander Werff. Periodic method of characteristics. Journal of Computational Physics, 11(2):296–305, February 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900089.

VanDooren:1974:SCC

[Van74] R. Van Dooren. Stabilization of Cowell’s classical finite difference method for numerical integration. Journal of Computational Physics, 16(2):186–192, October 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174901120.

VanLeer:1977:TUCa

[Van77a] Bram Van Leer. Towards the ultimate conservative difference scheme III. Upstream-centered finite-difference schemes for ideal compressible flow. Journal of Computational Physics, 23(3):263– 275, March 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900948.

VanLeer:1977:TUCb

[Van77b] Bram Van Leer. Towards the ultimate conservative difference scheme. IV. A new approach to numerical convection. Journal of Computational Physics, 23(3):276–299, March 1977. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917790095X.

Vesely:1977:PDP

[Ves77] Franz J. Vesely. N-particle dynamics of polarizable Stockmayertype molecules. Journal of Computational Physics, 24(4):361– 371, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900286.

Viecelli:1971:CMI

[Vie71] James A. Viecelli. A computing method for incompressible flows bounded by moving walls. Journal of Computational Physics, 8(1): 119–143, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900398.

Viecelli:1973:LCD

[Vie73] J. A. Viecelli. The linear Q and the calculation of decaying spherical shocks in solids. Journal of Computational Physics, 12(2):187– 201, February 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999173800105.

Vinokur:1974:CEG

[Vin74] Marcel Vinokur. Conservation equations of gasdynamics in curvilinear coordinate systems. Journal of Computational Physics, 14(2):105–125, February 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999174900084.

vanLeer:1974:TUC

[vL74] Bram van Leer. Towards the ultimate conservative difference scheme. II. Monotonicity and conservation combined in a secondorder scheme. Journal of Computational Physics, 14(4):361– 370, March 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900199.

vanLeer:1979:TUC

[vL79] Bram van Leer. Towards the ultimate conservative difference scheme. V. A second-order sequel to Godunov’s method. Journal of Computational Physics, 32(1):101–136, July 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999179901451.

Vojta:1973:NAN

[Voj73] V. Vojta. Notes on the article “Numerical Laplace transform inversion of a function arising in viscoelasticity”. Journal of Computational Physics, 13(4):611–612, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900363.

See [Lon72].

Vujisic:1975:FPR

[VP75] B. R. Vujisić and D. S. Pesić. A FORTRAN program for rotational analysis of the spectra of diatomic molecules. Journal of Computational Physics, 19(4):432–434, December 1975. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917590073X.

Vreugdenhil:1979:TLS

[Vre79] C. B. Vreugdenhil. Two-layer shallow-water flow in two dimensions, a numerical study. Journal of Computational Physics, 33(2):169– 184, November 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900147.

Vecheslavov:1977:APE

[VT77] V. V. Vecheslavov and N. I. Tolstobrova. Analysis-of plane electrostatic fields in doubly connected polygonal domains using conformal mappings. Journal of Computational Physics, 24(2):133– 140, June 1977. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999177900535.

Valleau:1977:MCS

[VW77] John P. Valleau and Stuart G. Whittington. Monte Carlo in statistical mechanics: Choosing between alternative transition matrices. Journal of Computational Physics, 24(2):150–157, June 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900559.

Walker:1975:RCR

[Wal75] R. B. Walker. Rapid computation of rotation functions. Journal of Computational Physics, 17(4):437–439, April 1975. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999175900479.

Ward:1978:NCL

[War78] Richard A. Ward. Numerical calculation of the limiting properties of Coulomb excitation functions. Journal of Computational Physics, 29(1):49–66, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901080.

Warnock:1979:NIE

[War79] Robert Lee Warnock. A nonlinear integral equation from the Ball– Zachariasen model of diffractive scattering: Numerical solution near a singularity of the Fréchet derivative. Journal of Computational Physics, 33(1):45–69, October 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900275.

Wasserstrom:1972:NMS

[Was72] E. Wasserstrom. A new method for solving eigenvalue problems. Journal of Computational Physics, 9(1):53–74, February 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999172900368.

Watterstrom:1971:RFP

[Wat71] E. Watterstrom. Root finding of polynomials as an initial value problem. Journal of Computational Physics, 8(2):304–308, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917190012X.

Watanabe:1975:NMT

[Wat75] D. S. Watanabe. A numerical method for the time-dependent Krook kinetic equation. Journal of Computational Physics, 19(1):29–42, September 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917590114X.

Wilks:1977:CEA

[WB77] G. Wilks and J. S. Bramley. On the computation of eigenvalues arising out of perturbations of the Blasius profile. Journal of Computational Physics, 24(3):303–319, July 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900390.

Wichterle:1971:DCE

[WCK71] Ivan Wichterle, Patsy S. Chappelear, and Rik Kobayashi. Determination of critical exponents from measurements of binary vapor-liquid equilibrium in the neighborhood of the critical line. Journal of Computational Physics, 7(3):606–620, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171901161.

Walkden:1978:LTD

[WCL78] F. Walkden, P. Caine, and G. T. Laws. A locally two-dimensional numerical method for calculating three-dimensional supersonic flows. Journal of Computational Physics, 27(1):103–122, April 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900281.

Wang:1978:SII

[WD78] T. S. Wang and D. Dobrott. Suppression of inherent instabilities in initial value methods. Journal of Computational Physics, 29(1): 93–100, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178901110.

Wilhelmson:1977:DSP

[WE77a] Robert B. Wilhelmson and James H. Ericksen. Direct solutions for Poisson’s equation in three dimensions. Journal of Computational Physics, 25(4):319–331, December 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900018.

Wiscombe:1977:ESF

[WE77b] W. J. Wiscombe and J. W. Evans. Exponential-sum fitting of radiative transmission functions. Journal of Computational Physics, 24(4):416–444, August 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999177900316.

Weiland:1978:CTD

[Wei78] C. Weiland. Calculation of three-dimensional stationary supersonic flow fields by applying the “progonka” process to a conservative formulation of the governing equations. Journal of Computational Physics, 29(2):173–198, November 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917890150X.

Westbrook:1978:GIM

[Wes78] Charles K. Westbrook. A generalized ICE method for chemically reactive flows in combustion systems. Journal of Computational Physics, 29(1):67–80, October 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901092.

Warme:1972:RXR

[WGS72] Paul K. Warme, Nobuhiro Go, and Harold A. Scheraga. Refinement of X-ray data on proteins. I. Adjustment of atomic coordinates to conform to a specified geometry. Journal of Computational Physics, 9(2):303–317, April 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999172900253.

Warming:1974:MEA

[WH74] R. F. Warming and B. J. Hyett. The modified equation approach to the stability and accuracy analysis of finite-difference methods. Journal of Computational Physics, 14(2):159–179, February 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174900114.

Wnek:1975:ENS

[WH75] Walter J. Wnek and Willard R. Haas. The equivalence and numerical solution of regression, optimal and boundary value problems involving differential equations. Journal of Computational Physics, 18(3):326–341, July 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999175900066.

See comment [Cel76].

Whitney:1970:FDM

[Whi70] James C. Whitney. Finite difference methods for the Fokker– Planck equation. Journal of Computational Physics, 6(3):483– 509, December 1970. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999170900446.

White:1973:NFAa

[Whi73a] John W. White. A new form of artificial viscosity. Journal of Computational Physics, 11(4):573–590, April 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901381.

White:1973:NFAb

[Whi73b] John W. White. A new form of artificial viscosity: Postscript. Journal of Computational Physics, 12(4):553–556, August 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173901058.

White:1974:NFA

[Whi74] John W. White. A new form of artificial viscosity for elastic solids. Journal of Computational Physics, 16(2):119–126, October 1974. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999174901077.

White:1979:ICS

[Whi79] R. B. White. An interactive code for solving differential equations, using phase integral methods. Journal of Computational Physics, 31(3):409–424, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900548.

Wilkins:1970:FDS

[Wil70] Mark L. Wilkins. Finite difference scheme for calculating problems in two space dimensions and time. Journal of Computational Physics, 5(3):406–414, June 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900719.

Williamson:1971:CFS

[Wil71a] David L. Williamson. A comparison of first- and second-order difference approximations over a spherical geodesic grid. Journal of Computational Physics, 7(2):301–309, April 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199917190091X.

Williamson:1971:IPD

[Wil71b] J. H. Williamson. Initial particle distributions for simulated plasma. Journal of Computational Physics, 8(2):258–267, October 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999171900064.

Wills:1971:URR

[Wil71c] John G. Wills. On the use of recursion relations in the numerical evaluation of spherical Bessel functions and Coulomb functions. Journal of Computational Physics, 8(1):162–166, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917190043X.

Wilson:1978:NLF

[Wil78] J. C. Wilson. A note on the leap frog schemes in any number of space variables. Journal of Computational Physics, 28(3):433– 436, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900645.

Wimp:1978:CBT

[Wim78] Jet Wimp. The computation of Borel-type sums arising in scattering theory. Journal of Computational Physics, 27(2):284– 290, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999178900104.

Winterbon:1978:NCC

[Win78] K. B. Winterbon. Note: Comment on “Comparison of some methods for evaluating infinite range oscillatory integrals”. Journal of Computational Physics, 26(3):447–448, March 1978. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999178900803. See [BEH76].

Winsor:1970:NMT

[WJD70] Niels K. Winsor, John L. Johnson, and John M. Dawson. A numerical model for toroidal plasma containment with flow. Journal of Computational Physics, 6(3):430–448, December 1970. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999170900410.

Whealton:1978:OIB

[WJW78] J. H. Whealton, E. F. Jaeger, and J. C. Whitson. Optics of ion beams of arbitrary perveance extracted from a plasma. Journal of Computational Physics, 27(1):32–41, April 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900244.

Whitlock:1979:MCC

[WK79] P. A. Whitlock and M. H. Kalos. Monte Carlo calculation of the radial distribution function of quantum hard spheres at finite temperatures. Journal of Computational Physics, 30(3):361– 388, March 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179901219.

Wood:1971:MCC

[WL71] W. W. Wood and F. Lado. Monte Carlo calculation of normal and abnormal diffusion in Ehrenfest’s wind-tree model. Journal of Computational Physics, 7(3):528–546, June 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999171901094.

Wait:1971:CSE

[WM71] R. Wait and A. R. Mitchell. Corner singularities in elliptic problems by finite element methods. Journal of Computational Physics, 8(1): 45–52, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900337.

Wentland:1972:NSH

[WM72] N. L. Wentland and J. R. Merrill. Numerical solutions to the helicon boundary value problem. Journal of Computational Physics, 9(1):138–158, February 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917290040X.

Wellford:1975:DFE

[WO75] L. C. Wellford, Jr. and J. T. Oden. Discontinuous finite-element approximations for the analysis of shock waves in nonlinearly elastic materials. Journal of Computational Physics, 19(2):179– 210, October 1975. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917590087X.

Wolniewicz:1978:NIS

[WO78] L. Wolniewicz and T. Orlikowski. On the numerical integration of the Schrödinger equation with a double-minimum potential. Journal of Computational Physics, 27(2):169–179, May 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999178900025.

Warme:1973:RRD

[WS73] Paul K. Warme and Harold A. Scheraga. Refinement of X-ray data on proteins. II. Adjustment of structure of specified geometry to relieve atomic overlaps. Journal of Computational Physics, 12(1): 49–64, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917390168X.

Wengle:1978:PSA

[WS78] Hans Wengle and John H. Seinfeld. Pseudospectral solution of atmospheric diffusion problems. Journal of Computational Physics, 26(1):87–106, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999178901018.

Webster:1974:LPE

[WSB74] Joseph B. Webster III, Brian G. Stephan, and Charles J. Bridgman. Legendre polynomial expansion for the Klein–Nishina formula for nonstationary electrons. Journal of Computational Physics, 14(1): 29–39, January 1974. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999174900035.

Whitson:1978:CII

[WSW78] J. C. Whitson, J. Smith, and J. H. Whealton. Calculations involving ion beam source. Journal of Computational Physics, 28(3):408– 415, September 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199917890061X.

Yee:1970:CDN

[Yee70] Samuel Y. K. Yee. Comments on “A direct numerical solution to a one-dimensional Helmholtz equation”. Journal of Computational Physics, 6(2):336–337, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900331.

See [O’B69].

Yee:1972:ENS

[Yee72] Samuel Y. K. Yee. An efficient numerical scheme for a baroclinic quasi-geostrophic model. Journal of Computational Physics, 9(3): 412–423, June 1972. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999172900034.

Yee:1976:EMF

[Yee76] Samuel Y. K. Yee. An efficient method for a finite-difference solution of the Poisson equation on the surface of a sphere. Journal of Computational Physics, 22(2):215–228, October 1976. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999176900760.

Yeh:1971:RKP

[YK71] Yin Yeh and R. N. Keeler. Reaction kinetics probed by dynamic light scattering. Journal of Computational Physics, 7(3): 566–575, June 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171901124.

Yen:1972:QTS

[YL72] Ching-Se Yen and Russell D. Larsen. Quasilinearization technique for solution of unsteady state diffusion problems. Journal of Computational Physics, 10(3):554–567, December 1972. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999172900526.

Youakim:1971:GRH

[YLY71] M. Y. Youakim, C. H. Liu, and K. C. Yeh. Generation of random hypersurfaces. Journal of Computational Physics, 8(3):382– 391, December 1971. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999171900180.

Yoshimine:1973:CHM

[Yos73] M. Yoshimine. Construction of the Hamiltonian matrix in large configuration interaction calculations. Journal of Computational Physics, 11(3):449–454, March 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173900855.

Zalesak:1979:FMF

[Zal79] Steven T. Zalesak. Fully multidimensional flux-corrected transport algorithms for fluids. Journal of Computational Physics, 31(3): 335–362, June 1979. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999179900512.

Zaroodny:1973:VSA

[ZG73] Serge J. Zaroodny and Michael D. Greenberg. On a vortex sheet approach to the numerical calculation of water waves. Journal of Computational Physics, 11(3):440–446, March 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999173900831.

Zabusky:1979:CDE

[ZHR79] Norman J. Zabusky, M. H. Hughes, and K. V. Roberts. Contour dynamics for the Euler equations in two dimensions. Journal of Computational Physics, 30(1):96–106, January 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999179900895.

Zinn:1973:FDS

[Zin73] John Zinn. A finite difference scheme for time-dependent spherical radiation hydrodynamics problems. Journal of Computational Physics, 13(4):569–590, December 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199917390034X.

Zwas:1973:ENT

[ZR73] Gideon Zwas and Joseph Roseman. The effect of nonlinear transformations on the computation of weak solutions. Journal of Computational Physics, 12(2):179–186, February 1973. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999173800099.